GUILE NCURSES

For version 1.6, 12 October 2014

Michael L. Gran Pradeep Padala Tom Dickey and the ncurses Team

This manual is for the GNU Guile-NCurses library, (version 1.6, 12 October 2014). It provides bindings for the ncurses library, which provides functionality for updating character screens. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1998-2010,2011,2013,2014 Free Software Foundation, Inc. Copyright \begin{matrix}c \\ ◯ \end{matrix} 2001-2005 Pradeep Padala

Permission is hereby granted, free of charge, to any person obtaining a copy of this software and associated documentation files (the "Software"), to deal in the Software without restriction, including without limitation the rights to use, copy, modify, merge, publish, distribute, distribute with modifications, sublicense, and/or sell copies of the Software, and to permit persons to whom the Software is furnished to do so, subject to the following conditions: The above copyright notice and this permission notice shall be included in all copies or substantial portions of the Software. THE SOFTWARE IS PROVIDED "AS IS", WITHOUT WARRANTY OF ANY KIND, EXPRESS OR IMPLIED, INCLUDING BUT NOT LIMITED TO THE WARRANTIES OF MERCHANTABILITY, FITNESS FOR A PAR-TICULAR PURPOSE AND NONINFRINGEMENT. IN NO EVENT SHALL THE ABOVE COPYRIGHT HOLDERS BE LIABLE FOR ANY CLAIM, DAMAGES OR OTHER LIABILITY, WHETHER IN AN ACTION OF CON-TRACT, TORT OR OTHERWISE, ARISING FROM, OUT OF OR IN CON-NECTION WITH THE SOFTWARE OR THE USE OR OTHER DEALINGS IN THE SOFTWARE. Except as contained in this notice, the name(s) of the above copyright holders shall not be used in advertising or otherwise to promote the sale, use or other dealings in this Software without prior written authorization.

Short Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Components of a Curses Interface . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Types and encodings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4 Curses Tutorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 5 Curses Reference. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Authors, Copyright, and Declarations. . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Components of a Curses Interface. . . . . . . . . . . . . 3

2.1 Terminal Capabilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Keyboards and Keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Characters, Attributes, Colors, and Renditions. . . . . . . . . . . . . . . . . . 3 2.4 Windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3 Types and encodings . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.1 Character types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 Opaque types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4 Curses Tutorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4.1 Hello World!!! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4.1.1 Using the Guile Curses module . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.1.1.1 About initscr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.1.1.2 The mysterious refresh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.1.1.3 About endwin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

4.2 Interactive sessions and guile-ncurses-shell . . . . . . . . . . . . . . . . . 9 4.3 The Gory Details . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.4 Initialization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

4.4.1 Initialization Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.4.1.1 raw! and cbreak! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.4.1.2 echo! and noecho!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.4.1.3 keypad! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.4.1.4 halfdelay! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

4.4.2 An example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.5 A word about windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.6 Output functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

4.6.1 The addch procedure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.6.1.1 Moving the cursor addch . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.6.1.2 Complex characters and addch . . . . . . . . . . . . . . . . . . . . . . 13

4.6.2 addstr class of functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.6.3 A simple addstr example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.6.4 A word about coordinates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

4.7 Input functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.7.1 getch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.7.2 getnstr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.7.3 An example of getnstr and getch . . . . . . . . . . . . . . . . . . . . . . . 15

4.8 Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 4.8.1 The details. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 4.8.2 attr-on! vs attr-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.8.3 attr-get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 4.8.4 chgat function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

4.9 Windows Tutorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 4.9.1 The basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 4.9.2 Let there be a window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 4.9.3 Other functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.9.4 Other border functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

4.10 Colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 4.10.1 The basics of color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 4.10.2 Changing color definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

4.11 Interfacing with the keyboard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.11.1 The basics of keys. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 4.11.2 A simple key usage example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

4.12 Interfacing with the mouse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4.12.1 The basics of the mouse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4.12.2 Getting the events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 4.12.3 Miscellaneous functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

4.13 Panels Library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 4.13.1 Panels Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 4.13.2 Compiling With the Panels Library. . . . . . . . . . . . . . . . . . . . . . 31 4.13.3 Moving and Resizing Panels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

4.14 Menu Library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.14.1 Menu basics. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 4.14.2 Menu Driver: The work horse of the menu system . . . . . . . 36 4.14.3 Menu Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.14.4 Scrolling Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

4.15 Forms Library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.15.1 The Basics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.15.2 A Form Example. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 4.15.3 Making it useful. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 4.15.4 Playing with fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

4.15.4.1 Fetching Size and Location of Field . . . . . . . . . . . . . . . . . 43 4.15.5 Field Display Attributes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 4.15.6 Field Option Bits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 4.15.7 Field Status. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.15.8 Variable-Sized Fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.15.9 Form Windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.15.10 Field Validation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 4.15.11 The Form Driver: The Work Horse of the System. . . . . . . 53

4.15.11.1 Page Navigation Requests . . . . . . . . . . . . . . . . . . . . . . . . . 53 4.15.11.2 Inter-Field Navigation Requests . . . . . . . . . . . . . . . . . . . 53

5 Curses Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

5.1 The naming of routines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 5.2 The basic curses library: (ncurses curses) . . . . . . . . . . . . . . . . . . . 55

5.2.1 Overview. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

5.2.1.1 The Coordinate System . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.2.2 Features determined at compile time. . . . . . . . . . . . . . . . . . . . . . 57 5.2.3 Curses screen initialization and manipulation routines. . . . . 57 5.2.4 Errors and Exceptions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 5.2.5 Making rendered characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 5.2.6 Outputting characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

5.2.6.1 Line Graphics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 5.2.7 Outputting strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 5.2.8 Character and window attribute control routines. . . . . . . . . . 65 5.2.9 Bell and screen flash routines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 5.2.10 Window background manipulation routines . . . . . . . . . . . . . . 67 5.2.11 Borders and lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 5.2.12 Clearing windows and parts of window . . . . . . . . . . . . . . . . . . 69 5.2.13 Color manipulation routines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 5.2.14 Deleting the character under the cursor. . . . . . . . . . . . . . . . . . 71 5.2.15 Deleting and inserting lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 5.2.16 Getting characters from the keyboard . . . . . . . . . . . . . . . . . . . 72 5.2.17 Handling unmapped keys. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 5.2.18 Receiving strings from the keyboard . . . . . . . . . . . . . . . . . . . . . 77 5.2.19 Cursor location and window coordinates . . . . . . . . . . . . . . . . . 77 5.2.20 Getting a rendered character from the window . . . . . . . . . . . 78 5.2.21 Getting a string of characters and attributes from a window

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 5.2.22 Input options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 5.2.23 Inserting a character before the cursor . . . . . . . . . . . . . . . . . . . 81 5.2.24 Inserting a string before the cursor . . . . . . . . . . . . . . . . . . . . . . 81 5.2.25 Getting a string of characters from the screen. . . . . . . . . . . . 81 5.2.26 Low-level routines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 5.2.27 Mouse handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 5.2.28 Moving the cursor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 5.2.29 Output options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 5.2.30 Overlay and manipulate overlapped windows . . . . . . . . . . . . 86 5.2.31 Create and display pads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 5.2.32 Print data to a terminal-attached ports . . . . . . . . . . . . . . . . . . 88 5.2.33 Refresh windows and lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 5.2.34 Reading and writing a curses screen to a file . . . . . . . . . . . . . 89 5.2.35 Scroll a window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 5.2.36 Terminal attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 5.2.37 Refresh routines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 5.2.38 Miscellaneous utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 5.2.39 Window creation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 5.2.40 Terminal resizing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

5.3 The Function Key Label Library: (ncurses slk). . . . . . . . . . . . . . 94 5.4 The curses panel library: (ncurses panel) . . . . . . . . . . . . . . . . . . . . 95 5.5 The curses menu library: (ncurses menu) . . . . . . . . . . . . . . . . . . . . . 96

5.5.1 Menus Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 5.5.2 Menu Item Procedures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 5.5.3 Menu Procedures. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

5.5.3.1 Menu Colors and Attributes . . . . . . . . . . . . . . . . . . . . . . . . . 98 5.5.3.2 Positioning a menu’s cursor. . . . . . . . . . . . . . . . . . . . . . . . . . 99 5.5.3.3 The menu driver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 5.5.3.4 Menu sizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 5.5.3.5 Menu mark strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

5.5.4 Menu creation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 5.5.5 Menu options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

5.5.5.1 Menu pattern buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 5.5.5.2 Writing or erasing menus from window . . . . . . . . . . . . . 103 5.5.5.3 Control spacing between menu items. . . . . . . . . . . . . . . . 103 5.5.5.4 Associate menus with window . . . . . . . . . . . . . . . . . . . . . . 103

5.6 The curses form library: (ncurses form). . . . . . . . . . . . . . . . . . . . . 104 5.6.1 Forms Overview. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 5.6.2 Positioning a form window cursor . . . . . . . . . . . . . . . . . . . . . . . 104 5.6.3 Form data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 5.6.4 Command-processing loop of the form system . . . . . . . . . . . 104 5.6.5 Making or breaking connections between forms and fields

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 5.6.6 Color and attribute control for form fields . . . . . . . . . . . . . . . 107 5.6.7 Field buffer control. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 5.6.8 Retrieving field characteristics . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 5.6.9 Retrieving field characteristics . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 5.6.10 Creating and destroying form fields. . . . . . . . . . . . . . . . . . . . . 109 5.6.11 Setting and getting field options. . . . . . . . . . . . . . . . . . . . . . . . 109 5.6.12 Data type validation for fields . . . . . . . . . . . . . . . . . . . . . . . . . . 110 5.6.13 Creating forms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 5.6.14 Form pagination functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 5.6.15 Setting and getting form options . . . . . . . . . . . . . . . . . . . . . . . 111 5.6.16 Setting the page number and field . . . . . . . . . . . . . . . . . . . . . . 112 5.6.17 Writing or erasing forms from associated subwindows … 112 5.6.18 Printable form request names . . . . . . . . . . . . . . . . . . . . . . . . . . 112 5.6.19 Form windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

1 Introduction

1.1 Authors, Copyright, and Declarations

Michael Gran assembled this book, but, wrote only small sections of text. Most of reference text of the curses library was taken directly from the documentation of ncurses itself and was modified to make it specific to this Guile binding. The ncurses library is maintained by Tom Dickey, and the manual pages in the library note that the Free Software Foundation holds the copyright.

I’ve put his name on the title page, because it would seem rude not to do so; however, one should not get the impression that he approves of the Guile Ncurses project or that he is involved with the Guile NCurses project, or, indeed, that he is even aware of it.

The license for the original ncurses documentation is as follows.

Copyright (c) 1998-2005,2006 Free Software Foundation, Inc. Permission is hereby granted, free of charge, to any person obtaining a copy of this software and associated documentation files (the "Software"), to deal in the Software without restriction, including without limitation the rights to use, copy, modify, merge, publish, distribute, distribute with modifications, sublicense, and/or sell copies of the Software, and to permit persons to whom the Software is furnished to do so, subject to the following conditions: The above copyright notice and this permission notice shall be included in all copies or substantial portions of the Software. THE SOFTWARE IS PROVIDED "AS IS", WITHOUT WARRANTY OF ANY KIND, EXPRESS OR IMPLIED, INCLUDING BUT NOT LIMITED TO THE WARRANTIES OF MERCHANTABILITY, FITNESS FOR A PAR-TICULAR PURPOSE AND NONINFRINGEMENT. IN NO EVENT SHALL THE ABOVE COPYRIGHT HOLDERS BE LIABLE FOR ANY CLAIM, DAMAGES OR OTHER LIABILITY, WHETHER IN AN ACTION OF CON-TRACT, TORT OR OTHERWISE, ARISING FROM, OUT OF OR IN CON-NECTION WITH THE SOFTWARE OR THE USE OR OTHER DEALINGS IN THE SOFTWARE. Except as contained in this notice, the name(s) of the above copyright holders shall not be used in advertising or otherwise to promote the sale, use or other dealings in this Software without prior written authorization.

The tutorial for the curses library is adapted from the ncurses Programming HowTo by Pradeep Padala. It was modified by Michael Gran to make it specific to the Guile binding.

Again, I’ve put his name on the title page, because it would seem rude not to do so; however, one should not get the impression that he approves of the Guile Ncurses project or that he is involved with the Guile Ncurses project, or, again, that he is even aware of it.

That document holds the following license.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2001, 2002, 2005 Pradeep Padala Permission is hereby granted, free of charge, to any person obtaining a copy of this software and associated documentation files (the “Software”), to deal in the Software without restriction, including without limitation the rights to use, copy, modify, publish, distribute, distribute with modifications, sublicense, and/or sell copies of the Software, and to permit persons to whom the Software is furnished to do so, subject to the following conditions. The above copyright notice and this permission notice shall be included in all copies or substantial portions of the Software. THE SOFTWARE IS PROVIDED "AS IS", WITHOUT WARRANTY OF ANY KIND, EXPRESS OR IMPLIED, INCLUDING BUT NOT LIMITED TO THE WARRANTIES OF MERCHANTABILITY, FITNESS FOR A PAR-TICULAR PURPOSE AND NONINFRINGEMENT. IN NO EVENT SHALL THE ABOVE COPYRIGHT HOLDERS BE LIABLE FOR ANY CLAIM, DAMAGES OR OTHER LIABILITY, WHETHER IN AN ACTION OF CON-TRACT, TORT OR OTHERWISE, ARISING FROM, OUT OF OR IN CON-NECTION WITH THE SOFTWARE OR THE USE OR OTHER DEALINGS IN THE SOFTWARE. Except as contained in this notice, the name(s) of the above copyright holders shall not be used in advertising or otherwise to promote the sale, use or other dealings in this Software without prior written authorization.

2 Components of a Curses Interface

The goal of Ncurses is to simplify the task of programming for character-cell terminals. A character-cell terminal is a combination of a display that shows text in a monospaced font and the keyboard and mouse that connect to it. These days, there are three common types of terminals. Actual hardware terminals, such as the DEC VT220 or the Boundless Technologies VT525, are dedicated thin clients that interact with a networked computer via a serial port. These are still somewhat common in legacy applications or industrial applications where computers would be too expensive or too fragile, but, they are rapidly becoming extinct. Second are the consoles of text-mode operating systems, such as the console of GNU/Linux when used without X/Windows or the command prompt of FreeDOS. The third type is the terminal emulation programs that some windowing systems have. The best known terminal emulation program is probably XTerm, which does a good emulation of a VT220 terminal.

The Ncurses library attempts to create a standardized solution for these problems.

• A program needs to know which keys are pressed and when.

• A program needs to know a terminal’s capabilities. Can text be bold, italic, or in color?

• A program needs to know how to exploit a terminal’s capabilities. How does a program tell the terminal to move the cursor, to change text color, or to erase the screen?

2.1 Terminal Capabilities

The first step in any Ncurses program is to understand the terminal’s capabilities and attributes. Ncurses includes a database of terminals and their capabilities called terminfo or perhaps termcap. When a Ncurses program is initialized, the type of terminal is queried, usually by examining the environment variable TERM, and the capabilities are read from the database.

2.2 Keyboards and Keys

When a key is pressed, one or more bytes are sent to the program by the keyboard driver. These bytes need to be converted by the program back into a key name. For ASCII characters, this conversion is trivial. A key press of c usually sends the ASCII encoding 142 (c) which can be understood to mean the letter “c”. For other keys on the keypad, like F1 or HOME, the encoding can take multiple bytes and can differ from terminal to terminal. On some terminals, for example, pressing F1 is the same as if the user had typed ESC O P.

The terminfo database provides information that helps associate multibyte codes to keys on the keypad, and there are curses functions that will do the association.

2.3 Characters, Attributes, Colors, and Renditions

As the goal of curses is to allow a program to take advantage of the capabilities of terminal, each character on the screen has a color pair and a set of attributes.

In curses, the colors are usually defined in terms of color pairs, which is a combination of a foreground and a background color. Each character written to a terminal will have an associated color pair, assuming the terminal supports color.

Also each character has a set of attributes, such as bold, underline, or blink.

A character’s color pair and its attributes are collectively called its rendition.

2.4 Windows

Curses uses the concept of a window to visually organize the terminal. A window is a rectangular area on the terminal. A terminal always has one default window which is the size of the physical screen. This may be further subdivided into other windows. The default window is called the standard screen.

Each window has a set of associated visual properties: the default rendition of characters added to it and whether it scrolls when text reaches the bottom of the screen, for example.

Each window forms a coordinate system for the display of characters. The top-left character location in the window is the origin, (0, 0). x coordinates increase as one moves right and y coordinates increase as one moves down. Because the origin begins at zero, the largest allowable x value will be the number of columns minus one.

3 Types and encodings

Curses has several types specific to how it operates: rendered strings and windows, for example. Some of the curses types are represented by standard scheme types, while others are represented as opaque types (SMOBs).

3.1 Character types

The Guile Ncurses library uses two basic character types, simple characters and complex characters. Simple characters are the native Guile characters, and complex characters are those used to interact with the NCurses library. When using this library, a programmer will often have to convert simple characters to complex characters and vice versa, so it is important to understand their differences and applications.

Simple characters are the Guile native character type.

For older versions of Guile, such as the 1.6.x and the 1.8.x versions, characters were limited to being 8-bit. The lower 128 characters were ASCII, and the upper 128 characters had a meaning based on the locale of the system. If the locale were set to ISO-8859-1, for example, the upper 128 characters would be common accented letters used in western European languages. In almost all cases, each character took up the same space on a character cell terminal.

In newer versions of Guile, such as 2.0.x, the native Guile character type is a Unicode codepoint, a 32-bit number. Characters based on Unicode include all or most of the glyphs for the most languages. There are challenges when using the non-ASCII characters. There are double-width characters, such as hiragana and katakana, that take up two character spaces on a character cell terminal because of their complexity. There are combining characters, such as accents, that aren’t intended to stand alone and actually modify the previous character in a string.

In the terminology of Unicode, a 32-bit integer that maps to a wide character is a code point, and documents that refer to Unicode code points typically write them like this: U+XXXX, where XXXX is a four-digit hexadecimal number.

simple characters

An simple character is a standard Guile character, such as #\x. It is also referred to as an unrendered characters. It is unrendered because it has no associated color or attributes.

simple strings

A simple string, also called an unrendered string, is a standard Guile string, such as "hello", which is a sequence of unrendered characters.

complex characters

Rendered, complex characters usually are a standard character, zero to three combining characters, attribute information, and color information. The attribute information describes if the character is bold, dim, reverse video, etc. The guile ncurses library defines a special record type for complex characters: the xchar. Each complex character may contain more than one simple character. The first character in the list should be a base character or a control character. A base character is usually a character than can be printed standalone: combining accents and other letter codepoints intended to modify another letter are not base characters. The remaining characters in the list, if any, are accents or other combining characters that modify the appearance of the base character. If the first character was a control character, no combining characters are allowed.

Here are examples of the constructors used to make complex characters. These constructors will be described in more detail later on.

;; A constructor for a letter 'x' using the default colors
> (normal #\x)
;; The display format of the resulting complex character
==> #<xchar #\x>
;; A constructor for a bold letter 'L' using default colors
> (bold #\L)
;; The display format of the resulting bold 'L' character
==> #<xchar bold #\L>
;; A bold letter 'x' printed white on a green background
> (init-pair! 2 COLOR_WHITE COLOR_GREEN)
> (bold-on (color 2 #\x))
==> #<xchar bold color-pair #2 [white on green] #\x>
;; A letter 'n' overwritten with a tilde
> (define ntilde (normal #\n))
> (set-xchar-chars! ntilde '(#\n #\~))
> ntilde
==> #<xchar #\n #\~>
rendered, complex strings

Rendered, complex strings are lists of rendered, complex characters.

An example of the constructor for a rendered complex string, and the display format of that string

;; The constructor for a complex string: the word 'hello' in
;; reverse video
> (inverse "hello")
;; The display format of the resulting string could be...
==> (#<xchar reverse color-pair #0 [white on black] #\h>
     #<xchar reverse color-pair #0 [white on black] #\e>
     #<xchar reverse color-pair #0 [white on black] #\l>
     #<xchar reverse color-pair #0 [white on black] #\l>
     #<xchar reverse color-pair #0 [white on black] #\o>)

3.2 Opaque types

These are types from system libraries that are complex or system-dependent. They are implemented as SMOBs.

The #<window> type

The #<window> type is an opaque type that contains information about the characters and renditions for a given window. It is created by the function newwin and is used and most curses functions.

The #<screen> type

The #<screen> type is an opaque types that contain information about the physical terminal. Variables of type #<screen> are created with the function newterm and used by the function set-term. This type is usually only found in programs that use multiple terminals.

The #<panel> type

The #<panel> type is an opaque type that represents a window’s place in a stack of windows. It is created with new-panel. It is used in functions that deal with overlapping windows

The #<item> type

The #<item> type is an opaque type that represents one menu item in a menu. It is created with new-item. It is part of the menu library.

The #<menu> type

The #<menu> type is an opaque type that represents a menu. It is created with new-menu from a list of #<item> types.

The #<field> type

The #<field> type is an opaque type that represents one text box or entry on a form.

The #<form> type

This represents a form, created using new-form from a list of elements of type

#<field>.

4 Curses Tutorial

4.1 Hello World!!!

Welcome to the world of curses. Before we plunge into the library and look into its various features, let’s write a simple program and say hello to the world.

4.1.1 Using the Guile Curses module

To use ncurses library functions, you have to load the (ncurses curses) module into the program.

Here is the Hello World program.

     #!/usr/bin/guile
     !#
     (use-modules (ncurses curses))
     (define stdscr (initscr))
     (addstr stdscr "Hello World!!!")
     (refresh stdscr)
     (getch stdscr)
     (endwin)

The first line of the example, #!/usr/bin/guile, gives the location of where Guile is installed on my system. This may differ on your system.

The above program prints “Hello World!!!” to the screen, waits for the user to press any key, and then exits. This program shows how to initialize curses and do screen manipulation and end curses mode. Let’s dissect it line by line.

4.1.1.1 About initscr

The procedure initscr initializes the terminal in curses mode. It clears the screen and presents a blank screen. To do any screen manipulation using the curses package, this has to be called first. This function initializes the curses system and allocates memory for screen handling and some other data structures. It returns a SMOB that represents the default window: the window that represents the entire screen. By convention, this window is denoted stdscr, the standard screen. Under extreme cases, this function might fail due to insufficient memory to allocate memory for curses library’s data structures.

The procedure initscr returns a #<window> that contains necessary information about the curses screen. The #<window> that is returned must be stored until the program is finished with the curses library. If that #<window> is garbage collected, the curses library cannot continue.

After this is done, we can do a variety of initializations to customize our curses settings

4.1.1.2 The mysterious refresh

The next line addstr prints the string “Hello World!!!” on to the screen. This function prints the data on a window called stdscr at the current (y, x) coordinates. Since our present coordinates are at 0,0 the string is printed at the top, left-hand corner of the window.

This brings us to the mysterious (refresh stdscr). Well, when we called addstr, the data is actually written to an imaginary window, which is not updated on the screen yet. The job of addstr is to update a few flags and data structures and write the data to a buffer corresponding to stdscr. In order to show it on the screen, we need to call refresh and tell the curses system to dump the contents on the screen.

The philosophy behind all this is to allow the programmer to do multiple updates on the imaginary screen or windows and to do a refresh once all the screen update is done. refresh checks the window and updates only the portion which has been changed. This improves performance and offers greater flexibility, too. But, it is sometimes frustrating to beginners. A common mistake committed by beginners is to forget to call refresh after they did some update through addstr.

4.1.1.3 About endwin

And finally, don’t forget to end the curses mode. Otherwise your terminal might behave strangely after the program quits. endwin frees the memory taken by the curses sub-system and its data structures and puts the terminal in normal mode. This function must be called after you are done with curses mode.

4.2 Interactive sessions and guile-ncurses-shell

If you would like to try out these functions interactively by typing them into the scheme shell, instead of typing them and running them as scripts, the program guile-ncurses-shell can be used. The problem with interactive Guile sessions using curses is that you are typing into the same screen that curses is trying to manage, which leads to confusing results. The program guile-ncurses-shell, which must be run on X, starts an interactive guile session and creates and xterm that will be managed by curses. The results of the curses function calls will appear in the xterm, instead of in the screen where the interactive guile session occurs.

Here we’ll try four functions: initscr sets up the screen, addstr writes a string to the screen, refresh redraws the screen, and endwin frees the screen.

Upon initialization guile-ncurses-shell automatically calls these functions.

     (use-modules (ncurses curses))
     (define %guile-ncurses-shell-stdscr (initscr))

It loads the ncurses module; initializes the screen; and saves the returned ncurses screen structure in the variable %guile-ncurses-shell-stdscr. Thus, you do not need to call (initscr) in your interactive session when you use guile-ncurses-shell.

The first thing you want to do is to redefine the name of the standard window to something more reasonable.

     (define win %guile-ncurses-shell-stdscr)

To check and see if the guile-ncurses-shell is working, you can write a string onto the created xterm.

     (addstr win "hello, world!")
     (refresh win)

The guile-ncurses-shell communicates with the xterm using a read port and a write port. If, for some obscure reason, you need to access the read or write port directly, they are stored

in the variables %guile-ncurses-shell-read-port and %guile-ncurses-shell-write-
port.

When exiting an interactive session, by C-D for example, guile-ncurses-shell automatically calls (endwin), so you do not need to call that yourself.

4.3 The Gory Details

Now that we’ve seen how to write a simple curses program, let’s get into the details. There are many functions that help customize what you see on the screen and many features which can be put to full use.

Here we go…

4.4 Initialization

We now know that to initialize curses the function initscr has to be called. There are functions which can be called after this initialization to customize our curses session. We may ask the curses system to set the terminal in raw mode or initialize color or initialize the mouse, etc. Let’s discuss some of the functions that are normally called immediately

after initscr.

4.4.1 Initialization Functions

4.4.1.1 raw! and cbreak!

Normally the terminal driver buffers the characters a user types until a newline or carriage return is encountered. But most programs require that the characters be available as soon as the user types them. The functions cbreak! and raw! are used to disable line buffering. The difference between these two functions is in the way control characters like suspend (Ctrl-Z), interrupt and quit (Ctrl-C) are passed to the program. In the raw! mode, these characters are directly passed to the program without generating a signal. In the cbreak! mode, these control characters are interpreted as any other character by the terminal driver, allowing Ctrl-C and Ctrl-Z to quit and suspend the program.

4.4.1.2 echo! and noecho!

These functions control the echoing of characters typed by the user to the terminal. noecho! switches off echoing. With echoing off, when a user presses a character, it is not displayed on the screen.

4.4.1.3 keypad!

This function enables the reading of function keys like F1, F2, arrow keys, etc. Almost every interactive program enables this, as arrow keys are a major part of any user interface. Do (keypad! stdscr #t) to enable this feature for the regular screen stdscr (assuming that stdscr is the variable you used to hold the output of initscr.)

4.4.1.4 halfdelay!

This function is useful when you want to ask the user for input, and if he doesn’t respond within a certain time, do something else. One possible example would be a timeout at a password prompt. halfdelay! enables half-delay mode, which is similar to cbreak! mode in that characters types are immediately available to the program. However, after a period of time if there is no input, it returns #f.

4.4.2 An example

Let’s write a program that will clarify the usage of these functions.

     #!/usr/bin/guile
     !#
     (use-modules (ncurses curses))
     (define stdscr (initscr))          ; Start curses mode
     (raw!)                                                          ; Line buffering disabled
     (keypad!  stdscr #t)                ; We get F1, F2, etc
     (noecho!)                                                    ; Don't echo when we get a keypress
     (addstr stdscr "Type any character to see it in bold\n")
     (let ((ch (getch stdscr)))         ; Read a key press, put it in 'ch'
       (addstr stdscr "The pressed key is ")
       (if (char?  ch)                    ; If a non-function key is pressed
           (addch stdscr (bold ch))      ; print its name
           (addchstr stdscr (bold (keyname ch)))) ; Or, print the function
                                                  ; key name

       (refresh stdscr)                      ; Print it on the real screen
       (getch stdscr)                        ; Wait for user input

       (endwin))                             ; End curses mode

Hopefully this program is easy to follow even though I used functions that aren’t explained yet. The procedure getch is used to get a character from the user. If it returns a character, the user pressed a character key. If it returns a number, the use pressed a function key. The (bold ch) adds the bold attribute to the character, and addch prints the character on the screen.

4.5 A word about windows

Before we plunge into the myriad ncurses functions, let me clear a few things about windows. Windows are explained in detail in the following sections.

A window is an imaginary screen defined by the curses system. A window does not mean a bordered window which you usually see in GNOME or KDE system. When curses is initialized, it creates a default window conventionally named stdscr which represents your 80x25 (or the size of window in which you are running) screen. If you are doing simple tasks like printing a few strings, reading input, etc., you can safely use this single window for all of your purposes. You can also create windows and call functions which explicitly work on the specified window.

For example, if you call

     (addstr stdscr "Hi There!!!")
     (refresh stdscr)

It prints the string on stdscr at the present cursor position. Similarly, the call to

refresh works on stdscr only.

Say you have created multiple windows, then you have to call the functions separately on each window. When you call refresh, you need to call it on the window that was updated.

     (addstr win "Hi There!!!")
     (refresh win)

For many functions, there are optional key parameters.

     (addstr stdscr string)             ; Print on stdscr at present
                                        ; cursor location
     (addstr stdscr string #:y y #:x x) ; Move to (y, x) then print string

4.6 Output functions

I guess you can’t wait any more to see some action. Back to our odyssey of curses functions. Now that curses is initialized, let’s interact with the world.

There are three primary functions which you can use to output characters and strings to the screen.

1. addch: output rendered characters

2. addstr: output unrendered strings

3. addchstr: output rendered strings

4.6.1 The addch procedure

The addch function puts a single rendered character into the current cursor location and advances the position of the cursor. Attributes are explained in detail in later sections of the document. If a character is associated with an attribute (bold, reverse video, etc.), when curses prints the character, it is printed with that rendition.

In order to combine a character with some attributes, you have two options:

1. You can explicitly construct a rendered character by passing a simple character through

the rendition functions for the desired attributes:

         (bold #\x)
         (blink-on (bold-on #\x))
         (color 2 #\x)

2. By using functions like attr-set!, attr-on!, attr-off!, you can manipulate the current attributes of the given window. Once set, the characters printed in the window are associated with the attributes until it is turned off.

Additionally, curses provides some special characters for character-based graphics. You can draw tables, horizontal or vertical lines. Try looking for the procedures beginning with

acs-.

4.6.1.1 Moving the cursor addch

The optional key parameters #:y y #:x x can be used to move the cursor to a given point, and then print. Thus the calls

     (move stdscr row col)
     (addch stdscr ch)

can be replaced by

     (addch stdscr ch #:y row #:x col)

4.6.1.2 Complex characters and addch

There are really two version of the NCurses library: a standard version and a wide version. When guile ncurses was compiled, it was associated with either the standard version libncurses or the wide version libncursesw. The wide version has greater capability to print non-Latin characters than the standard version.

For every C function that operates on characters, there is a parallel function that operates on wide characters. The guile ncurses library hides all of that complexity, and presents the same API regardless of whether it used libncurses or libncursew.

At this point, a C programmer familiar with ncurses might be wondering how to call add-wch to print, for example, a Chinese character. The guile ncurses library abstracts both the C ncurses function addch and the C ncurses function add-wch as the Guile function

addch.

So, if you version of Guile is capable of Unicode characters (such as Guile version 2.0.x), and if you version of NCurses is the wide version libncursesw, then you can use this library to print non-Latin characters.

First off, if you want to use wide characters, you need to call (setlocale LC_ALL "") before the call to initscr. The locale that is set must be an encoding that has greater than 8-bit characters, such as UTF-8. Also, you terminal must be capable of printing non-Latin characters.

Then, to add a rendered, complex character to the screen, use addch and friends as before

     ;; Bold U+041B Cyrillic Capital Letter El
     (addch stdscr (bold #\))

4.6.2 addstr class of functions

These functions output entire strings to the screen.

They can be used effectively in conjunction with the format procedure from (ice-9

format).

4.6.3 A simple addstr example

     #!/usr/bin/guile
     !#
     (use-modules (ncurses curses)
                  (srfi srfi-1))
     (define stdscr (initscr))
     (let* ((mesg "Just a string")
            (len (string-length mesg))
            (siz (getmaxyx stdscr))
            (row (first siz))

            (col (second siz)))
       ;; Print the message centered in the window
       (move stdscr
             (round (/ row 2))
             (round (/ (- col len) 2)))
       (addstr stdscr mesg)
       ;; Use "format" to generate a message, and then print it
       (addstr stdscr
               (format #f "This screen has ~a rows and ~a columns ~%"
                       row col)
               #:y (- row 2)
               #:x 0)
       (addstr stdscr "Try resizing your window (if possible) ")
       (addstr stdscr "and then run this program again")
       (refresh stdscr)
       ;; Wait for a keypress
       (getch stdscr)
       (endwin))

The above program demonstrates how easy it is to combine addstr and move to print at a specific location on the screen. It also shows how to use the equivalent key parameters #:y and #:x. They do exactly the same thing. It also shows how (format #f ...) can be used in conjunction with addstr to do formatted output.

The example introduces the new function getmaxyx. It gives the number of columns and the number of rows in a given window. getmaxyx does this by returning a list of two elements, y and x.

4.6.4 A word about coordinates

The move function takes the y coordinate first and then x as its arguments. A common mistake by beginners is to pass x, y in that order.

If you use the #:y and #:x key parameters that most output procedures have in lieu of using the move statement, you can use enter the coordinates in whichever order seems most natural to you.

     ;; either
     (addstr win "Hi" #:x 0 #:y 10)
     ;; or
     (addstr win "Hi" #:y 10 #:x 0)

4.7 Input functions

Well, printing without taking input is boring. Let’s see functions which allow us to get input from the user.

1. getch: get a character 2. getnstr: get a string

4.7.1 getch

The procedure getch reads a single character from the terminal. But there are several subtle facts to consider. For example, if you don’t use the function cbreak!, curses will not read your input characters contiguously, but, will begin to read them only after RET has been pressed or after and EOF has been encountered. In order to avoid this, the cbreak! function must be used so that characters are immediately available to your program. Another widely used function is noecho!. As the name suggests, when this function is set (used), the characters that are keyed in by the user will not show up on the screen. The two functions cbreak! and noecho! are typical examples of key management.

4.7.2 getnstr

The getnstr procedure is used to get strings from the terminal. In essence, it performs the same task as would be achieved by a series of calls to getch until a NL, CR, or EOF is received. The resulting string of characters is returned. The procedure always takes a length parameter which is the maximum length of string that it will allow to be input.

4.7.3 An example of getnstr and getch

     #!/usr/bin/guile
     !#
     (use-modules (ncurses curses)
                  (ice-9 format))
     (define stdscr (initscr))
     (let* ((mesg "Enter a string:  ")
            (len (string-length mesg))
            (siz (getmaxyx stdscr))
            (row (car siz))
            (col (cadr siz)))
       (addstr stdscr mesg
               #:y (round (/ row 2))
               #:x 0)
       (refresh stdscr)
       (let ((str (getnstr stdscr 80)))
         (addstr stdscr
                 (format #f "You entered:  ~s~%" str)
                 #:y (- row 2)
                 #:x 0)
         (getch stdscr)))
     (endwin)

4.8 Attributes

Attributes can be used to print characters with some special effects. Attributes, when set prudently, can present information in an easy, understandable manner.

The following program takes a scheme file as input and prints the file with comments in bold. It does so by turning on the A_BOLD attribute using attr-on! when a semicolon is detected, and then turning of the A_BOLD attribute using attr-off! when a newline is detected. Have a look.

#!/usr/bin/guile
!#
(use-modules (ncurses curses)
             (ice-9 format))
;; A helper function that return the cursor's current row
(define (getrow win)
  (car (getyx win)))
;; The program should be passed a filename from the command line
(if (not (eqv?  2 (length (command-line))))
    (begin
      (format #t "Usage:  ~a <scm file name>~%" (car (command-line)))
      (primitive-exit 1)))
(let* ((filename (cadr (command-line)))
       (fport (open-input-file filename))
       (stdscr (initscr)))
  ;; Read one char at a time from the file
  (let loop ((ch (read-char fport)))
    (if (not (eof-object?  ch))
        (begin
          ;; Wait for a key press once a page
          ;; of text has been printed
          (if (eqv?  (getrow stdscr) (- (lines) 1))
              (begin
                (addstr stdscr "<-Press any key->")
                (refresh stdscr)
                (getch stdscr)
                (clear stdscr)
                (move stdscr 0 0)))
          ;; Bold all text between a semicolon
          ;; and the end of a line
          (cond
           ((eqv?  ch #\;)
            (attr-on!  stdscr A_BOLD))
           ((eqv?  ch #\nl)

                 (attr-off!  stdscr A_BOLD)))
               (addch stdscr (normal ch))
               (refresh stdscr)
               (loop (read-char fport)))
             ;; Clean up and exit
             (begin
               (addstr stdscr "<-Press any key->")
               (refresh stdscr)
               (getch stdscr)
               (endwin)
               (close-input-port fport)))))

One important thing to note is that in this program, addch is always passed a normal, un-bolded, character. Note the line

               (addch stdscr (normal ch))

But yet, the character printed by addch may still appear as bold on the screen. This is because the character attributes passed to addch combine with the character attributes set by attr-on!. If attr-on! has set the window’s default attributes to bold, that will merge with the attributes passed to addch.

The function also introduces the useful function getyx. It returns the coordinates of the present cursor as a list of two elements.

The above program is really a simple one which doesn’t do much. Along these lines once could write a more useful program which reads a scheme file, parses it, and prints it in different colors.

4.8.1 The details

Let’s get into more details of attributes. The functions attr-on!, attr-off!, attr-set!, and their sister functions attr-get etc., can be used to switch attributes on and off, get attributes, and produce a colorful display.

The functions attr-on! and attr-off! take a bit-mask of attributes and switch them on or off respectively. The following video attributes can be passed to these functions.

A_NORMAL Normal display (no highlight)

A_STANDOUT

Best highlighting mode of the terminal

A_UNDERLINE

Underlining

A_REVERSE

Reverse video

A_BLINK Blinking

A_DIM Half-bright

A_BOLD Extra bright or bold

A_PROTECT

Protected mode

A_INVIS Invisible or blank mode

A_ALTCHARSET

Alternate character set

A_CHARTEXT

Bit-mask to extract a character

A_HORIZONTAL

Unsupported for now and left for future use

A_LEFT Unsupported for now and left for future use

A_LOW Unsupported for now and left for future use

A_RIGHT Unsupported for now and left for future use

A_TOP Unsupported for now and left for future use

A_VERTICAL

Unsupported for now and left for future use

(color-pair n)

A procedure that returns the bit mask for a color-pair

Colors are explained in the next sections.

We can logior any number of attributes to get a combined effect. If you wanted the character in window win to have reverse video with blinking characters, you can use

     (attr-on! win (logior A_REVERSE A_BLINK))
4.8.2 attr-on! vs attr-set!

Then what is the difference between attr-on! and attr-set!? attr-set! sets the attributes of a window whereas attr-on! just switches on the attribute given to it. So attr-set! fully overrides whatever attributes the window previously had and sets it to the new attribute(s). Similarly, attr-off! just switches off the attributes(s) given to it as an argument. This gives us the flexibility of managing attributes easily. But, if you use them carelessly, you may lose track of what attributes the window has and garble the display. This is especially true while managing menus with colors and highlighting. So decide on a consistent policy and stick to it. You can always use standend! which is equivalent to (attr-set! win A_NORMAL) which turns off all attributes and brings you back to normal mode.

4.8.3 attr-get

The function attr-get gets the current attributes and color pair of the window. Though we might not use this as often as the above functions, this is useful in scanning areas of a screen. Say we wanted to do some complex update on the screen and we are not sure what attribute each character is associated with. Then this function can be used with either attr-set! or attr-on! to produce the desired effect.

4.8.4 chgat function

The function chgat, which is short for change attributes, can be used to set attributes for a group of characters already on the screen without moving the cursor. It changes the attributes of a given number of characters starting at the current cursor location.

You can pass it -1 as the character count to update until the end of the current line. The following example changes the attributes of characters from the current position to

the end of the line to reverse video on a window named win1.

(chgat win1 -1 A_REVERSE 0)

The are optional key parameters #:y and #:x can be used with chgat. The following example will print a string on the screen. Then it will set the first 5

characters of the string to blink and change color to cyan.

#!/usr/bin/guile
!#
(use-modules (ncurses curses))
(define stdscr (initscr))
;; Prep the color functions
(start-color!)
;; Label cyan on black as color-pair #1
(init-pair!  1 COLOR_CYAN COLOR_BLACK)
(addstr stdscr "Blink Don't Blink")
(chgat   stdscr              ; window
         5                   ; num of chars
         A_BLINK             ; attributes
         1                   ; use color pair #1
         #:y 0               ; start y
         #:x 0)              ; start x
;; Move the cursor out of the way
(move stdscr 1 0)
(refresh stdscr)
(getch stdscr)
(endwin)

This example also introduces us to the color world of curses. Colors will be explained in detail later. Use 0 for white on black.

Now wait… Did you try running this little script? Did it work? Blinking is one of those features that may not be implemented on your terminal. As of the moment of this writing, Gnome terminal doesn’t do blinking. The standard xterm does do blinking, but, it doesn’t blink at the location of the cursor.

4.9 Windows Tutorial

Windows form the most important concept in curses. You have seen the standard window stdscr used in most of the previous examples. Now to design even a simplest GUI, you need to resort to windows. The main reason you may want to use windows is to manipulate parts of the screen separately, for better efficiency, by updating only the windows that need to be changed, and for a better design. I would say the last reason is the most important in going for windows. You should always strive for a better and easy-to-manage design in your programs. If you are writing big, complex GUIs, this is of pivotal importance before you start doing anything.

Now, anyone that has worked with a modern GUI has an intuitive understanding of what a “window” is. You will need to unlearn this knowledge when programming curses windows.

First off, a curses window doesn’t necessarily have any sort of border or decoration that separates it from the rest of the terminal. You can make a curses window and then draw a border on it, but, that is up to the programmer.

Second, curses windows don’t normally overlap. They are usually distinct, non-overlapping regions of the screen.

4.9.1 The basics

A window can be created by calling the function newwin. It doesn’t draw anything on the screen initially. It allocates memory for a structure to manipulate the window and updates the structure with data regarding the window, like it’s size, position, etc.

The function newwin, like initscr, returns a #<window>. Just like the stdscr created by initscr, the window created by newwin can be used in any procedure that takes a window argument. The window returned by newwin needs to remain in scope for as long as the window is to be used. If it is garbage collected, the memory associated with the window will eventually be freed.

The memory associated with a window can be more quickly freed by explicitly calling delwin. It will deallocate the memory associated with the window structure. Any further attempts to then use that window will cause an error.

4.9.2 Let there be a window

What fun is it if a window is created and we can’t see it? So the fun part begins by displaying the window. The functions box and border can be used to draw a border around the window. Let’s explore these functions in more detail in this example.

This example isn’t meant as a practical example of what windows are good for. It just shows how to make, draw, erase, and destroy them.

     #!/usr/bin/guile
     !#
     (use-modules (ncurses curses))
     ;; This procedure makes a new window and draws a box
     ;; around it
     (define (create-newwin height width starty startx)

  ((lambda (win)                        ; Make a lambda proc that
     (box win (acs-vline) (acs-hline))  ; Makes a box,
     (refresh win)                      ; Draws the window
     win)                               ; Returns the window to the caller

   (newwin height width starty startx))) ; Create a window and apply it
                                        ; to the lambda function
;; This procedure erases the box around a window and then deletes it
(define (destroy-win win)
  (let ((s (normal #\sp)))

    (border win s s s s s s s s)        ; Draw a box of spaces

    (refresh win)
    (delwin win)))
;; This prodecure deletes a window than then draw a new one someplace
;; else
(define (move-win win height width starty startx)
  (destroy-win win)
  (create-newwin height width starty startx))

;; Program Begins

(define stdscr (initscr))               ; Start curses
(cbreak!)                               ; Line buffering disabled
(keypad!  stdscr #t)                    ; Check for function keys

(let* ((height 3)
       (width 10)
       (starty (round (/ (- (lines) height) 2)))
       (startx (round (/ (- (cols) width) 2))))
  (addstr stdscr "Press F1 to exit")
  (refresh stdscr)
  (let loop ((starty starty)
             (startx startx)
             (my-win (create-newwin height width starty startx))
             (ch (getch stdscr)))
    (cond
     ((eqv?  ch KEY_LEFT)
      (loop starty
            (- startx 1)
            (move-win my-win height width starty (- startx 1))
            (getch stdscr)))
     ((eqv?  ch KEY_RIGHT)

           (loop starty
                 (+ startx 1)
                 (move-win my-win height width starty (+ startx 1))
                 (getch stdscr)))
          ((eqv?  ch KEY_UP)
           (loop (- starty 1)
                 startx
                 (move-win my-win height width (- starty 1) startx)
                 (getch stdscr)))
          ((eqv?  ch KEY_DOWN)
           (loop (+ starty 1)
                 startx
                 (move-win my-win height width (+ starty 1) startx)
                 (getch stdscr)))
          ((eqv?  ch (key-f 1))
           #f)
          (else
           (loop starty startx my-win (getch stdscr)))))
       (endwin))

Don’t scream. I know it is a big example. But there are some important things to explain here. This program creates a rectangular window that can be moved with left, right, up, and down arrow keys. It repeatedly creates and destroys windows as a user presses a key. Don’t go beyond the screen limits. Checking for limits is left as an exercise for the reader. Let’s dissect it line by line.

The create-newwin function creates a window with newwin and draws a box around it with box. For the horizontal lines in the box, I chose the special drawing character acshline. The vertical lines are the special drawing character acs-vline. For the corners of the box, the box procedure will use a guess of the best available corners for the terminal.

Most terminals will have special box drawing characters available. The procedure acshline and acs-vline will return these special drawing characters. If the terminal you are using does not have box drawing characters available, acs-hline and acs-vline will return the hyphen “-” and the vertical bar “|”.

The procedure destroy-win first erases the window from the screen by painting a border of blanks and then calling delwin to deallocate memory related to it. Depending on the key the user presses, startx and starty are changed, and a new window is created.

In the destroy-win, as you can see, I used border instead of box. The reason is this: border draws a border around the window and the characters given to it as the four corners and the four lines. To put it clearly, if you called border as below:

     (border win
             (normal #\|)

             (normal #\|)
             (normal #\-)
             (normal #\-)
             (normal #\+)
             (normal #\+)
             (normal #\+)
             (normal #\+))

it produces something like this

   +-----+
   |     |
   |     |
   |     |
   +-----+

It wouldn’t have been sufficient to use (box win (normal #\sp) (normal #\sp)) to erase the box, because the box procedure still would have drawn the four corners of the box.

4.9.3 Other functions

You can also see in the above examples that I have used the procedures cols, and lines. These procedures return the size of the stdscr.

The function getch as usual gets the key from the keyboard, and then the cond expression acts on the key pressed. This type of cond expression is very common in any GUI-based program.

4.9.4 Other border functions

The box and border procedures are useful, but, they only draw borders around the outside of windows. If you want to draw lines on the screen is other locations than the border of windows, you can use the hline and vline procedures.

The following little program shows how to draw a box at any location. To draw a box, it needs to draw four corners, two horizontal lines, and two vertical lines. It uses hline and vline. These two functions are simple. They create a horizontal or vertical line of the specified length at the specified position. The program uses more of the special drawing characters like (acs-urcorner), which is the upper-right corner of a box.

     #!/usr/bin/guile
     !#
     (use-modules (ncurses curses))
     (define stdscr (initscr))
     ;; Draw a box the hard way
     (define (box2 win y x height width)
       ;; top
       (move win y x)
       (addch win (acs-ulcorner))

(move win y (1+ x))
(hline win (acs-hline) (- width 2))
(move win y (+ x width -1))
(addch win (acs-urcorner))
;; left side
(move win (+ y 1) x)
(vline win (acs-vline) (- height 2))
;; right side
(move win (+ y 1) (+ x width -1))
(vline win (acs-vline) (- height 2))
;; bottom
(move win (+ y height -1) x)
(addch win (acs-llcorner))
(move win (+ y height -1) (1+ x))
(hline win (acs-hline) (- width 2))
(move win (+ y height -1) (+ x width -1))
(addch win (acs-lrcorner)))
(let* ((stdscr (initscr))
     (height 3)
     (width 10)
     (y (round (/ (- (lines) height) 2)))
     (x (round (/ (- (cols) width) 2))))
(box2 stdscr y x height width)
(refresh stdscr)
(sleep 3)
(endwin))

4.10 Colors

4.10.1 The basics of color

Life seems dull with no colors. Curses has a nice mechanism to handle colors. Let’s get into the thick of things with a small program.

#!/usr/bin/guile
!#
(use-modules (ncurses curses))
(define stdscr (initscr))
(if (has-colors?)
  (begin
    (start-color!)
    (init-pair!  1 COLOR_GREEN COLOR_YELLOW)
    (attr-on!  stdscr (logior A_BOLD (color-pair 1)))

           (addstr stdscr "Voila!!  In color...")
           (refresh stdscr)
           (sleep 3)
           (endwin)
           0)
         (begin
           (endwin)
           (display "Your terminal does not support color")
           1))

As you can see, to start using color, you should first call start-color!. After that you can use color capabilities of your terminal. To find out whether a terminal has color capabilities or not, you can use has-colors?, which returns #f if the terminal does not support color.

Curses initializes all the color support for the terminal when start-color! is called. Usually, a color terminal will have at least eight colors available that can be accessed by constants like COLOR_BLACK, etc. Now to actually start using colors, you have to define pairs. Colors are always used in pairs. That means you have to use the function initpair! to define the foreground and background for the pair number you give. After that, that pair number can be used as a normal attribute with the color-pair procedure. This may seem to be cumbersome at first.

The following colors are defined. You can use these as parameters for the various color functions.

1. COLOR BLACK

2. COLOR RED

3. COLOR GREEN

4. COLOR YELLOW

5. COLOR BLUE

6. COLOR MAGENTA

7. COLOR CYAN

8. COLOR WHITE

4.10.2 Changing color definitions

Most terminals don’t support defining new colors or changing the colors that already exist. For those terminals that do, the function init-color! can be used to change the definition of colors by adjusting the amount of red, green, and blue in each. To set the red to a darker red, you could use this function

     ;; Param 1: color name
     ;; Param 2, 3, 4; RGB content, min=0 max=1000
     (init-color! COLOR_RED 300 0 0)

If your terminal cannot change the color definitions, the procedure will throw and error. The function can-change-color? can be used to find out whether that terminal has the capability of changing color content or not. The RGB content is scaled from 0 to 1000. Initially, red is likely defined with content 680 (r), 0 (g), 0 (b).

The functions color-content and pair-content can be used to find the color content of a given color and foreground/background combination of a color pair.

4.11 Interfacing with the keyboard

4.11.1 The basics of keys

No GUI is complete without a strong user interface and to interact with the user, a curses program should be sensitive to key presses or mouse actions done by the user. Let’s deal with keys first.

As you have seen in almost all of the above examples, it’s very easy to get key input from the user. A simple way of getting key presses is the getch procedure. The cbreak! mode should be enabled to read keys when you are interested in reading individual key hits rather than complete lines of text (which usually end with a CR). keypad! should be enabled to get the function keys, arrow keys and so on.

getch returns a character if the key pressed maps to a standard character or it returns an integer code for those keys that don’t map to characters, such as PAGE UP or DELETE. This integer code which can be matched with the KEY_constants. For example, if the user presses F1, the integer returned is 265. The procedure (key-f 1) returns 265.

For example, if you call getch like this

     (let ((x (getch win)))
     ...

getch will wait for the user to press a key, (unless you specified a timeout), and when the user presses a key, the corresponding character or integer is returned. If it is an integer, then you can check the value returned with the KEY_constants or the result of the function key procedure key-f.

The following code piece will do that job

     (let ((ch (getch win)))
         (if (eqv? ch KEY_LEFT)
             (addstr win "Left arrow is pressed")))

Let’s write a small program which creates a menu that can be navigated by up and down arrows.

4.11.2 A simple key usage example

     #!/usr/bin/guile
     !#
     (use-modules (ncurses curses))
     ;; Draw the menu on to the window WIN, using the list of CHOICES,
     ;; highlighting the currently selected entry.
     (define (print-menu win highlight choices)
       (let ((x 2)
             (y 2)
             (n-choices (length choices)))

;; The menu border
(box win (acs-vline) (acs-hline))
;; The menu entries from CHOICES, with the current entry
;; highlighted
(let loop ((i 0)
       (y y)
       (n n-choices))
(cond
((eqv?  highlight i)
(attr-on!  win A_REVERSE)
(move win y x)
(addstr win (list-ref choices i))
(attr-off!  win A_REVERSE))
(else
(move win y x)
(addstr win (list-ref choices i))))
(if (< i (- n-choices 1))
  (loop (+ 1 i)
        (+ 1 y)
        n-choices)
  (refresh win)))))

;; This loop lets the user select a menu.  It returns the number of the
;; selected item.
(define (get-selection menu-win highlight choices)
(let ((n-choices (length choices)))
(let loop ((highlight highlight)
       (ch (getch menu-win)))
(cond
((eqv?  ch KEY_UP)
((lambda (h)
   (print-menu menu-win h choices)
   (loop h (getch menu-win)))
 (if (eqv?  highlight 0)
     (- n-choices 1)
     (- highlight 1))))
((eqv?  ch KEY_DOWN)
((lambda (h)
   (print-menu menu-win h choices)

(loop h (getch menu-win)))
(if (eqv?  highlight (- n-choices 1))
  0
  (+ highlight 1))))
;; If enter or return is pressed, return the current selected
;; menu item
((or (eqv?  ch #\nl) (eqv?  ch KEY_ENTER))
highlight)
(else
(loop highlight (getch menu-win)))))))

(define (main)
(let ((stdscr(initscr)))
(cbreak!)
(noecho!)
(let* ((menu-width 30)
(menu-height 10)
(startx (round (/ (- (cols) menu-width) 2)))
(starty (round (/ (- (lines) menu-height) 2)))
(menu-win (newwin menu-height menu-width starty startx))
(choices '("Choice 1"
           "Choice 2"
           "Choice 3"
           "Choice 4"
           "Exit"))
(highlight 0))
(begin
(keypad!  menu-win #t)
(print-menu menu-win 0 choices)
(let ((choice (get-selection menu-win highlight choices)))
(move stdscr 23 0)
(addstr stdscr (format #f "You chose ~s ~%"
             (list-ref choices choice)))
(clrtoeol stdscr)
(refresh stdscr)
(sleep 2)
(endwin)
0)))))
(main)

4.12 Interfacing with the mouse

Now that you have seen how to get keys, let’s do the same thing from the mouse. Usually each UI allows the user to interact with both keyboard and mouse.

4.12.1 The basics of the mouse

Before you do anything else, the events you want to receive have to be enabled with mousemask. You pass it a bit mask of events you would like to listen. By default, all the events are turned off. The bit mask ALL_MOUSE_EVENTS can be used to get all the events.

For current technology, the following events are of use.

Name  Description
BUTTON1_PRESSED  mouse button 1 down

BUTTON1_RELEASED  mouse button 1 up

BUTTON1_CLICKED  mouse button 1 clicked
BUTTON1_DOUBLE_CLICKED  mouse button 1 double clicked
BUTTON2_PRESSED  mouse button 2 down

BUTTON2_RELEASED  mouse button 2 up

BUTTON2_CLICKED  mouse button 2 clicked
BUTTON2_DOUBLE_CLICKED  mouse button 2 double clicked
BUTTON_SHIFT  shift was down during button state change
BUTTON_CTRL  control was down during button state change
BUTTON_ALT  alt was down during button state change
ALL_MOUSE_EVENTS  report all button state changes

REPORT_MOUSE_POSITION  report mouse movement

There similar constants for mouse button #3 and #4.

There are some important things to note.

1. Mouse buttons 1, 2, and 3 may be right, center, and left respectively, instead of right, left, and center.

2. The BUTTON_SHIFT, BUTTON_ALT, and BUTTON_CTRL codes will probably not work, as they will probably be intercepted by the window manager.

3. The mouse driver will not send both a click/release pair and a pressed message.

4.12.2 Getting the events

Once a class of mouse events have been enabled, getch and friends return KEY_MOUSE every time some mouse event happens. Then, the mouse event can be retrieved with getmouse.

getmouse returns a list of five elements: id, x, y, z, and flags. The flags contain information about mouse button events.

Schematically, decoding mouse events could look like this. (Note that I’ve used some of the srfi-1 list functions in this example.)

     (set! c (getch win))
     (if (eqv? c KEY_MOUSE)
         (let* ((m (getmouse))

                (mouse-x (second m))
                (mouse-y (third m))
                (mouse-flag (fifth m)))
            (cond
             ((logtest BUTTON1_PRESSED mouse-flag)
              ; do button1-pressed response here
             )
             ((logtest BUTTON1_RELEASED mouse-flag)
              ; do button1-released response here
             ))))

4.12.3 Miscellaneous functions

The function mouse-trafo can be used to convert mouse coordinates to screen relative coordinates.

The mouseinterval function sets the maximum time (in thousandths of a second) that can elapse between press and release events in order for them to be recognized as a click. The default is one-fifth of a second.

4.13 Panels Library

Curses was originally build around the idea of having tiled text windows, where they did not overlap. The API is not set up to do the bookkeeping for overlapping windows.

Now that you are proficient in curses, you wanted to do some thing big. You created a lot of overlapping windows to give a professional windows-type look. Unfortunately, it soon becomes difficult to manage these. The multiple refreshes, updates plunge you into a nightmare. The overlapping windows create blotches, whenever you forget to refresh the windows in the proper order.

Don’t despair. There’s an elegant solution provided in panels library.

When your interface design is such that windows may dive deeper into the visibility stack or pop to the top at runtime, the resulting book-keeping can be tedious and difficult to get right. Hence the panels library.

If you have lot of overlapping windows, then panels library is the way to go. It obviates the need of doing series of noutrefresh, doupdate and relieves the burden of doing it correctly (bottom up). The library maintains information about the order of windows, their overlapping and update the screen properly. So why wait? Let’s take a close peek into panels.

4.13.1 Panels Basics

The #<panel> object is a window that is implicitly treated as part of a deck including all other panel objects. The deck is treated as a stack with the top panel being completely visible and the other panels may or may not be obscured according to their positions. So the basic idea is to create a stack of overlapping panels and use panels library to display them correctly. There is a function similar to refresh which, when called , displays panels in the correct order. Functions are provided to hide or show panels, move panels, change its size etc.. The overlapping problem is managed by the panels library during all the calls to these functions.

The general flow of a panel program goes like this:

1. Create the windows (with newwin) to be attached to the panels.

2. Create panels with the chosen visibility order. Stack them up according to the desired visibility. The function new-panel is used to create panels.

3. Call update-panels to write the panels to the virtual screen in correct visibility order. Do a doupdate to show it on the screen.

4. Manipulate the panels with show-panel, hide-panel, move-panel etc. Make use of

helper functions like panel-hidden and panel-window.

5. When you are done with the panel use del-panel to delete the panel.

Let’s make the concepts clear, with some programs. The following is a simple program which creates 3 overlapping panels and shows them on the screen.

4.13.2 Compiling With the Panels Library

To use panels library functions, you have to use the module (ncurses panel).

#!/usr/bin/guile
!#
(use-modules (ncurses curses)
           (ncurses panel))
(define stdscr (initscr))
(let* ((win1 (newwin 5 10 5 5))
     (win2 (newwin 5 10 7 7))
     (panel1 (new-panel win1))
     (panel2 (new-panel win2)))
(box win1 (acs-vline) (acs-hline))
(addstr win1 "box 1"  #:y 1 #:x 1)
(box win2 (acs-vline) (acs-hline))
(addstr win2 "box 2" #:y 1 #:x 1)
(update-panels)
(doupdate)
(sleep 1)
;; Move panel 1 to the bottom
(addstr win1 "bottom" #:y 1 #:x 1)
(bottom-panel panel1)
(update-panels)
(doupdate)
(sleep 1)
;; Move panel 1 to the top
(addstr win1 "top   " #:y 1 #:x 1)
(top-panel panel1)
(update-panels)

(doupdate)
(sleep 1)
;; Hide panel 1
(addstr win1 "hide   " #:y 1 #:x 1)
(hide-panel panel1)
(update-panels)
(doupdate)
(sleep 1)
;; Unhide panel 1
(addstr win1 "show   " #:y 1 #:x 1)
(show-panel panel1)
(update-panels)
(doupdate)
(sleep 1)
;; Move panel 1
(addstr win1 "move   " #:y 1 #:x 1)
(move-panel panel1 2 2)
(update-panels)
(doupdate)
(sleep 1))
(endwin)

As you can see, above program follows a simple flow. The windows are created with newwin and then they are attached to panels with new-panel. As we attach one panel after another, the stack of panels gets updated. To put them on screen update-panels and doupdate are called.

The panels can be brought to the front with top-panel or pushed to the back with bottom-panel. They can be removed with hide-panel, and then unhidden with show-

panel.

4.13.3 Moving and Resizing Panels

The procedure move-panel can be used to move a panel to the desired location. It does not change the position of the panel in the stack. Make sure that you use move-panel instead mvwin on the window associated with the panel.

Resizing a panel is slightly complex. There is no straight forward function just to resize the window associated with a panel. A solution to resize a panel is to create a new window with the desired sizes, change the window associated with the panel using replace-panel!. Don’t forget to delete the old window. The window associated with a panel can be found by using the function panel-window.

4.14 Menu Library

The menus library provides a nice extension to basic curses, through which you can create menus. It provides a set of functions to create menus. But they have to be customized to give a nicer look, with colors etc. Let’s get into the details.

A menu is a screen display that assists the user to choose some subset of a given set of items. To put it simple, a menu is a collection of items from which one or more items can be chosen. Some readers might not be aware of multiple item selection capability. The menu library provides functionality to write menus from which the user can chose more than one item as the preferred choice. This is dealt with in a later section. Now it is time for some rudiments.

4.14.1 Menu basics

To create menus, you first create items, and then post the menu to the display. After that, all the processing of user responses is done in an elegant function menu-driver which is the work horse of any menu program.

The general flow of control of a menu program looks like this.

1. Initialize curses

2. Create items using new-item. You can specify a name and description for the items.

3. Create the menu with new-menu by specifying the items with which it is to be attached.

4. Post the menu with menu-post and refresh the screen.

5. Process the user requests with a loop and do necessary updates to menu with menu-

    driver.

6. Unpost the menu with menu-unpost.

7. End curses.

Let’s see a program which prints a simple menu and updates the current selection with up, down arrows.

To use menu library functions, you have to use the module (ncurses menu).

     #!/usr/local/bin/guile
     -s
     !#
     (use-modules (srfi srfi-1)
                  (ncurses curses)
                  (ncurses menu))
     (define stdscr (initscr))
     (cbreak!)
     (noecho!)
     (keypad! stdscr #t)
     (let* (;; Labels for the menu items
            (names '("Choice 1" "Choice 2" "Choice 3" "Choice 4" "Exit"))
            (descriptions '("Description 1" "Description 2" "Description 3"

               "Description 4" ""))
;; Create menu items for each label
(my-items (map (lambda (name desc) (new-item name desc))
              names
              descriptions))
;; Create the menu
(my-menu (new-menu my-items)))
;; Draw the menu
(move stdscr (- (lines) 2) 0)
(addstr stdscr "Press 'q' to Quit")
(post-menu my-menu)
(refresh stdscr)
;; Process the up and down arrow keys.  Break the loop if F1 is
;; pressed.  Ignore other keys.
(let loop ((c (getch stdscr)))
(cond
;; Move down the menu when down arrow is pressed and then loop.
((eqv? c KEY_DOWN)
(begin
(menu-driver my-menu REQ_DOWN_ITEM)
(loop (getch stdscr))))
;; Move up the menu when the up arrow is pressed and then loop.
((eqv? c KEY_UP)
(begin
(menu-driver my-menu REQ_UP_ITEM)
(loop (getch stdscr))))
;; When enter is pressed, return the selection and quit.
((or (eqv? c KEY_ENTER)
  (eqv? c #\cr)
  (eqv? c #\nl))
(begin
(unpost-menu my-menu)
(move stdscr (- (lines) 4) 0)
(addstr stdscr
        (format #f "You selected item #~a: ~a"
                (item-index (current-item my-menu))
                (item-name (current-item my-menu))))
(refresh stdscr)
(sleep 2)))
;; If 'Q' or 'q'  is pressed, quit.  Otherwise, loop.

          ((not (or (eqv? c #\Q) (eqv? c #\q)))
           (loop (getch stdscr)))))
       (endwin))

This program demonstrates the basic concepts involved in creating a menu using menus library. First we create the items using new-item and then attach them to the menu with new-menu function. After posting the menu and refreshing the screen, the main processing loop starts. It reads user input and takes corresponding action. The function menu-driver is the main work horse of the menu system. The second parameter to this function tells what’s to be done with the menu. According to the parameter, menu-driver does the corresponding task. The value can be either a menu navigational request, an ASCII character, or a KEY_MOUSE special key associated with a mouse event.

The menu driver accepts following navigational requests.

REQ_LEFT_ITEM

Move left to an item.

REQ_RIGHT_ITEM

Move right to an item.

REQ_UP_ITEM

Move up to an item.

REQ_DOWN_ITEM

Move down to an item.

REQ_SCR_ULINE

Scroll up a line.

REQ_SCR_DLINE

Scroll down a line.

REQ_SCR_DPAGE

Scroll down a page.

REQ_SCR_UPAGE

Scroll up a page.

REQ_FIRST_ITEM

Move to the first item.

REQ_LAST_ITEM

Move to the last item.

REQ_NEXT_ITEM

Move to the next item.

REQ_PREV_ITEM

Move to the previous item.

REQ_TOGGLE_ITEM

Select/deselect an item.

REQ_CLEAR_PATTERN

Clear the menu pattern buffer.

REQ_BACK_PATTERN

Delete the previous character from the pattern buffer.

REQ_NEXT_MATCH

Move to the next item matching the pattern match.

REQ_PREV_MATCH

Move to the previous item matching the pattern match.

Don’t get overwhelmed by the number of options. We will see them slowly one after another. The options of interest in this example are REQ_UP_ITEM and REQ_DOWN_ITEM. These two options when passed to menu driver, menu driver updates the current item to one item up or down respectively.

4.14.2 Menu Driver: The work horse of the menu system

As you have seen in the above example, menu-driver plays an important role in updating the menu. It is very important to understand various options it takes and what they do. As explained above, the second parameter to menu-driver can be either a navigational request, a printable character or a KEY_MOUSE key. Let’s dissect the different navigational requests.

REQ_LEFT_ITEM and REQ_RIGHT_ITEM

A menu can be displayed with multiple columns for more than one item. This can be done by using the menu-format function. When a multicolumnar menu is displayed these requests cause the menu driver to move the current selection to left or right.

REQ_UP_ITEM and REQ_DOWN_ITEM

These two options you have seen in the above example. These options when given, makes the menu-driver to move the current selection to an item up or down.

REQ_SCR_* options
The four options REQ_SCR_ULINE, REQ_SCR_DLINE, REQ_SCR_DPAGE, REQ_SCR_

UPAGE are related to scrolling. If all the items in the menu cannot be displayed in the menu sub window, then the menu is scrollable. These requests can be given to the menu-driver to do the scrolling either one line up, down or one page down or up respectively.

REQ_FIRST_ITEM, REQ_LAST_ITEM, REQ_NEXT_ITEM and REQ_PREV_ITEM

Easy enough.

REQ_TOGGLE_ITEM

This request when given, toggles the present selection. This option is to be used only in a multivalued menu. So to use this request the option O_ONEVALUE must be off. This option can be made off or on with set_menu_opts.

Pattern Requests

Every menu has an associated pattern buffer, which is used to find the nearest match to the ASCII characters entered by the user. Whenever ASCII characters are given to menu driver, it puts in to the pattern buffer. It also tries to find the nearest match to the pattern in the items list and moves current selection to that item. The request REQ_CLEAR_PATTERN clears the pattern buffer. The request REQ_BACK_PATTERN deletes the previous character in the pattern buffer. In case the pattern matches more than one item then the matched items can be cycled through REQ_NEXT_MATCH and REQ_PREV_MATCH which move the current selection to the next and previous matches respectively.

Mouse Requests

In case of KEY_MOUSE requests, according to the mouse position an action is taken accordingly. The action to be taken is explained in the man page as, If the second argument is the KEY_MOUSE special key, the associated mouse event is translated into one of the above pre-defined requests. Currently only clicks in the user window (e.g. inside the menu display area or the decoration window) are handled. If you click above the display region of the menu, a REQ_SCR_ ULINE is generated, if you doubleclick a REQ_SCR_UPAGE is generated and if you tripleclick a REQ_FIRST_ITEM is generated. If you click below the display region of the menu, a REQ SCR DLINE is generated, if you doubleclick a REQ_ SCR_DPAGE is generated and if you tripleclick a REQ_LAST_ITEM is generated. If you click at an item inside the display area of the menu, the menu cursor is positioned to that item.

Each of the above requests will be explained in the following lines with several examples whenever appropriate.

4.14.3 Menu Windows

Every menu created is associated with a window and a subwindow. The menu window displays any title or border associated with the menu. The menu subwindow displays the menu items currently available for selection. But we didn’t specify any window or sub window in the simple example. When a window is not specified, stdscr is taken as the main window, and then menu system calculates the subwindow size required for the display of items. Then items are displayed in the calculated sub window. So let’s play with these windows and display a menu with a border and a title.

#!/usr/bin/guile
-s
!#
(use-modules (srfi srfi-1)
        (ncurses curses)
        (ncurses menu))
(define stdscr (initscr))
(start-color!)
(cbreak!)
(noecho!)
(keypad! stdscr #t)
(init-pair! 1 COLOR_RED COLOR_BLACK)
(let* (;; Labels for the menu items
  (names '("Choice 1" "Choice 2" "Choice 3" "Choice 4" "Exit"))

     (descriptions '("Description 1" "Description 2" "Description 3"
                     "Description 4" ""))
     ;; Create menu items for each label
     (my-items (map (lambda (name desc) (new-item name desc))
                    names
                    descriptions))
     ;; Create the menu
     (my-menu (new-menu my-items))
     ;; Make a windows to hold the menu
     (my-menu-win (newwin 10 40 4 4)))
(keypad! my-menu-win #t)
;; Set the main window and subwindow
(set-menu-win! my-menu my-menu-win)
(set-menu-sub! my-menu (derwin my-menu-win 6 38 3 1))
;; Set the menu mark string
(set-menu-mark! my-menu " * ")
;; Print a border around the main window, and a title
(box my-menu-win 0 0)
(attr-on! my-menu-win (color-pair 1))
(move my-menu-win 1 16)
(addstr my-menu-win "My Menu")
(attr-off! my-menu-win (color-pair 1))
(move my-menu-win 2 0)
(addch my-menu-win (acs-ltee))
(move my-menu-win 2 1)
(hline my-menu-win (acs-hline) 38)
(move my-menu-win 2 39)
(addch my-menu-win (acs-rtee))
(move stdscr (- (lines) 2) 0)
(addstr stdscr "F1 to exit")
(refresh stdscr)
;; Post the menu
(post-menu my-menu)
(refresh my-menu-win)
;; Draw the menu
(move stdscr (- (lines) 2) 0)

(addstr stdscr "F1 to Exit")
;; Process the up and down arrow keys.  Break the loop if F1 is
;; pressed.  Ignore other keys.
(let loop ((c (getch my-menu-win)))
(cond
;; Move down the menu when down arrow is pressed and then loop.
((eqv? c KEY_DOWN)
(begin
(menu-driver my-menu REQ_DOWN_ITEM)
(loop (getch my-menu-win))))
;; Move up the menu when the up arrow is pressed and then loop.
((eqv? c KEY_UP)
(begin
(menu-driver my-menu REQ_UP_ITEM)
(loop (getch my-menu-win))))
;; When enter is pressed, return the selection and quit.
((or (eqv? c KEY_ENTER)
  (eqv? c #\cr)
  (eqv? c #\nl))
(begin
(unpost-menu my-menu)
(move stdscr (- (lines) 4) 0)
(addstr stdscr
        (format #f "You selected item #~a: ~a"
                (item-index (current-item my-menu))
                (item-name (current-item my-menu))))
(refresh stdscr)
(sleep 2)))
;; If F1 is pressed, quit.  Otherwise, loop.
((not (eqv? c (key-f 1)))
(loop (getch stdscr)))))
(endwin))

This example creates a menu with a title, border, a fancy line separating title and the items. As you can see, in order to attach a window to a menu the function set-menu-win! has to be used. Then we attach the sub window also. This displays the items in the sub window. You can also set the mark string which gets displayed to the left of the selected

item with set-menu-mark!.

4.14.4 Scrolling Menus

If the subwindow given for a window is not big enough to show all the items, then the menu will be scrollable. When you are on the last item in the present list, if you send REQ_DOWN_ITEM, it gets translated into REQ_SCR_DLINE and the menu scrolls by one item. You can manually give REQ_SCR_operations to do scrolling.

4.15 Forms Library

Well, if you have seen those forms on web pages that take input from users and do various kinds of things, you might be wondering how one would create such form in a text mode display. It’s quite difficult to write those nifty forms in plain ncurses. The forms library tries to provide a basic framework to build and maintain forms with ease. It has lots of features that manage validation, dynamic expression of fields, and so on. Let’s see it in full flow.

A form is a collection of fields. Each field can be either a label (static text) or a dataentry location. The forms library also provides functions to divide forms into multiple pages.

4.15.1 The Basics

Forms are created in much the same way as menus. First, the fields related to the form are created with new-field. You can set options for the fields, so that they can be displayed with some fancy attributes, validated before the field looses focus, etc. Then the fields are attached to the form. After this, the form can be posted to display and is ready to receive inputs. Along similar lines to menu-driver, the form is manipulated with form-driver. We can send requests to form-driver to move focus to a certain field, move the cursor to the end of the field, and so on. After the user enters the values in the fields and validation is done, the form can be unposted and the memory allocated can be freed.

The general flow of control of a form program looks like this

1. Initialize curses

2. Create fields using new-field. You can specify the height and width of the field, and its position on the form.

3. Create the forms with new-form by specifying the fields to be attached.

4. Post the form with form-post and refresh the screen.

5. Process the user requests with a loop and do necessary updates to the form with form-

    driver.

6. Unpost the menu with form-unpost

7. If desired, explicitly free the memory allocated to menu with free-form.

8. If desired, explicitly free the memory allocated to menu item with free-field.

9. End curses.

The forms library is similar to the menu library. The following examples will explore various aspects of form processing. Let’s stat the journey with a simple example first.

4.15.2 A Form Example

To use form library functions, you have to include (ncurses form).

#!/usr/bin/guile
!#
(use-modules (srfi srfi-1)
             (ncurses curses)
             (ncurses form))
;; Initialize curses
(define stdscr (initscr))
(cbreak!)
(noecho!)
(keypad!  stdscr #t)
;; Initialize the fields
(define field (list
               (new-field 1 10 4 18 0 0)
               (new-field 1 10 6 18 0 0)))
;; Set field options
;; Print a line for the options
(set-field-back!  (first field) A_UNDERLINE)
;; Don't go to the next field when this field is filled up
(field-opts-off!  (first field) O_AUTOSKIP)
(set-field-back!  (second field) A_UNDERLINE)
(field-opts-off!  (second field) O_AUTOSKIP)
;; Create the new form and post it
(define my-form (new-form field))
(post-form my-form)
(refresh stdscr)
(addstr stdscr "Value 1:" #:y 4 #:x 10)
(addstr stdscr "Value 2:" #:y 6 #:x 10)
(refresh stdscr)
;; Loop through to get user requests
(let loop ((ch (getch stdscr)))
  (if (not (eqv?  ch (key-f 1)))
      (cond
       ((eqv?  ch KEY_DOWN)
        (begin
          ;; Go to the end of the next field
          (form-driver my-form REQ_NEXT_FIELD)
          (form-driver my-form REQ_END_LINE)
          (loop (getch stdscr))))
       ((eqv?  ch KEY_UP)
        (begin

               ;; Go to the end of the previous field
               (form-driver my-form REQ_PREV_FIELD)
               (form-driver my-form REQ_END_LINE)
               (loop (getch stdscr))))
            (else
             (begin
               ;; Print any normal character
               (form-driver my-form ch)
               (loop (getch stdscr)))))))
     ;; Unpost the form
     (unpost-form my-form)
     (endwin)

The above example is pretty straightforward. It creates two fields with new-field. The procedure new-field takes height, width startx, starty, the number of off-screen rows, and number of additional working buffers. The fifth argument number of off-screen rows specified how much of the field to be shown. If it is zero, the entire field is always displayed; otherwise, the form will be scrollable when the user accesses undisplayed part of the field. The forms library allocates one buffer per field to store the data which the user enters. Using the last parameter to new-field we can specify it to allocate some additional buffers. These can be used for any purpose you like.

After creating the fields, the background attribute of both of them is set to an underscore

with set-field-back!. The AUTOSKIP option is turned off using field-opts-off!. If this

option is turned of, focus will move to the next field in the form once the active field is filled up completely.

After attaching the fields to the form, it is posted. Here on, user inputs are processed in the loop, by making corresponding request to form-driver. The details of all the request to form-driver are explained later.

4.15.3 Making it useful

The previous example does create a form and allow one to move between the fields, but, it doesn’t perform the useful function of allowing the program to fetch the contents of those fields. Also, the basic editing features like BS and DELETE are not connected.

To make it useful, it needs a couple of important changes. First, it passes more of the important editing commands like BS and DELETE to the form driver for processing. Second, it fetches the final content of the fields as one would normally do in a form application.

A more complete editing loop might look like this

     ;; Loop through to get user requests
     (let loop ((ch (getch win)))
       (if (not (eqv? ch (key-f 1)))
           (begin
             (cond
              ((eqv? ch KEY_DOWN)
               (begin

            ;; Go to the end of the next field
            (form-driver my-form REQ_NEXT_FIELD)
            (form-driver my-form REQ_END_LINE)))
         ((eqv? ch KEY_UP)
          (begin
            ;; Go to the end of the previous field
            (form-driver my-form REQ_PREV_FIELD)
            (form-driver my-form REQ_END_LINE)))
         ((eqv? ch KEY_LEFT)
          (form-driver my-form REQ_PREV_CHAR))
         ((eqv? ch KEY_RIGHT)
          (form-driver my-form REQ_NEXT_CHAR))
         ((eqv? ch KEY_DC)
          (form-driver my-form REQ_DEL_CHAR))
         ((eqv? ch KEY_BACKSPACE)
          (form-driver my-form REQ_DEL_PREV))
         (else
          (form-driver my-form ch)))
        (loop (getch win)))))

And, gathering the contents of the form can be accomplished like this:

;; Move the cursor to ensure that the last characters typed by the
;; user get committed to the field buffer.
(form-driver my-form REQ_LAST_FIELD)
;; Unpost the form
(unpost-form my-form)
;; Store the contents of the field
(let ((result1 (field-buffer (first field) 0))
      (result2 (field-buffer (second field) 0)))
  (endwin)
  (display (string-append "You typed " result1 " and " result2))
  (newline))

4.15.4 Playing with fields

Each form field is associated with lots of attributes. They can be manipulated to get the required effect.

4.15.4.1 Fetching Size and Location of Field

The parameters we have given at the time of creation of a field can be retrieved with

field-info.
(field-info field)

The procedure field-info takes a field and returns a list: height, width, top, left, offscreen, and number of buffers.

The location of the field can be moved to a different position with move-field, but, only if this field hasn’t become part of a form. (This function is thus mostly useless.)

(move-field field top left)

The justification to be done for the field can be fixed using the function set-field-just!

(set-field-just! field justification)
The justification mode value is the either NO_JUSTIFICATION, JUSTIFY_RIGHT, JUSTIFY_
LEFT, or JUSTIFY_CENTER.

The procedure field-just returns the justification mode.

4.15.5 Field Display Attributes

As you have seen, in the above example, display attributes for the fields can be set with set-field-fore! and set-field-back!. These functions set foreground and background attributes of the fields. You can also specify a pad character that will be filled in the unfilled portion of the field. The pad character is set with a call to set-field-pad!. Default pad value is space. The functions field-fore , field-back, and field-pad can be used to query the present foreground background attributes and pad character for the field.

Though the functions seem quite simple, using colors with set-field-fore! may be frustrating in the beginning. Let me first explain about foreground and background attributes of a field. The foreground attribute is associated with the character. That means a character in the field is printed with the attribute you have set with set-field-fore!. Background attribute is the attribute used to fill background of field, whether any character is there or not. So what about colors? Since colors are always defined in pairs, what is the right way to display colored fields? Here’s an example clarifying color attributes.

#!/usr/bin/guile
!#
(use-modules (srfi srfi-1)
             (ncurses curses)
             (ncurses form))
;; Initialize curses
(define stdscr (initscr))
(start-color!)
(cbreak!)
(noecho!)
(keypad!  stdscr #t)
;; Initialize the color pairs
(init-pair!  1 COLOR_WHITE COLOR_BLUE)
(init-pair!  2 COLOR_WHITE COLOR_BLUE)
;; Initialize the fields
(define field (list
               (new-field 1 10 4 18 0 0)
               (new-field 1 10 6 18 0 0)))

(set-field-fore!  (first field) (color-pair 1))
(set-field-back!  (first field) (color-pair 2))
(field-opts-off!  (first field) O_AUTOSKIP)
(set-field-back!  (second field) A_UNDERLINE)
(field-opts-off!  (second field) O_AUTOSKIP)
;; Create the new form and post it
(define my-form (new-form field))
(post-form my-form)
(refresh stdscr)
(addstr stdscr "Value 1:" #:y 4 #:x 10)
(addstr stdscr "Value 2:" #:y 6 #:x 10)
(refresh stdscr)
;; Loop through to get user requests
(let loop ((ch (getch stdscr)))
  (if (not (eqv?  ch (key-f 1)))
      (cond
       ((eqv?  ch KEY_DOWN)
        (begin
          ;; Go to the end of the next field
          (form-driver my-form REQ_NEXT_FIELD)
          (form-driver my-form REQ_END_LINE)
          (loop (getch stdscr))))
       ((eqv?  ch KEY_UP)
        (begin
          ;; Go to the end of the previous field
          (form-driver my-form REQ_PREV_FIELD)
          (form-driver my-form REQ_END_LINE)
          (loop (getch stdscr))))
       (else
        (begin
          ;; Print any normal character
          (form-driver my-form ch)
          (loop (getch stdscr)))))))
;; Unpost the form
(unpost-form my-form)
(endwin)

Play with the color pairs and try to understand the foreground and background attributes. Color pair 1 is used for the character that have been typed into the form, and color pair 2 is used for the empty spaces in the form.

4.15.6 Field Option Bits

There is also a large collection of field option bits you can set to control various aspects of forms processing. You can manipulate them with the procedures set-field-opts!,

field-opts-on!, field-opts-off!, and field-opts.

The function set-field-opts! can be used to directly set attributes of a field or you can chose to switch a few attributes on and off with field-opts-on! and field-opts-off!. You can query the attributes of a field with field-opts. The following is a list of available options. By default, all options are on.

O_VISIBLE

Controls whether the field is visible on the screen. Can be used during form processing to hide or pop up fields depending on the value of parent fields.

O_ACTIVE

Controls whether the field is active during forms processing (i.e. visited by form navigation keys). Can be used to make labels or derived fields with buffer values alterable by the forms application, not the user.

O_PUBLIC

Controls whether data is displayed during field entry. If this option is turned off on a field, the library will accept and edit data in that field, but it will not be displayed and the visible field cursor will not move. You can turn off the O_PUBLIC bit to define password fields.

O_EDIT

Controls whether the field’s data can be modified. When this option is off, all editing requests except REQ_PREV_CHOICE and REQ_NEXT_CHOICE will fail. Such read-only fields may be useful for help messages.

O_WRAP

Controls word-wrapping in multiline fields. Normally, when any character of a (blank-separated) word reaches the end of the current line, the entire word is wrapped to the next line (assuming there is one). When this option is off, the word will be split across the line break.

O_BLANK

Controls field blanking. When this option is on, entering a character at the first field position erases the entire field (except for the just-entered character).

O_AUTOSKIP

Controls automatic skip to next field when this one fills. Normally, when the forms user tries to type more data into a field than will fit, the editing location jumps to the next field. When this option is off, the user’s cursor will hang at the end of the field. This option is ignored in dynamic fields that have not reached their size limit.

O_NULLOK

Controls whether validation is applied to blank fields. Normally, it is not; the user can leave a field blank without invoking the usual validation check on exit. If this option is off on a field, exit from it will invoke a validation check.

O_PASSOK

Controls whether validation occurs on every exit, or only after the field is modified. Normally the latter is true. Setting O_PASSOK may be useful if your field’s validation function may change during forms processing.

O_STATIC

Controls whether the field is fixed to its initial dimensions. If you turn this off, the field becomes dynamic and will stretch to fit entered data.

A field’s options cannot be changed while the field is currently selected. However, options may be changed on posted fields that are not current.

The option values are bit-masks and can be composed with logior in the obvious way. You have seen the usage of watching off O_AUTOSKIP option. The following example clarified usage of some more options. One field is a constant, uneditable field, that you cannot move to by using the arrow keys. The other field is a password-like field where the characters are hidden from view.

#!/usr/local/bin/guile
!#
(use-modules (srfi srfi-1)
             (ncurses curses)
             (ncurses form))
;; Constants
(define STARTX 15)
(define STARTY 4)
(define WIDTH 25)
(define N_FIELDS 2)
;; Initialize curses
(define stdscr (initscr))
(cbreak!)
(noecho!)
(keypad! stdscr #t)
;; Initialize the fields
(define field (map-in-order
               (lambda (y)
                 (new-field 1
                            WIDTH
                            (+ (* y 2) STARTY)
                            STARTX
                            0
                            0))
               (iota N_FIELDS)))

(set-field-back! (first field) A_UNDERLINE)
;; Don't go to the next field when this field is filled up
;; This field is a static label
(field-opts-off! (first field) O_ACTIVE)
;; This field is like a password field
(field-opts-off! (second field) O_PUBLIC)
(field-opts-off! (second field) O_AUTOSKIP)
;; Create the new form and post it
(define my-form (new-form field))
(post-form my-form)
(refresh)
(set-field-just! (first field) JUSTIFY_CENTER)
(set-field-buffer! (first field) 0 "This is a static field")
(addstr "Field 1:" #:y STARTY #:x (- STARTX 10))
(addstr "Field 2:" #:y (+ STARTY 2) #:x (- STARTX 10))
(refresh)
;; Loop through to get user requests
(let loop ((ch (getch)))
  (if (not (eqv? ch (key-f 1)))
      (cond
       ((eqv? ch KEY_DOWN)
        (begin
          ;; Go to the end of the next field
          (form-driver my-form REQ_NEXT_FIELD)
          (form-driver my-form REQ_END_LINE)
          (loop (getch))))
       ((eqv? ch KEY_UP)
        (begin
          ;; Go to the end of the previous field
          (form-driver my-form REQ_PREV_FIELD)
          (form-driver my-form REQ_END_LINE)
          (loop (getch))))
       (else
        (begin
          ;; Print any normal character
          (form-driver my-form ch)
          (loop (getch)))))))
;; Unpost the form
(unpost-form my-form)

     (endwin)

4.15.7 Field Status

The field status specifies whether the field has been edited. It is initially #f and when the user enters something and the data buffer gets modified, it becomes #t. So a field’s status can be queried to find out if it has been modified or not.

The two procedures that work with field status are set-field-status! and field-status?. The procedure field-status? returns the current state, and set-field-status! can be used to set the state.

4.15.8 Variable-Sized Fields

If you want a dynamically changing field with variable width, this is the feature you want to put to full use. This will allow the user to enter more data than the original size of the field and let the field grow. According to the field orientation it will scroll horizontally or vertically to incorporate the new data.

To make a field dynamically growable, the option O_STATIC should be turned off. This can be done with

     (field-opts-off! field O_STATIC)

It is usually not advisable to allow a field to grow infinitely.. You can set a maximum limit to the growth of the field with set-max-field.

The field info for a dynamically growable field can be retrieved with the procedure

dynamic-field-info.

Recall the library routine new-field; a new field created with height set to one will be defined to a one line field. A new field created with height greater than one will be defined to be a multi-line field.

A one line field with O_STATIC turned off (dynamically growable) will contain a single fixed row, but, the number of columns can increase if the user enters more data than the initial field will hole. The number of columns displayed will remain fixed and the additional data will scroll horizontally.

A multiline field with O_STATIC turned off (dynamically growable) will contain a fixed number of columns, but, the number of rows can increase if the user enters more data than the initial field will hold. The number of rows displayed will remain fixed and the additional data will scroll vertically.

4.15.9 Form Windows

The form windows concept is pretty much similar to menu windows. Every form is associated with a main window and a subwindow. The form main window displays any title or border associated or whatever the user wishes. The subwindow contains all the fields and displays them according to their position. This gives the flexibility of manipulating fancy form displays easily.

Since this pretty similar to menu windows, I am providing a very similar example.

     #!/usr/local/bin/guile
     !#

(use-modules (srfi srfi-1)
             (ncurses curses)
             (ncurses form))
;; Helper procedure to center a text
(define (print-in-middle win starty startx width str color)
  (let ((length (string-length str)))
    (attr-on! win color)
    (addstr win str
            #:y starty
            #:x (+ startx (/ (- width length) 2)))
    (attr-off! win color)
    (refresh)))

;; Initialize curses
(define stdscr (initscr))
(start-color!)
(cbreak!)
(noecho!)
(keypad! stdscr #t)
;;Initialize a few color pairs
(init-pair! 1 COLOR_RED COLOR_BLACK)
(define field (list
               (new-field 1 10 6 1 0 0)
               (new-field 1 10 8 1 0 0)))
;; Set field options
(set-field-back! (first field) A_UNDERLINE)
(field-opts-off! (first field) O_AUTOSKIP)
(set-field-back! (second field) A_UNDERLINE)
(field-opts-off! (second field) O_AUTOSKIP)
;; Create a new form
(define my-form (new-form field))
;; Calculate the area associated with the form
(define xy (scale-form my-form))
(define rows (car xy))
(define cols (cadr xy))

;; Create the window to be associated with the form
(define my-form-win (newwin (+ 4 rows)
                            (+ 4 cols)
                            4
                            4))
(keypad! my-form-win #t)
;; Set main window and subwindow
(set-form-win! my-form my-form-win)
(set-form-sub! my-form (derwin my-form-win rows cols 2 2))
;; Print a border around the main window and print a title
(box my-form-win 0 0)
(print-in-middle my-form-win 1 0 (+ cols 4) "My Form" (color-pair 1))
(post-form my-form)
(refresh my-form-win)
(addstr stdscr "Use UP, DOWN arrow keys to switch between fields"
        #:y (- (lines) 2) #:x 0)
(refresh stdscr)
;; Loop through to get user requests
(let loop ((ch (getch my-form-win)))
  (if (not (eqv? ch (key-f 1)))
      (cond
       ((eqv? ch KEY_DOWN)
        (begin
          ;; Go to the end of the next field
          (form-driver my-form REQ_NEXT_FIELD)
          (form-driver my-form REQ_END_LINE)
          (loop (getch my-form-win))))
       ((eqv? ch KEY_UP)
        (begin
          ;; Go to the end of the previous field
          (form-driver my-form REQ_PREV_FIELD)
          (form-driver my-form REQ_END_LINE)
          (loop (getch my-form-win))))
       (else
        (begin
          ;; Print any normal character
          (form-driver my-form ch)
          (loop (getch my-form-win)))))))
;; Unpost the form
(unpost-form my-form)

(endwin)

4.15.10 Field Validation

By default, a field will accept any data input by the user. It is possible to attach validation to the field. Then any attempt by the user to leave the field while it contains data that doesn’t match the validation type will fail.. Some validation types also have a charactervalidity check for each type a character is entered in the field.

Validation can be attached by using the set-field-type! procedure and queried with the field-type procedure.

The form driver validates the data in a field only when data is entered by the enduser. Validation does not occur when the application program changes the field value with

set-field-buffer!.

The validation types are as follows.

TYPE_ALPHA

This field type accepts alphabetic data; no blanks, no digits, no special characters. It takes a width argument that sets a minimum width of the data. The user has to enter at least that number of characters be fore he can leave the field. Typically you’ll want to set this to the field width. If it’s greater than the field width, the validation check will always fail. A minimum width of zero makes field completion optional.

     (set-field-type! field TYPE_ALPHA width)
TYPE_ALNUM

This field type accepts alphabetic data and digits. No blanks, no special characters. It also has a width argument.

     (set-field-type! field TYPE_ALPHA width)
TYPE_ENUM

This type allows you to restrict a field’s values to be among a specified set of string values (for example, the two-letter postal codes for US states). It takes as input a list of strings. It can be set up as either case sensitive or case insensitive. When the user exits, the validation procedure tries to complete the data in the buffer to a valid entry. If a complete choice string has been entered, it is, of course, valid. But it is also possible to enter a prefix of a valid string and have it completed for you. By default, if you enter such a prefix and it matches more than on value in the string list, the prefix will be completed to the first matching value. If the check-unique option is chosen, the prefix match must be unique to be valid.

     (set-field-type! field TYPE_ENUM valuelist checkcase checkunique)
TYPE_INTEGER

Valid characters consist of an optional leading minus and digits. A range check is performed on exit. If the range maximum is less than or equal to the minimum, the range is ignored.

     (set-field-type! field TYPE_INTEGER zero-padding min max)

If the value passes its range check, it is padded with as many leading zero digits as necessary to meet the padding requirement.

TYPE_NUMERIC

This accepts a decimal number.

               (set-field-type! field TYPE_NUMERIC digits-of-precision min max)

Valid characters consist of a leading minus and digits, possibly including a decimate point. The range check is performed on exit. If the value passes its range check, it is padded with as many trailing zero digits as necessary to meet the padding argument.

TYPE_REGEXP

This field type accepts a regular expression, and thus can be used to perform other types of range checks.

4.15.11 The Form Driver: The Work Horse of the System

The form-driver procedure plays a very important role in the forms system. All types of requests to forms should be funneled through form-driver.

As you have seen some of the examples above, you have to be in a loop looking for user input and then decide whether it is a field data or a form request. The form requests are then passes to form-driver to do the work.

The requests roughly can be divided into the following categories. Different request and their usage is explained below.

4.15.11.1 Page Navigation Requests

These requests cause page-level moves through the form, triggering display of a new form screen. A form can be made of multiple pages. If you have a big form with lots of fields and logical sections, then you can divide the form into pages. The function set-new-page sets a new page at the field specified.

Also, the following requests, when passed to the form driver, all you to move to different pages.

• REQ NEXT PAGE Move to the next form page.

• REQ PREV PAGE Move to the previous form page.

• REQ FIRST PAGE Move to the first form page.

• REQ LAST PAGE Move to the last form page.

These are cyclic. The moving to the next page after the last page put you at the top.

4.15.11.2 Inter-Field Navigation Requests

• REQ NEXT FIELD Move to the next field

• REQ PREV FIELD Move to the previous field

• REQ FIRST FIELD Move to the first field.

• REQ LAST FIELD Move to the last field.

• REQ SNEXT FIELD Move to the sorted next field.

• REQ SPREV FIELD Move to the sorted previous field.

• REQ SFIRST FIELD Move to sorted first field.

• REQ SLAST FIELD Move to sorted last field.

• REQ LEFT FIELD Move left to field.

• REQ RIGHT FIELD Move right to field.

• REQ UP FIELD Move up one field.

• REQ DOWN FIELD Move down one field. These requests treat fields as cyclic, moving off the end of the page will return you to

the top.

5 Curses Reference

The guile-ncurses package provides the following modules:

1. (ncurses curses): Basic functionality and input/output.

2. (ncurses panel): A library for managing overlapping windows.

3. (ncurses menu): A library for menu creation.

4. (ncurses form): A library for creating text-based form screens.

5. (ncurses slk): The soft keys functions.

5.1 The naming of routines

When you look at this library, it may seem that exclamation points and question marks were sprinkled at random on the ends of curses functions. There is actually a thought process behind it: it isn’t just chaos.

• A procedure that returns #t or #f and does not change the behavior of the system in any way ends in a question mark, e.g. has-ic?.

• But, a procedure that returns #t or #f and does change the behavior of the system does not end in a question mark, e.g. beep.

• A procedure that changes any property of curses but does not write to a window or screen ends in an exclamation point, e.g. bkgdset!.

• Any procedure that writes to a window or screen does not in an exclamation point,

    e.g. addch.

• Curses integer constants are capitalized and use underscores, for example A_BOLD.

• Curses procedures are not capitalized, even if they were capitalized macros in the C Ncurses library, for example lines and tabsize.

5.2 The basic curses library: (ncurses curses)

These are the functions in the (ncurses curses) module.

The procedures in the curses library closely follow the functions in the C-language API. Most of the procedures in the library have approximately the same name and arguments as the functions in the C-language API.

5.2.1 Overview

The GNU Guile-Ncurses library routines give the user a terminal-independent method of updating character screens with reasonable optimization.

The package supports overall screen, window and pad manipulation; output to windows and pads; reading terminal input; control over terminal and curses input and output options; environment query routines; color manipulation; use of soft label keys; terminfo capabilities; and access to low-level terminal-manipulation routines.

To initialize the routines, the routine initscr or newterm must be called before any of the other routines that deal with windows and screens are used. The routine endwin must be called before exiting. To get character-at-a-time input without echoing (most interactive, screen oriented programs want this), the following sequence should be used:

     (cbreak!)
     (noecho!)

Most programs would additionally use the sequence:

     (nonl!)
     (intrflush! #f)
     (keypad! stdscr #t))

Before a curses program is run, the tab stops of the terminal should be set and its initialization strings, if defined, must be output. This can be done by executing the tput init command after the shell environment variable TERM has been exported. tset is usually responsible for doing this. [See terminfo for further details.]

The Guile-Ncurses library permits manipulation of data structures, the #<window> type, which can be thought of as two-dimensional arrays of characters representing all or part of a CRT screen. A default window, conventionally called stdscr, which is the size of the terminal screen, is supplied. The #<window> information for stdscr is returned by calling the procedure initscr. Other windows may be created with newwin.

Note that (ncurses curses) does not easily handle overlapping windows. Overlapping windows usually require the use of the (ncurses panel) library. Without the panel library, you can either position everything on screen by hand or divide the screen into tiled windows. Chose one of the two strategies and stick with it. Mixing the two will result in unpredictable, and undesired, effects.

Windows are referred to by variables declared as #<window>. These data structures are manipulated with routines described here. Among those, the most basic routines are move and addch. move places the cursor at a location in a buffer, and addch puts a character at that location.

After using routines to manipulate a window, refresh is called, telling curses to make the user’s CRT screen look like array of characters in a #<window> buffer. Other information about the character may also be stored with each character.

Special windows called pads may also be manipulated. These are windows which are not constrained to the size of the screen and whose contents need not be completely displayed.

In addition to drawing characters on the screen, video attributes and colors may be supported, causing the characters to show up in such modes as underlined, in reverse video, or in color on terminals that support such display enhancements. Line drawing characters may be specified to be output. On input, curses is also able to translate arrow and function keys that transmit escape sequences into single values. The video attributes, line drawing characters, and input values use names, such as A_REVERSE, (acs-hline), and KEY_LEFT.

If the environment variables LINES and COLUMNS are set, or if the program is executing in a window environment, line and column information in the environment will override information read by terminfo.

If the environment variable TERMINFO is defined, any program using curses checks for a local terminal definition before checking in the standard places. For example, if TERM is set to xterm, then the compiled terminal definition is found in /usr/share/terminfo/x/xterm

(The “x” is copied from the first letter of “xterm” to avoid creation of huge directories.)

However, if TERMINFO is set to $HOME/myterms, curses first checks $HOME/myterms/x/xterm,

and if that fails, it then checks the standard location.

This is useful for developing experimental definitions or when write permission in

/usr/share/terminfo is not available.

The getter procedures (lines) and (cols) are defined in (ncurses curses) and will be return the size of the screen at the time initscr was called.

5.2.1.1 The Coordinate System

The move routine and routines that take #:y and #:x arguments use a screen based coordinate system. The coordinate y always refers to the row (of the window), and x always refers to the column. The upper left-hand corner is always (0,0). y increases as one moves down the window, and x increases as one moves left.

5.2.2 Features determined at compile time

When the Guile-Ncurses library was compiled, tests were run to see which features could be enabled. There are two variables that describe the capabilities of this build of Guile-Ncurses.

[Constant] %wide-ncurses

This is set to #t if Guile-Ncurses is using the features of the wide ncurses library libncursesw. Otherwise, it is #f.

[Constant] %ucs4-chars

This is set to #t if Guile-Ncurses expects that characters are stored as UCS4 codepoints, such as in Guile 2.0.x. It is #f if Guile-Ncurses expects that characters are 8-bit only and stored according to the current 8-bit locale.

5.2.3 Curses screen initialization and manipulation routines

initscr is normally the first curses routine to call when initializing a program. A few special routines sometimes need to be called before it; these are slk-init, %filter, ripoffline, use-env. For multiple-terminal applications, newterm may be called before initscr.

[Procedure] initscr

The initscr code determines the terminal type and initializes all curses data structures. initscr also causes the first call to refresh to clear the screen. The #<window> returned by initscr should be held in a variable that exists for the lifetime of a curses program. If it is garbage collected, the window will become unusable. If errors occur, initscr throws and error and exits; otherwise, a #<window> referencing stdscr is returned.

[Procedure] cols [Procedure] lines

These procedures, if called after initscr has been called, will return the size of the screen at the time initscr was called.

[Procedure] newterm type outport inport The procedure newterm initializes curses for a given terminal type type on specific curses port inport and outport. The newterm function relies on some relatively obscure GNU C library functions and thus may not be available on non-GNU systems.

The routine newterm should be called once for each terminal. It returns a variable of type #<screen> which should be saved as a reference to that terminal. type is the type of the terminal to be used in place of $TERM. outport is a port that will receive the output to the terminal.

IMPORTANT: These ports must be file ports. The underlying ncurses library extracts the file descriptor from the port and uses that to write to the screen. Also, these ports will be closed by newterm and cannot be reused.

newterm will internally use a duplicate port to the file pointed to by inport. inport won’t be used except to get a file descriptor for the underlying file.

Immediately following the call to newterm, one should create a window for the terminal by using the stdscr procedure. So the standard way to initialize it would be something like.

     (newterm "vt220" outport inport)
     (define stdscr (stdscr))

This routine will throw an error if the terminal could not be created.

A program that outputs to more than one terminal should use the newterm routine for each terminal instead of initscr. A program that needs to inspect capabilities, so it can continue to run in a line-oriented mode if the terminal cannot support a screen-oriented program, would also use newterm. The program must also call endwin for each terminal being used before exiting from curses. If newterm is called more than once for the same terminal, the first terminal referred to must be the last one for which endwin is called.

[Procedure] endwin

A program should always call endwin before exiting or escaping from curses mode temporarily. This routine restores TTY modes, moves the cursor to the lower lefthand corner of the screen and resets the terminal into the proper non-visual mode. Calling refresh or doupdate after a temporary escape causes the program to resume visual mode.

Its return value is #t if the terminal can be restored to its behavior (as in resetshell-mode), or #f otherwise. If the terminal was created by newterm, then this will usually return #f since input file had no previous mode.

[Procedure] isendwin?

The isendwin? routine returns #t if endwin has been called without any subsequent calls to refresh, and #f otherwise.

[Procedure] set-term new

The set-term routine is used to switch between different terminals. If the program has created multiple terminal using the newterm procedure, then set-term can be called to set one of those terminal to be the current terminal on which all the curses procedures will operate. The screen reference new becomes the new current terminal. This is the only routine which manipulates #<screen> types; all other routines affect only the current terminal.

The return value is unspecified.

[Procedure] delscreen screen

The delscreen routine frees storage associated with the screen data structure. The endwin routine does not do this, so delscreen should be called after endwin if a particular screen is no longer needed. Trying to use a screen after it has been freed will likely result in “bad state” errors. The return value is unspecified.

5.2.4 Errors and Exceptions

The curses library uses SRFI-34’s raise and SRFI-35 conditions to signal errors. All errors in the curses library are of type &curses-error.

[Procedure] curses-error? err

Returns true if err is a &curses-error

The wrong type error has two fields: arg and expected-type.

[Procedure] curses-wrong-type-arg-error? err
[Procedure] curses-wrong-type-arg-error:arg err
[Procedure] curses-wrong-type-arg-error:expected-type err

The predicate and accessors for the wrong type error.

5.2.5 Making rendered characters

Many curses functions take rendered, complex characters: characters with associated color and attribute information.

For those who are familiar with the C API for ncurses, you know that a rendered character is either a chtype, which is a 32-bit integer containing an 8-bit char and 24-bits of color and rendering information, or is a cchar_t, which is one or more wchar_t characters plus associated color and rendering information. Guile-Ncurses abstracts away the difference between these two types and presents a consistent API for both. The Guile-Ncurses complex rendered character will get converted automatically to chtype or cchar_t when necessary.

There is a family of functions to convert unrendered, simple characters or strings to rendered complex characters and strings.

[Procedure] blink x [Procedure] bold x [Procedure] dim x [Procedure] horizontal x [Procedure] invis x [Procedure] left x [Procedure] low x [Procedure] normal x [Procedure] protect x [Procedure] inverse x [Procedure] right x [Procedure] standout x [Procedure] top x [Procedure] underline x [Procedure] vertical x

These procedures take x, which can be either a simple character, a complex character, a simple string, or a complex string, and returns a rendered character or string with the attribute blink, bold, dim, horizontal, invisible, left, low, normal, protect, inverse, right, top, underline, or vertical, respectively. If the input x was a rendered character or a rendered string, the old attributes are replaced. If x was a rendered character or string with an associated color pair, the returned character or string will have the same associated color pair. Note that whether these attributes can actually be visualized depends on the capabilities of the terminal itself. Most terminals can handle bold, dim, inverse, and sometimes blink, underline and invisible. The rest should probably not be used.

[Procedure] blink-on x [Procedure] bold-on x [Procedure] dim-on x

[Procedure] horizontal-on x

[Procedure] invis-on x [Procedure] left-on x [Procedure] low-on x [Procedure] normal-on x [Procedure] protect-on x [Procedure] inverse-on x [Procedure] right-on x [Procedure] standout-on x [Procedure] top-on x

[Procedure] underline-on x

[Procedure] vertical-on x

These procedures take x, which can be either a simple character, a complex character, a simple string, or a complex string. If x is a simple character or simple string, it sets its rendering to blink, bold, dim, horizontal, invisible, left, low, normal, protect, inverse, right, top, underline, or vertical, respectively. If the input x was a rendered character or a rendered string, these attributes are added to the rendered character. Most terminals can’t actually visualize multiple attributes on a single character: some terminals can handle the combination of blink and bold, bold and underline, or inverse and bold.

[Procedure] blink-off x [Procedure] bold-off x [Procedure] dim-off x

[Procedure] horizontal-off x

[Procedure] invis-off x [Procedure] left-off x [Procedure] low-off x [Procedure] normal-off x [Procedure] protect-off x [Procedure] inverse-off x [Procedure] right-off x

[Procedure] standout-off x

[Procedure] top-off x

[Procedure] underline-off
[Procedure] vertical-off x

These procedures take x, which can be either a simple character, a complex character, a simple string, or a complex string. If x is a simple character or simple string, it sets its rendering to normal. If the input x was a rendered character or a rendered string and had the given attribute, these attributes is removed from the resulting rendered character.

[Procedure] color n x

These procedure takes x, which can be either a simple character, a complex character, a simple string, or a complex string. It returns a rendered character or string with an associated color pair n. This procedure should only be used if start-color has been called.

There are a set of primitives to operate directly on complex characters.

[Procedure] xchar? c

Returns #t if c is a complex character.

[Procedure] xchar-attr c [Procedure] xchar-color c [Procedure] xchar-chars c

These three procedures return the attributes, color pair number, and list of constiuent characters of a complex character, respectively.

[Procedure] set-xchar-attr! c attr

[Procedure] set-xchar-color! c color-pair-number

[Procedure] set-xchar-chars! c list-of-chars

These procedures directly set the attributes, color-pair number, and the list of characters of a complex character, respectively.

Complex characters can have a base character and a set of accent characters that overwrite the base character. That is why set-xchar-chars! takes a list of characters, instead of a single character. But, for these complex characters to be rendered correctly on the terminal, you need both a terminal that can do overstrike characters and you need to have compiled with the wide Ncurses library. If either of these conditions are not true, only the first character in the list of characters will appear on the screen.

There are also a couple of low-level functions to do conversion between C characters and Guile characters.

[Procedure] %scheme-char-from-c-char c

Given an 8-bit integer c that represents a C character in the current locale, this returns the associated scheme character.

[Procedure] %scheme-char-to-c-char c

This returns an 8-bit integer that is the C representation of character c in the current locale. If the character cannot be represented in the current locale, it will return the integer 63, which is the ASCII code for the question mark.

[Procedure] %scheme-char-from-c-wchar c

Given an integer c that represents a wchar t representation of a C wide character, this returns the associated scheme character.

[Procedure] %scheme-char-to-c-wchar c

This returns an integer that is the C wchar_t representation of character c. If the character cannot be represented in the current locale, it will return the integer #xFFFD, which is the Unicode codepoint for the replacement character.

5.2.6 Outputting characters

There are two primary functions that add a rendered character to a curses window, and then advance the cursor.

[Procedure] addch win ch #:key y x The addch routine puts the rendered character ch into the given window at its current window position, which is then advanced. If the advance is at the right margin, the cursor automatically wraps to the beginning of the next line. At the bottom of the current scrolling region, if scrollok! is enabled, the scrolling region is scrolled up one line. If ch is a tab, newline, or backspace, the cursor is moved appropriately within the window. Backspace moves the cursor one character left; at the left edge of a window it does nothing. Newline does a clrtoeol, then moves the cursor to the window left margin on the next line, scrolling the window if on the last line. Tabs are considered to be at every eighth column. The tab interval may be altered with the set-tabsize! procedure. If ch is any control character other than TAB, CR, or BS, it is drawn in ^X notation. Calling inch after adding a control character does not return the character itself, but instead returns the ^-representation of the control character. Since ch is a rendered character, the rendition will be applied to the character when it is put on the screen. This function returns #t on success and #f otherwise. A #f usually indicates an attempt to add a character that won’t fit on the screen or that would cause a wrap or scroll in a window that doesn’t wrap or scroll.

[Procedure] echochar win ch #:key y x The echochar routine is equivalent to a call to addch followed by a call to refresh. The knowledge that only a single character is being output allows for efficiency. This function returns #t on success and #f otherwise. A #f usually indicates an attempt to add a character that won’t fit on the screen or that would cause a wrap or scroll in a window that doesn’t wrap or scroll.

[Procedure] tabsize

Returns the current tabsize for this screen. The default is eight.

[Procedure] set-tabsize! tab

Sets the tabsize for this screen to tab characters. Its return value is unspecified.

5.2.6.1 Line Graphics

The line-graphics procedures may be used to add line drawing characters to the screen with routines of the addch family. Not all terminals have line-graphics characters. There is a default character listed below is used if the acsc capability doesn’t define a terminalspecific replacement for it. The names are taken from VT100 nomenclature. Each of these procedures takes no arguments and returns a rendered character with the ’normal’ attribute and with the default color pair. The following prints a less-than-or-equal-to sign to the window win, for example.

     (addch win (normal (acs-lequal)))

The set of line-graphics characters is given in Table 5.1.

name  default description
acs-block  #  solid square block
acs-board  #  board of squares

acs-btee       +  bottom tee
acs-bullet  o  bullet

acs-ckboard  :  checker board (stipple)
acs-darrow  v  arrow pointing down
acs-degree  ’  degree symbol

acs-diamond    +  diamond

acs-gequal     >  greater-than-or-equal-to
acs-hline  -  horizontal line
acs-lantern  #  lantern symbol
acs-larrow     <  arrow pointing left
acs-lequal     <  less-than-or-equal-to
acs-llcorner   +  lower left-hand corner
acs-lrcorner   +  lower right-hand corner

acs-ltee       +  left tee
acs-nequal  !  not-equal

acs-pi  *  greek pi

acs-plminus  #  plus/minus
acs-plus       +  plus

acs-rarrow     >  arrow pointing right

acs-rtee       +  right tee

acs-s1  -  scan line 1
acs-s3  -  scan line 3
acs-s7  -  scan line 7
acs-s9  scan line 9
acs-sterling f  pound-sterling symbol

acs-ttee       +  top tee

acs-uarrow     ^arrow pointing up
acs-ulcorner   +  upper left-hand corner
acs-urcorner   +  upper right-hand corner
acs-vline      |  vertical line

Table 5.1: the line-graphics procedures

5.2.7 Outputting strings

[Procedure] addstr win #:key y x n This routine writes the characters of the unrendered simple string str on the given window. When called with n it writes at most n characters. If n is -1, then the entire string will be added, up to the maximum number of characters that will fit on the line, or until a terminating null is reached.

It returns #f upon failure and #t on success. Failure usually indicates at attempt to write beyond the borders of the window or to write in a place that would cause scroll or wrap on a window that doesn’t support it.

[Procedure] addchstr win chstr #:key y x n These routines copy the rendered, complex string chstr into the window image structure at and after the current cursor position. If n is defined, copy at most n elements, but no more than will fit on the line. If n equals -1 or if it is undefined, then the whole string is copied, to the maximum number of characters that will fit on the line. The window cursor is not advanced, and these routines work faster than addstr. On the other hand, they do not perform any kind of checking (such as for the newline, backspace, or carriage return characters), they do not advance the current cursor position, they do not expand other control characters to ^-escapes, and they truncate the string if it crosses the right margin, rather than wrapping it around to the newline. It returns #f upon failure and #t on success.

5.2.8 Character and window attribute control routines

These routines manipulate the current attributes of the named window. The current attributes of a window apply to all characters that are written into the window with addch and addstr. Attributes are a property of the character, and move with the character through any scrolling and insert/delete line/character operations. To the extent possible, they are displayed as appropriate modifications to the graphic rendition of characters put on the screen.

[Procedure] attr-set! win attrs #:optional color

[Procedure] attr-on! win attrs [Procedure] attr-off! win attrs [Procedure] standend! win [Procedure] standout! win

The routine attr-set! sets the current rendition (attributes and color pair) of the given window to attrs. Optionally, you may split the color-pair information as a third parameter. attrs is one of the attribute constants: there is an attribute constant for each of the attributes mentioned in Section 5.2.5 [Making rendered characters], page 59: A BLINK, A BOLD, A DIM, A INVIS, A NORMAL, A PROTECT, A REVERSE, A STANDOUT, A UNDERLINE, A HORIZONTAL, A LEFT, A LOW, A RIGHT, A TOP, A VERTICAL. The following two calls are equivalent:

     (attr-set! win (logior A_BOLD (color-pair 1)))

or

     (attr-set! win A_BOLD 1)

The routine attr-off! turns off the named attributes without turning any other attributes on or off. The routine attr-on! turns on the named attributes without affecting any others. The routine standout! is the same as (attr-on! A_STANDOUT).

The routine standend! is the same as (attr-set! A_NORMAL) or (attr-set! 0),

that is, it turns off all attributes.

The return value of these routines are undefined. They can throw an exception if curses is in a bad state.

The attr-set! and related routines do not affect the attributes used when erasing portions of the window. For functions which modify the attributes used for erasing and clearing See Section 5.2.10 [Window background manipulation routines], page 67.

[Procedure] color-set! win color-pair-number The routine color-set! sets the current color of the given window to the foreground/background combination described by the color-pair-number. The return value is unspecified, but, it can throw an error if the color pair number was too large.

[Procedure] attr-get

The routine attr-get returns the current attribute and color pair for the given window. They are returned as a list containing two elements. The first element is a bitmap containing all the attributes. The second element is the color pair alone.

[Procedure] chgat win n attr color #:key y x The routine chgat changes the attributes of a given number of characters starting at the current cursor location of stdscr or of win if it is given. It does not update the cursor and does not perform wrapping. A character count of -1 or greater than the remaining window width means to change attributes all the way to the end of the current line. If y and x are given, the function does a cursor move before acting. In these functions, the color argument is a color-pair index as in the first argument of init-pair!. See Section 5.2.13 [Color manipulation routines], page 69. The return value is unspecified.

The attributes can be passed to the routines attr-on!, attr-off!, and attr-set!, or logior’d with the characters passed to addch. For the color part of the rendition, use color-pair. See Table 5.2.

[Procedure] color-pair n

Returns a bit mask to apply the color pair n to a rendition. color-pair values can only be logior’ed with attributes if the pair number is less than 256. The alternate functions such as color-set! can pass a color pair value directly.

For example, the following two calls are equivalent ways of setting the default attribute of the screen to be bold and have color-pair #1.

(attr-set! win (logior A_BOLD (color-pair 1)))

or

(attr-set! win A_BOLD 1)

[Procedure] pair-number attrs

This function is the inverse operation of color-pair. It is rarely useful.

     (color-pair 1) ==> 256
     (pair-number 256) ==> 1

Name  Description
A NORMAL  Normal display (no highlight)
A STANDOUT  Best highlighting mode of the terminal.
A UNDERLINE  Underlining
A REVERSE  Reverse video
A BLINK  Blinking
A DIM  Half bright
A BOLD  Extra bright or bold
A PROTECT  Protected mode
A INVIS  Invisible or blank mode
A ALTCHARSET Alternate character set
A CHARTEXT  Bit-mask to extract a character

Table 5.2: Attributes

5.2.9 Bell and screen flash routines

[Procedure] beep [Procedure] flash

The beep and flash routines are used to alert the terminal user. The routine beep sounds an audible alarm on the terminal, if possible; otherwise it flashes the screen (visible bell). The routine flash flashes the screen, and if that is not possible, sounds the alert. If neither alert is possible, nothing happens. Nearly all terminals have an audible alert (bell or beep), but only some can flash the screen.

These routines return #t if they succeed in beeping or flashing, #f otherwise.

5.2.10 Window background manipulation routines

[Procedure] bkgdset! win ch

The bkgdset! routine manipulates the background of the named window. The window background is set to the rendered complex character ch. The attribute part of the ch is combined (logior’d) with all non-blank characters that are written into the window with addch. Both the character and attribute parts of the ch are combined with the blank characters. The background becomes a property of the character and moves with the character through any scrolling and insert/delete line/character operations.

To the extent possible on a particular terminal, the attribute part of the background is displayed as the graphic rendition of the character put on the screen.

The return value is undefined.

[Procedure] bkgd win ch The bkgd function sets the background property of the current or specified window and then applies this setting to every character position in that window: The rendition of every character on the screen is changed to the new background rendition. Wherever the former background character appears, it is changed to the new background character.

The return value is undefined.

[Procedure] getbkgd win

The getbkgd function returns the given window’s current background rendered character.

5.2.11 Borders and lines

[Procedure] border win ls rs ts bs tl tr bl br The border routine draws a box around the edges of a window. Other than win, each argument is a rendered character, representing a side or a corner of the border. The arguments are ls - left side, rs - right side, ts - top side, bs - bottom side, tl - top left-hand corner, tr - top right-hand corner, bl - bottom left-hand corner, and br - bottom right-hand corner.

If any of these arguments is zero, then the corresponding default values are used instead. The default values are the line drawing characters that create a box, and

they would be ls = acs-vline, rs = acs-vline, ts = acs-hline, bs = acs-hline,
tl = acs-ulcorner, tr = acs-urcorner, bl = acs-llcorner, br = acs-lrcorner.

The characters used should have the standard character width. Double-width characters should not be used in borders.

It returns #f on failure and #t on success.

[Procedure] box win verch horch box is a shorthand for the following call:

     (border win verch verch horch horch 0 0 0 0)

[Procedure] hline win ch n #:key y x The hline procedure draws a horizontal (left to right) line using ch starting at the current cursor position in the window. The current cursor position is not changed. The line is at most n characters long, or as many as fit into the window.

The optional parameters y and x cause the cursor to be moved to that position before drawing the line.

The return value is unspecified.

[Procedure] vline win ch n #:key y x The vline procedure draws a vertical (top to bottom) line using ch starting at the current cursor position in the window. The current cursor position is not changed. The line is at most n characters long, or as many as fit into the window. If the key parameters y and x are set, it moves the cursor before drawing the line. The return value is unspecified.

5.2.12 Clearing windows and parts of window

The following family of functions clear all or part of a window. When clearing a window, the blanks created by erasure have the current background rendition (as set by bkgdset!) merged into them.

[Procedure] erase win

The erase routine copies blanks to every position in the window, clearing the screen. It returns #f on failure and #t on success.

[Procedure] clear win

The clear routine is like erase, but they also call clearok!, so that the screen is cleared completely on the next call to refresh for that window and repainted from scratch. Its return value is unspecified.

[Procedure] clrtobot win

The clrtobot routine erase from the cursor to the end of screen. That is, they erase all lines below the cursor in the window. Also, the current line to the right of the cursor, inclusive, is erased. Its return value is unspecified.

[Procedure] clrtoeol win

The clrtoeol routine erase the current line to the right of the cursor, inclusive, to the end of the current line. It returns #f on failure and #t on success. Failure could occur if the cursor position is offscreen.

5.2.13 Color manipulation routines

Curses supports color attributes on terminals with that capability. To use these routines start-color! must be called, usually right after initscr. Colors are always used in pairs (referred to as color-pairs). A color-pair consists of a foreground color (for characters) and a background color (for the blank field on which the characters are displayed). A programmer initializes a color-pair with the routine init-pair!. After it has been initialized, colorpair can be used as a new video attribute.

If a terminal is capable of redefining colors, the programmer can use the routine initcolor! to change the definition of a color. The routines has-colors? and can-change-color? return #t or #f, depending on whether the terminal has color capabilities and whether the programmer can change the colors. The routine color-content allows a programmer to extract the amounts of red, green, and blue components in an initialized color. The routine pair-content allows a programmer to find out how a given color-pair is currently defined.

[Procedure] start-color!

The start-color! routine requires no arguments. It must be called if the programmer wants to use colors, and before any other color manipulation routine is called. It is good practice to call this routine right after initscr. start-color! initializes eight basic colors (black, red, green, yellow, blue, magenta, cyan, and white), and prepares the procedures colors and color-pairs (respectively defining the maximum number of colors and color-pairs the terminal can support). It also restores the colors on the terminal to the values they had when the terminal was just turned on. The return value is unspecified.

[Procedure] colors

Returns the maximum number of colors the terminal can support.

[Procedure] color-pairs

Returns the maximum number of color-pairs the terminal can support.

[Constant] COLOR_BLACK

[Constant] COLOR_RED

[Constant] COLOR_GREEN
[Constant] COLOR_YELLOW
[Constant] COLOR_BLUE
[Constant] COLOR_MAGENTA
[Constant] COLOR_CYAN
[Constant] COLOR_WHITE

These variables contain the color number of the eight default colors.

[Procedure] init-color! color red green blue The init-color! routine changes the definition of a color. It takes four arguments: the number of the color to be changed followed by three RGB values (for the amounts of red, green, and blue components). The value of the first argument must be between 0 and the value returned by the colors procedure. Each of the last three arguments must be a value between 0 and 1000. When init-color! is used, all occurrences of that color on the screen immediately change to the new definition. It returns #f on failure and #t on success. Failure may indicate that, for the current terminal, the colors cannot be modified.

[Procedure] init-pair! color-pair fore-color back-color The init-pair! routine changes the definition of a color-pair. It takes three arguments: the number of the color-pair to be changed, the fore-ground color number, and the background color number. The value of the first argument must be between 1 and color-pairs - 1. The value of the second and third arguments must be between 0 and colors. Color pair 0 is assumed to be white on black, but is actually whatever the terminal implements before color is initialized. If the color-pair was previously initialized, the screen is refreshed and all occurrences of that color-pair are changed to the new definition. Color pair 0 is set via the assume-default-colors routine. After use-default-colors has been called, the special color number -1 can be used as fore-color or back-color. Color number -1 is set to contain the default foreground or background color for this screen, which is usually white on black. It returns #f on failure and #t on success. Failure may indicate an attempt to change a color pair beyond the allow range of color pairs.

[Procedure] has-colors?

The has-colors? routine returns #t if the terminal can manipulate colors; otherwise, it returns #f. This routine facilitates writing terminal-independent programs. For example, a programmer can use it to decide whether to use color or some other video attribute.

[Procedure] can-change-color?

The can-change-color? routine returns #t if the terminal supports colors and can change their definitions; otherwise, it returns #f.

[Procedure] color-content color

The color-content routine returns the intensity of the red, green, and blue (RGB) components in a color. Given color as its argument, it returns a list of three integers indicating the red, green, and blue components in the given color. The value of the color argument must be between 0 and the value returned by the colors procedure. The values returned are between 0 (no component) and 1000 (maximum amount of component).

[Procedure] pair-content pair

The pair-content routine allows programmers to find out what colors color-pair pair consists of. The value of pair must be between 1 and color-pairs - 1. It returns a list of two elements: the foreground color number and the background color number. The values are between 0 and the value returned by colors.

[Procedure] use-default-colors

This procedure allow routines that use color numbers, like init-pair! to use the special color number -1 to indicate the default color. When -1 is used in lieu of the foreground color, it indicated the default foreground color. When -1 is used as the background color, it indicated the default background color. Thus, it has different meanings depending on the context. It returns #t if successful or #f is the terminal lacks the capability to set color pair 0.

[Procedure] assume-default-colors fg bg

This procedure allow the modification of special color pair 0, which cannot be modified by init-pair!. Color pair 0 is the default foreground and background color for the window. It returns #t if successful or #f is the terminal lacks the capability to set color pair 0.

5.2.14 Deleting the character under the cursor

[Procedure] delch win #:key y x This routine deletes the character under the cursor; all characters to the right of the cursor on the same line are moved to the left one position and the last character on the line is filled with a blank. The cursor position does not change (after moving to y, x, if specified). (This does not imply use of the hardware delete character feature.) It returns #f on failure and #t on success.

5.2.15 Deleting and inserting lines

[Procedure] deleteln win

The deleteln procedure deletes the line under the cursor in the window; all lines below the current line are moved up one line. The bottom line of the window is cleared. The cursor position does not change. It returns #f on failure and #t on success.

[Procedure] insdelln win n

The insdelln routine, for positive n, inserts n lines into the specified window above the current line. The n bottom lines are lost. For negative n, it deletes n lines (starting with the one under the cursor), and move the remaining lines up. The bottom n lines are cleared. The current cursor position remains the same. It returns #f on failure and #t on success.

[Procedure] insertln win

The insertln routine inserts a blank line above the current line and the bottom line is lost. It returns #f on failure and #t on success.

5.2.16 Getting characters from the keyboard

[Procedure] getch win #:key y x The getch routine reads a character from the terminal associated with the given window. In no-delay mode, if no input is waiting, the value #f is returned. In delay mode, the program waits until the system passes text through to the program. Depending on the setting of cbreak, this is after one character (cbreak mode), or after the first newline (nocbreak! mode). In halfdelay! mode, the program waits until a character is typed or the specified timeout has been reached. Unless noecho! has been set, then the character will also be echoed into the designated window according to the following rules: If the character is the current erase character, left arrow, or backspace, the cursor is moved one space to the left and that screen position is erased as if delch had been called. If the character value is any other KEY_ define, the user is alerted with a beep call. Otherwise the character is simply output to the screen. If keypad! is #t, and a function key is pressed, the integer keycode for that function key is returned instead of the raw characters. There is a list of possible function keys. See Table 5.3. When a character that could be the beginning of a function key is received (which, on modern terminals, means an escape character), curses sets a timer. If the remainder of the sequence does not come in within the designated time, the character is passed through; otherwise, the function key value is returned. For this reason, many terminals experience a delay between the time a user presses the escape key and the escape is returned to the program.

If the window is not a pad, and it has been moved or modified since the last call to refresh, refresh will be called before another character is read.

[Procedure] ungetch ch

The ungetch routine places ch back onto the input queue to be returned by the next call to getch. There is just one input queue for all windows.

It returns #f if there is no more space in the buffer, or #t otherwise.

[Procedure] key-f n

This procedure returns the keycode for function key n.

The following function keys, might be returned by getch if keypad! has been enabled.

Note that almost all of these function keys do not exist on modern keyboards. The standard PC keyboard cannot be depended upon to have more than (key-f 1) through

(key-f 12), KEY_PPAGE (Page Up), KEY_NPAGE (Page Down), KEY_HOME, KEY_END, KEY_IC
(Insert), KEY_DC (Delete), KEY_BACKSPACE, KEY_DC (Delete), KEY_UP, KEY_DOWN, KEY_LEFT,
and KEY_RIGHT.

Also, a common terminal is more likely to return C-M than KEY_ENTER when the RET key is pressed.

name  description

KEY_BREAK  Break key

KEY_DOWN  Arrow down
KEY_UP  Arrow up
KEY_LEFT  Arrow left
KEY_RIGHT  Arrow right

KEY_HOME  Home key
KEY_BACKSPACE  Backspace

KEY_F0  Function key zero
KEY_DL  Delete line
KEY_IL  Insert line
KEY_DC  Delete character
KEY_IC  Insert char or enter insert mode
KEY_EIC  Exit insert char mode
KEY_CLEAR  Clear screen
KEY_EOS  Clear to end of screen
KEY_EOL  Clear to end of line
KEY_SF  Scroll 1 line forward
KEY_SR  Scroll 1 line backward (reverse)

KEY_NPAGE  Next page

KEY_PPAGE  Previous page

KEY_STAB  Set tab
KEY_CTAB  Clear tab

KEY_CATAB  Clear all tabs
KEY_ENTER  Enter or send
KEY_SRESET  Soft (partial) reset
KEY_RESET  Reset or hard reset
KEY_PRINT  Print or copy
KEY_LL  Home down or bottom (lower left)
KEY_A1  Upper left of keypad
KEY_A3  Upper right of keypad
KEY_B2  Center of keypad
KEY_C1  Lower left of keypad
KEY_C3  Lower right of keypad
KEY_BTAB  Back tab key
KEY_BEG  Beg(inning) key

KEY_CANCEL  Cancel key
KEY_CLOSE  Close key

KEY_COMMAND  Cmd (command) key

KEY_COPY  Copy key
KEY_CREATE  Create key
KEY_END  End key
KEY_EXIT  Exit key
KEY_FIND  Find key
KEY_HELP  Help key
KEY_MARK  Mark key
KEY_MESSAGE  Message key

Table 5.3: the keypad constants, part 1

name  description
KEY_MOUSE  Mouse event read

KEY_MOVE  Move key

KEY_NEXT  Next object key

KEY_OPEN  Open key
KEY_OPTIONS  Options key

KEY_PREVIOUS  Previous object key

KEY_REDO  Redo key

KEY_REFERENCE  Ref(erence) key

KEY_REFRESH  Refresh key
KEY_REPLACE  Replace key

KEY_RESIZE  Screen resized

KEY_RESTART  Restart key
KEY_RESUME  Resume key
KEY_SAVE  Save key

KEY_SBEG  Shifted beginning key
KEY_SCANCEL  Shifted cancel key
KEY_SCOMMAND  Shifted command key
KEY_SCOPY  Shifted copy key
KEY_SCREATE  Shifted create key
KEY_SDC  Shifted delete char key
KEY_SDL  Shifted delete line key

KEY_SELECT  Select key

KEY_SEND  Shifted end key
KEY_SEOL  Shifted clear line key
KEY_SEXIT  Shifted exit key
KEY_SFIND  Shifted find key
KEY_SHELP  Shifted help key
KEY_SHOME  Shifted home key
KEY_SIC  Shifted input key
KEY_SLEFT  Shifted left arrow key
KEY_SMESSAGE  Shifted message key
KEY_SMOVE  Shifted move key
KEY_SNEXT  Shifted next key
KEY_SOPTIONS  Shifted options key
KEY_SPREVIOUS  Shifted prev key
KEY_SPRINT  Shifted print key
KEY_SREDO  Shifted redo key
KEY_SREPLACE  Shifted replace key
KEY_SRIGHT  Shifted right arrow
KEY_SRESUME  Shifted resume key
KEY_SSAVE  Shifted save key
KEY_SSUSPEND  Shifted suspend key
KEY_SUNDO  Shifted undo key

KEY_SUSPEND  Suspend key
KEY_UNDO  Undo key

Table 5.4: the keypad constants, part 2

[Procedure] has-key? ch

The has-key? routine takes a key value ch from the above list, and returns #t or #f according to whether the current terminal type recognizes a key with that value. Note that a few values do not correspond to a real key, e.g., KEY_RESIZE and KEY_MOUSE. See resizeterm for more details about KEY_RESIZE. For a discussion of KEY_MOUSE see Section 5.2.27 [Mouse handling], page 82.

Use of the ESC key by a programmer for a single character function is discouraged, as it will cause a delay of up to one second while the keypad code looks for a following function-key sequence.

Note that some keys may be the same as commonly used control keys, e.g., KEY ENTER versus C-M, KEY BACKSPACE versus C-H. Some curses implementations may differ according to whether they treat these control keys specially (and ignore the terminfo), or use the terminfo definitions. Ncurses uses the terminfo definition. If it says that KEY ENTER is C-M, getch will return KEY_ENTER when you press C-M.

   When using getch, wgetch, mvgetch, or mvwgetch, nocbreak mode (nocbreak!) and

echo mode (echo!) should not be used at the same time. Depending on the state of the TTY driver when each character is typed, the program may produce undesirable results.

Historically, the set of keypad macros was largely defined by the extremely function-key-rich keyboard of the AT&T 7300, aka 3B1, aka Safari 4. Modern personal computers usually have only a small subset of these. IBM PC-style consoles typically support little more than

KEY_UP, KEY_DOWN, KEY_LEFT, KEY_RIGHT, KEY_HOME, KEY_END, KEY_NPAGE, KEY_PPAGE, and

function keys 1 through 12. The ins key is usually mapped to KEY_IC

5.2.17 Handling unmapped keys

For most every modern-day terminal, you’ll find that if you get the TERM environment variable set correctly, then each of the function keys on the keyboard will be interpreted as a key code.

But in rare circumstances, you may find that a function key on your keyboard is not being mapped to a keycode. As an example, on some keyboards there is a special “menu” key in between Ctrl and Alt. On my keyboard, pressing “menu” returns “ESC [ 2 9 ~”, which getch returns as a that 5 separate characters, instead of a single keycode like

KEY_MENU.

When this happens, the problem is most often that you’ve set your TERM incorrectly. Next most likely is that your terminfo or termcap database is out of date. That is where you should first seek your solution.

But, if that fails, and you need a quick-and-dirty workaround, define-key can help. It lets you map an escape sequence to a key code.

Continuing my example, if I choose to map this menu key to a keycode, I can choose the keycode of a key that that doesn’t appear on my keyboard, like KEY_FIND, and associate with that string, using define-key.

[Procedure] define-key defn keycode This procedure defines a new, custom keycode. When the string in defn is input, routines like getch will return the keycode instead, if the keypad is on. If defn is an empty string, the keycode will be cleared.

If keycode is an existing keycode, its defn replaces its previous definition.

The return value is #t on success.

For the example, I can use the command

 (define-key (string #\esc #\[ #\2 #\9 #\~) KEY_FIND)

From that point on, when the Menu key is pressed, getch will return the integer KEY_

FIND.

[Procedure] key-defined defn

If the string defn is a character sequence that is bound to a keycode, that keycode is returned. Otherwise #f is returned.

5.2.18 Receiving strings from the keyboard

[Procedure] getnstr win n #:key y x The procedure getnstr is equivalent to a series of calls to getch, until a newline or carriage return is received (the terminating character is not included in the returned string). The resulting string is returned.

getnstr reads at most n characters, thus preventing a possible overflow of the input buffer. Any attempt to enter more characters (other than the terminating newline or carriage return) causes a beep. Function keys also cause a beep and are ignored.

The user’s erase and kill characters are interpreted. If keypad mode is on for the window, KEY_LEFT and KEY_BACKSPACE are both considered equivalent to the user’s kill character.

Characters input are echoed only if echo! is currently on. In that case, backspace is echoed as deletion of the previous character (typically a left motion).

If a SIGWINCH interrupts the function, it will return the integer value of KEY RESIZE instead of a string.

5.2.19 Cursor location and window coordinates

[Procedure] getyx win [Procedure] getcurx win [Procedure] getcury win

The getyx procedure returns the current cursor position as a list containing two integers. The first is the y position and the second is the x position.

The other two procedures return the x y positions respectively.

[Procedure] getparyx win [Procedure] getparx win [Procedure] getpary win

If win is a subwindow, the getparyx procedure places the beginning coordinates of the subwindow relative to the parent window into two element list (y x). Otherwise, (-1 -1) is returned.

The other two procedures return the x and y respectively.

[Procedure] getbegyx win [Procedure] getbegx win [Procedure] getbegy win

The procedure getbegyx returns the beginning coordinates of a window as two element list (y x).

[Procedure] getmaxyx win [Procedure] getmaxx win [Procedure] getmaxy win

The procedure getmaxyx returns the size of a window as two element list (y x).

5.2.20 Getting a rendered character from the window

These functions examine a character on the screen and return it, along with its rendition.

[Procedure] inch

These routines return the rendered complex character at the current position in the named window.

5.2.21 Getting a string of characters and attributes from a window

[Procedure] inchnstr win #:key y x n These routines return a list of rendered characters, starting at the current cursor position in the named window and ending at the right margin of the window or until n characters have been found.

5.2.22 Input options

[Procedure] cbreak! [Procedure] nocbreak!

Normally, the TTY driver buffers typed characters until a NL or RET is typed. The cbreak! routine disables line buffering and erase/kill character processing (interrupt and flow control characters are unaffected), making characters typed by the user immediately available to the program. The nocbreak! routine returns the terminal to normal (cooked) mode. These routines will return #t on success and #f on failure. Failure may indicate that the underlying screen data is invalid. The routine may also fail if cbreak! was called on an ordinary file (such as might be used with newterm) instead of a TTY. Initially the terminal may or may not be in cbreak mode, as the mode is inherited; therefore, a program should call cbreak! or nocbreak! explicitly. Most interactive programs using curses set the cbreak! mode. Note that cbreak! overrides raw!.

[Procedure] echo! [Procedure] noecho!

The echo! and noecho! routines control whether characters typed by the user are echoed by getch as they are typed. Echoing by the TTY driver is always disabled, but initially getch is in echo mode, so characters typed are echoed. Authors of most interactive programs prefer to do their own echoing in a controlled area of the screen, or not to echo at all, so they disable echoing by calling noecho!. See Section 5.2.16 [Getting characters from the keyboard], page 72 for a discussion of how these routines interact with echo! and noecho!. The return values are unspecified.

[Procedure] halfdelay! tenths

The halfdelay! routine is used for half-delay mode, which is similar to cbreak! mode in that characters typed by the user are immediately available to the program. However, after blocking for tenths tenths of seconds, #f is returned if nothing has been typed. The value of tenths must be a number between 1 and 255. Use nocbreak! to leave half-delay mode. The return value is unspecified.

[Procedure] intrflush! bf

If the intrflush option is enabled, (bf is #t), when an interrupt key is pressed on the keyboard (interrupt, break, quit) all output in the TTY driver queue will be flushed, giving the effect of faster response to the interrupt, but causing curses to have the wrong idea of what is on the screen. Disabling (bf is #f), the option prevents the flush. The default for the option is inherited from the TTY driver settings. The return value is unspecified.

[Procedure] keypad! win bf

The keypad! option enables the keypad of the user’s terminal. If enabled (bf is #t), the user can press a function key (such as an arrow key) and getch returns a single value representing the function key, as in KEY_LEFT. If disabled (bf is #f), curses does not treat function keys specially and the program has to interpret the escape sequences itself. If the keypad in the terminal can be turned on (made to transmit) and off (made to work locally), turning on this option causes the terminal keypad to be turned on when getch is called. The default value for keypad is #f. The return value is unspecified. This procedure could throw a “bad state” error.

[Procedure] meta! bf

If Guile-Ncurses was compiled with the standard version of the ncurses library, and not the wide version, it operates on 8-bit characters. Initially, whether the terminal returns 7 or 8 significant bits on input depends on the control mode of the TTY driver. To force 8 bits to be returned, invoke (meta! #t). This is equivalent, under POSIX, to setting the CS8 flag on the terminal. To force 7 bits to be returned, invoke (meta! #f). This is equivalent, under POSIX, to setting the CS7 flag on the terminal. The window argument, win, is always ignored. If the terminfo capabilities smm (meta-on) and rmm (meta-off) are defined for the terminal. The code for smm is sent to the terminal when (meta! #t) is called and rmm is sent when (meta! #f) is called. The return value is unspecified.

[Procedure] nodelay! win bf

The nodelay! option causes getch to be a non-blocking call. If no input is ready, getch returns #f. If disabled (bf is #f), getch waits until a key is pressed. The return value is unspecified.

[Procedure] notimeout! win bf

While interpreting an input escape sequence, getch sets a timer while waiting for the next character. If (notimeout! win #t) is called, then getch does not set a timer. The purpose of the timeout is to differentiate between sequences received from a function key and those typed by a user. The return value is unspecified.

[Procedure] raw! [Procedure] noraw!

The raw! and noraw! routines place the terminal into or out of raw mode. Raw mode is similar to cbreak! mode, in that characters typed are immediately passed through to the user program. The differences are that in raw mode, the interrupt, quit, suspend, and flow control characters are all passed through uninterpreted, instead of generating a signal. The behavior of the BREAK key depends on other bits in the TTY driver that are not set by curses. The return value is unspecified.

[Procedure] noqiflush! [Procedure] qiflush!

When the noqiflush! routine is used, normal flush of input and output queues associated with the INTR, QUIT and SUSP characters will not be done. When qiflush! is called, the queues will be flushed when these control characters are read. You may want to call noqiflush! in a signal handler if you want output to continue as though the interrupt had not occurred, after the handler exits. The return value is unspecified.

[Procedure] timeout! win delay

The timeout! routine sets blocking or non-blocking read for a given window. If delay is negative, blocking read is used (i.e., waits indefinitely for input). If delay is zero, then non-blocking read is used (i.e., read returns #f if no input is waiting). If delay is positive, then read blocks for delay milliseconds, and returns #f if there is still no input. Hence, these routines provide the same functionality as nodelay!, plus the additional capability of being able to block for only delay milliseconds (where delay is positive).

[Procedure] typeahead! fd

The curses library does “line-breakout optimization” by looking for typeahead periodically while updating the screen. If input is found, and it is coming from a TTY, the current update is postponed until refresh or doupdate is called again. This allows faster response to commands typed in advance. Normally, the input file port passed to newterm, or stdin in the case that initscr was used, will be used to do this typeahead checking. The typeahead! routine specifies that the (integer) file descriptor fd is to be used to check for typeahead instead. If fd is -1, then no typeahead checking is done. The routine returns #t if the mode could be set and #f on failure.

There are a set of procedures to test the input options of a given window.

[Procedure] is-keypad? win [Procedure] is-meta? win [Procedure] is-nodelay? win [Procedure] is-immedok? win [Procedure] is-notimeout? win

These test the input options of the window win and return #t if they are set.

5.2.23 Inserting a character before the cursor

[Procedure] insch win ch #:key y x These routines insert the character ch before the character under the cursor. All characters to the right of the cursor are moved one space to the right, with the possibility of the rightmost character on the line being lost. The insertion operation does not change the cursor position.

It returns #f upon failure and #t upon successful completion.

5.2.24 Inserting a string before the cursor

[Procedure] insstr win str #:key y x n This routine inserts a character string (as many characters as will fit on the line) before the character under the cursor. All characters to the right of the cursor are shifted right with the possibility of the rightmost characters on the line being lost. The cursor position does not change (after moving to y, x, if specified). If n is specified, it inserts a leading substring of at most n characters, or if n<=0, then the entire string is inserted.

Special characters are handled as in addch.

It returns #f upon failure and #t upon successful completion.

5.2.25 Getting a string of characters from the screen

[Procedure] instr win #:key y x n This routine returns a unrendered string, extracted starting at the current cursor position in the named window. Attributes are stripped from the characters. If n is specified, it returns a leading substring at most n characters long.

5.2.26 Low-level routines

The following routines give low-level access to various curses capabilities. Theses routines typically are used inside library routines.

[Procedure] def-prog-mode
[Procedure] def-shell-mode

The def-prog-mode and def-shell-mode routines save the current terminal modes as the "program" (in curses) or "shell" (not in curses) state for use by the resetprog-mode and reset-shell-mode routines. This is done automatically by initscr. There is one such save area for each screen context allocated by newterm.

They return #t on success and #f on failure.

[Procedure] reset-prog-mode
[Procedure] reset-shell-mode

The reset-prog-mode and reset-shell-mode routines restore the terminal to "program" (in curses) or "shell" (out of curses) state. These are done automatically by endwin and, after an endwin, by doupdate, so they normally are not called.

Returns #t on success or #f on failure. Failure could indicate that this terminal was created with newterm and thus doesn’t have a previous state.

[Procedure] resetty [Procedure] savetty

The resetty and savetty routines save and restore the state of the terminal modes. savetty saves the current state in a buffer and resetty restores the state to what it was at the last call to savetty.

Returns #t on success or #f on failure. Failure could indicate that this terminal was created with newterm and thus doesn’t have a previous state.

[Procedure] curs-set visibility The curs-set routine changes the visibility of the cursor. If visibility is 0, the cursor is set to invisible. For 1, the cursor is visible. For 2, the cursor is very visible. If the terminal supports the visibility requested, the previous cursor state is returned; otherwise, #f is returned.

[Procedure] napms ms

The procedure is used to sleep for ms milliseconds.

The return value is unspecified.

5.2.27 Mouse handling

These functions provide an interface to mouse events. Mouse events are represented by KEY_MOUSE pseudo-key values in the wgetch input stream.

To make mouse events visible, use the mousemask function.

[Procedure] has-mouse?

Returns #t if the mouse driver has been successfully installed.

[Procedure] mousemask mask

This will set the mouse events to be reported. By default, no mouse events are reported. The function will return a mask to indicate which of the specified mouse events can be reported; on complete failure it returns 0.

As a side effect, setting a zero mousemask may turn off the mouse pointer; setting a nonzero mask may turn it on. Whether this happens is device-dependent.

See Table 5.5 for the mouse event types mask that may be defined.

name  description
BUTTON1_PRESSED  mouse button 1 down

BUTTON1_RELEASED  mouse button 1 up

BUTTON1_CLICKED  mouse button 1 clicked
BUTTON1_DOUBLE_CLICKED  mouse button 1 double clicked
BUTTON1_TRIPLE_CLICKED  mouse button 1 triple clicked
BUTTON2_PRESSED  mouse button 2 down

BUTTON2_RELEASED  mouse button 2 up

BUTTON2_CLICKED  mouse button 2 clicked
BUTTON2_DOUBLE_CLICKED  mouse button 2 double clicked
BUTTON2_TRIPLE_CLICKED  mouse button 2 triple clicked
BUTTON3_PRESSED  mouse button 3 down

BUTTON3_RELEASED  mouse button 3 up

BUTTON3_CLICKED  mouse button 3 clicked
BUTTON3_DOUBLE_CLICKED  mouse button 3 double clicked
BUTTON3_TRIPLE_CLICKED  mouse button 3 triple clicked
BUTTON4_PRESSED  mouse button 4 down

BUTTON4_RELEASED  mouse button 4 up

BUTTON4_CLICKED  mouse button 4 clicked
BUTTON4_DOUBLE_CLICKED  mouse button 4 double clicked
BUTTON4_TRIPLE_CLICKED  mouse button 4 triple clicked
BUTTON5_PRESSED  mouse button 5 down

BUTTON5_RELEASED  mouse button 5 up

BUTTON5_CLICKED  mouse button 5 clicked
BUTTON5_DOUBLE_CLICKED  mouse button 5 double clicked
BUTTON5_TRIPLE_CLICKED  mouse button 5 triple clicked
BUTTON_SHIFT  shift was down during button state change
BUTTON_CTRL  control was down during button state change
BUTTON_ALT  alt was down during button state change
ALL_MOUSE_EVENTS  report all button state changes

REPORT_MOUSE_POSITION  report mouse movement

Table 5.5: Mouse mask constants

Once a class of mouse events have been made visible in a window, calling the wgetch function on that window may return KEY_MOUSE as an indicator that a mouse event has been queued. To read the event data and pop the event off the queue, call getmouse.

[Procedure] getmouse

This will return either a list of mouse information, or #f. If it does return a list, it will have the following form:

(id ; id to distinguish multiple devices x y z ; event coordinates bstate) ; button state bits When getmouse returns a list, the data deposited as y and x in the list will be screenrelative character-cell coordinates. The returned state mask will have exactly one bit set to indicate the event type.

[Procedure] ungetmouse mouse-event The ungetmouse function behaves analogously to ungetch. It pushes a KEY_MOUSE event onto the input queue, and associates with that event the given state data and screen-relative character-cell coordinates in the mouse-event list, where mouse-event is a list of five elements as described above.

[Procedure] mouse-trafo win y x to-screen The mouse-trafo function transforms a given pair of coordinates y, x from stdscr-relative coordinates to coordinates relative to the given window win or vice versa. Please remember, that stdscr-relative coordinates are not always identical to window-relative coordinates due to the mechanism to reserve lines on top or bottom of the screen for other purposes (see slk-init, for example). If the parameter to-screen is #t, the procedure returns either a list of two elements (y, x) which is the location inside the window win, or #f if the location was not inside the window. If to-screen is #f, the return a list of two elements of where the window-relative location y, x would be in stdscr-relative coordinates. The mouse-trafo procedure performs the same translation as wmouse-trafo, using stdscr for win.

[Procedure] wenclose? win y x

The wenclose? function tests whether a given pair of screen-relative character-cell coordinates is enclosed by the given window win, returning #t if it is and #f otherwise. It is useful for determining what subset of the screen windows enclose the location of a mouse event.

[Procedure] mouseinterval erval

The mouseinterval function sets the maximum time (in thousands of a second) that can elapse between press and release events for them to be recognized as a click. Use (mouseinterval 0) to disable click resolution. This function returns the previous interval value. Use (mouseinterval -1) to obtain the interval without altering it. The default is one sixth of a second.

5.2.28 Moving the cursor

[Procedure] move win y x This routine moves the cursor associated with the window to line y and column x. This routine does not move the physical cursor of the terminal until refresh is called. The position specified is relative to the upper left-hand corner of the window, which

is (0 0).

It returns #f upon failure and #t on successful completion.

5.2.29 Output options

These routines set options that change the style of output within curses. All options are initially #f, unless otherwise stated. It is not necessary to turn these options off before calling endwin.

[Procedure] clearok! win bf

If clearok! is called with bf as #t, the next call to refresh with this window will clear the screen completely and redraw the entire screen from scratch. This is useful when the contents of the screen are uncertain, or in some cases for a more pleasing visual effect. The return value is unspecified.

[Procedure] idlok! win bf

If idlok! is called with #t as second argument, curses considers using the hardware insert/delete line feature of terminals so equipped. Calling idlok! with bf as #f disables use of line insertion and deletion. This option should be enabled only if the application needs insert/delete line, for example, for a screen editor. It is disabled by default because insert/delete line tends to be visually annoying when used in applications where it isn’t really needed. If insert/delete line cannot be used, curses redraws the changed portions of all lines. The return value is unspecified.

[Procedure] idcok! win bf

If idcok! is called with bf as #f, curses no longer considers using the hardware insert/delete character feature of terminals so equipped. Use of character insert/delete is enabled by default. Calling idcok! with #t as second argument re-enables use of character insertion and deletion. The return value is unspecified.

[Procedure] immedok! win bf

If immedok! is called with #t as argument, any change in the window image, such as the ones caused by addch, clrtobot, scrl, etc., automatically cause a call to refresh. However, it may degrade performance considerably, due to repeated calls to refresh. It is disabled by default. The return value is unspecified.

[Procedure] leaveok! win bf

Normally, the hardware cursor is left at the location of the window cursor being refreshed. The leaveok! option allows the cursor to be left wherever the update happens to leave it. It is useful for applications where the cursor is not used, since it reduces the need for cursor motions.

[Procedure] setscrreg! win top bot

The setscrreg! routine allows the application programmer to set a software scrolling region in a window. top and bot are the line numbers of the top and bottom margin of the scrolling region. (Line 0 is the top line of the window.) If this option and scrollok! are enabled, an attempt to move off the bottom margin line causes all lines in the scrolling region to scroll one line in the direction of the first line. Only the text of the window is scrolled. (Note that this has nothing to do with the use of a physical scrolling region capability in the terminal, like that in the VT100. If idlok! is enabled and the terminal has either a scrolling region or insert/delete line capability, they will probably be used by the output routines.)

[Procedure] getscrreg win

Returns a two-element list that contains the line numbers of the top and bottom of the scroll region for the window win.

[Procedure] scrollok! win bf

The scrollok! option controls what happens when the cursor of a window is moved off the edge of the window or scrolling region, either as a result of a newline action on the bottom line, or typing the last character of the last line. If disabled, (bf is #f), the cursor is left on the bottom line. If enabled, (bf is #t), the window is scrolled up one line (Note that to get the physical scrolling effect on the terminal, it is also necessary to call idlok!). The return value is unspecified.

[Procedure] nl! [Procedure] nonl!

The nl! and nonl! routines control whether the underlying display device translates RET (return) into NL (newline) on input, and whether it translates NL into RET and LF (linefeed) output (in either case, the call (addch (normal #\nl)) does the equivalent of return and line feed on the virtual screen). Initially, these translations do occur. If you disable them using nonl!, curses will be able to make better use of the line-feed capability, resulting in faster cursor motion. Also, curses will then be able to detect the return key. The return values are unspecified;

There are a set of procedures to test the output options of a given window.

[Procedure] is-cleared? win [Procedure] is-idlok? win [Procedure] is-idcok? win [Procedure] is-immedok? win [Procedure] is-leaveok? win [Procedure] is-scrollok? win

These test the output options of the window win and return #t if they are set.

5.2.30 Overlay and manipulate overlapped windows

[Procedure] overlay srcwin dstwin [Procedure] overwrite srcwin dstwin

The overlay and overwrite routines overlay srcwin on top of dstwin. scrwin and dstwin are not required to be the same size; only text where the two windows overlap is copied. The difference is that overlay is non-destructive (blanks are not copied) whereas overwrite is destructive. The return values are unspecified.

[Procedure] copywin srcwin dstwin sminrow smincol dminrow dmincol dmaxrow

dmaxcol overlay The copywin routine provides a finer granularity of control over the overlay and overwrite routines. Like in the prefresh routine, a rectangle is specified in the destination window, (dminrow, dmincol) and (dmaxrow, dmaxcol), and the upperleft-corner coordinates of the source window, (sminrow, smincol). If the argument overlay is #t, then copying is non-destructive, as in overlay. Returns #t on success or #f on failure. #f could indicate that some part of the window would be placed offscreen.

5.2.31 Create and display pads

A pad is like a window, except that it is not restricted by the screen size, and is not necessarily associated with a particular part of the screen. Pads can be used when a large window is needed, and only a part of the window will be on the screen at one time. Automatic refreshes of pads (e.g., from scrolling or echoing of input) do not occur. It is not legal to call refresh with a pad as an argument; the routines prefresh or pnoutrefresh should be called instead. Note that these routines require additional parameters to specify the part of the pad to be displayed and the location on the screen to be used for the display.

[Procedure] newpad nlines ncols The newpad routine creates and returns #<window> pad with the given number of lines, nlines, and columns, ncols.

[Procedure] is-pad? win

Returns #t if win is a pad. #f otherwise. If the underlying ncurses implementation is not capable of reporting whether a window is a pad, this function will always return #t. This can happen in older versions of ncurses that were compiled with the NCURSES_OPAQUE option enabled. To see if this is-pad? procedure actually works, you can check the constant %ispad-broken, which will be #f is is-pad? actually works.

[Procedure] subpad orig nlines ncols begin-y begin-x The subpad routine creates and returns a pointer to a subwindow within a pad with the given number of lines, nlines, and columns, ncols. Unlike subwin, which uses screen coordinates, the window is at position (begin-x, begin-y) on the pad. The window is made in the middle of the window orig, so that changes made to one window affect both windows. During the use of this routine, it will often be necessary to call touchwin or touchline on orig before calling prefresh.

[Procedure] prefresh pad pminrow pmincol sminrow simincol smaxrow smaxcol [Procedure] pnoutrefresh pad pminrow pmincol sminrow smincol smaxrow smaxcol

The prefresh and pnoutrefresh routines are analogous to refresh and noutrefresh except that they relate to pads instead of windows. The additional parameters are needed to indicate what part of the pad and screen are involved. pminrow and pmincol specify the upper left-hand corner of the rectangle to be displayed in the pad. sminrow, smincol, smaxrow, and smaxcol specify the edges of the rectangle to be displayed on the screen. The lower right-hand corner of the rectangle to be displayed in the pad is calculated from the screen coordinates, since the rectangles must be the same size. Both rectangles must be entirely contained within their respective structures. Negative values of pminrow, pmincol, sminrow, or smincol are treated as if they were zero. Returns #f upon failure and #t upon successful

[Procedure] pechochar pad ch

The pechochar routine is functionally equivalent to a call to addch followed by a call to refresh or a call to addch followed by a call to prefresh. The knowledge that only a single character is being output is taken into consideration and, for noncontrol characters, a considerable performance gain might be seen by using these routines instead of their equivalents. In the case of pechochar, the last location of the pad on the screen is reused for the arguments to prefresh. Returns #f upon failure and #t upon successful completion.

5.2.32 Print data to a terminal-attached ports

Some old-school hardware terminals had serial or parallel ports attached to them. There is a procedure to send data to those ports. Usually those ports were used for old serial dot-matrix printers. This function is likely obsolete.

[Procedure] mcprint str

This procedure takes the given str, converts it into a byte stream using the current locale, and then sends the data to a terminal’s printer port. A terminal that has a printer port must terminfo information that describes it, such as mc5p or mc4 capabilities. Explaining this is beyond the scope of this library. Those hardware terminal ports and the printers that used them were slow, and the application must take the responsibility to not flood the port. The procedure returns #f if the write fails. Otherwise, it returns the number of bytes actually printed.

5.2.33 Refresh windows and lines

[Procedure] refresh win

The refresh routine (or noutrefresh and doupdate) must be called to get actual output to the terminal, as other routines merely manipulate data structures. The routine refresh copies the named window to the physical terminal screen, taking into account what is already there to do optimizations. The refresh routine is the same, using stdscr as the default window. Unless leaveok! has been enabled, the physical cursor of the terminal is left at the location of the cursor for that window.

[Procedure] noutrefresh win

Copies the named window win to the virtual screen

[Procedure] doupdate

Compares the virtual screen to the physical screen and updates it where necessary.

The noutrefresh and doupdate routines allow multiple updates with more efficiency than refresh alone. In addition to all the window structures, curses keeps two data structures representing the terminal screen: a physical screen, describing what is actually on the screen, and a virtual screen, describing what the programmer wants to have on the screen.

The routine refresh works by first calling noutrefresh, which copies the named window to the virtual screen, and then calling doupdate, which compares the virtual screen to the physical screen and does the actual update. If the programmer wishes to output several windows at once, a series of calls to refresh results in alternating calls to noutrefresh and doupdate, causing several bursts of output to the screen. By first calling noutrefresh for each window, it is then possible to call doupdate once, resulting in only one burst of output, with fewer total characters transmitted and less CPU time used. If the win argument to refresh is the global variable curscr, the screen is immediately cleared and repainted from scratch.

The phrase "copies the named window to the virtual screen" above is ambiguous. What actually happens is that all touched (changed) lines in the window are copied to the virtual screen. This affects programs that use overlapping windows; it means that if two windows overlap, you can refresh them in either order and the overlap region will be modified only when it is explicitly changed.

[Procedure] redrawwin win

[Procedure] redrawln win beg-line num-lines The redrawln routine indicates to curses that some screen lines are corrupted and should be thrown away before anything is written over them. It touches the indicated lines (marking them changed). The routine redrawwin touches the entire window. Returns #t on success and #f on failure.

5.2.34 Reading and writing a curses screen to a file

[Procedure] scr-dump filename

The scr-dump routine dumps the current contents of the virtual screen to the file filename.

[Procedure] scr-restore filename

The scr-restore routine sets the virtual screen to the contents of filename, which must have been written using scr-dump. The next call to doupdate restores the screen to the way it looked in the dump file.

[Procedure] scr-init filename

The scr-init routine reads in the contents of filename and uses them to initialize the curses data structures about what the terminal currently has on its screen. If the data is determined to be valid, curses bases its next update of the screen on this information rather than clearing the screen and starting from scratch. scr-init is used after initscr or a system call to share the screen with another process which has done a scr-dump after its endwin call. The data is declared invalid if the terminfo capabilities rmcup and nrrmc exist; also if the terminal has been written to since the preceding scr-dump call.

[Procedure] scr-set filename

The scr-set routine is a combination of scr-restore and scr-init. It tells the program that the information in filename is what is currently on the screen, and also what the program wants on the screen. This can be thought of as a screen inheritance function.

To read (write) the contents of a window from (to) a port, use the getwin and putwin routines.

All routines return the integer #f upon failure and #t upon success.

5.2.35 Scroll a window

[Procedure] scroll win

The scroll routine scrolls the window up one line. This involves moving the lines in the window data structure. As an optimization, if the scrolling region of the window is the entire screen, the physical screen may be scrolled at the same time.

[Procedure] scrl win n

For positive n, the scrl and wscrl routines scroll the window up n lines (line i+n becomes i); otherwise scroll the window down n lines. This involves moving the lines in the window character image structure. The current cursor position is not changed. For these functions to work, scrolling must be enabled via scrollok!. It returns #f upon failure, and #t upon successful completion.

5.2.36 Terminal attributes

[Procedure] baudrate

The baudrate routine returns the output speed of the terminal. The number returned is in bits per second, for example 9600, and is an integer. It could return #f if this screen is no longer valid.

[Procedure] erasechar

The erasechar routine returns the user’s current erase character. If the terminal is a dumb terminal and has no erase character, it will return #f.

[Procedure] has-ic?

The has-ic? routine is true if the terminal has insert- and delete- character capabilities.

[Procedure] has-il?

The has-il? routine is true if the terminal has insert- and delete-line capabilities, or can simulate them using scrolling regions. This might be used to determine if it would be appropriate to turn on physical scrolling using scrollok!.

[Procedure] killchar

The killchar routine returns the user’s current line kill character. If the terminal has no killchar, it will return #f.

[Procedure] longname

The longname routine returns a string that is a verbose description of the current terminal.

[Procedure] termattrs

The termattrs function return a logior of all video attributes supported by the terminal using the standard A_XXX constants. The term-attrs function is nearly identical, except that it tests rarely used attributes that might someday having to

do with wide characters: A_HORIZONTAL, A_LEFT, A_LOW, A_RIGHT, A_TOP, and A_
VERTICAL.

[Procedure] termname

The termname routine returns the terminal name used by setupterm.

5.2.37 Refresh routines

[Procedure] touchwin win

[Procedure] touchline win start count The touchwin and touchline routines throw away all optimization information about which parts of the window have been touched, by pretending that the entire window has been drawn on. This is sometimes necessary when using overlapping windows, since a change to one window affects the other window, but the records of which lines have been changed in the other window do not reflect the change. The routine touchline only pretends that count lines have been changed, beginning with line start.

[Procedure] is-linetouched? win line

Returns #t if the specified window and line was modified since the last call to refresh.

[Procedure] is-wintouched? win

Returns #t if the specified window was modified since the last call to refresh>

[Procedure] untouchwin win

The untouchwin routine marks all lines in the window as unchanged since the last

call to refresh.

[Procedure] wtouchln win y n changed The wtouchln routine makes n lines in the window, starting at line y, look as if they have (changed=1) or have not (changed=0) been changed since the last call to

refresh.

5.2.38 Miscellaneous utilities

[Procedure] unctrl ch

The unctrl routine returns a character string which is a printable representation of the character ch, ignoring attributes. ch is either a simple Guile character or an xchar. Control characters are displayed in the ^X notation. Printing characters are displayed as is. Compare this to keyname.

[Procedure] keyname ch

The keyname routine returns a character string corresponding to the key ch. ch is either a simple Guile character or an integer key constant like KEY_HOME, and can not be an xchar. Control characters are displayed in the ^X notation. Values above 128 are either meta characters, shown in the M-X notation, or the names of function keys, or null.

[Procedure] use-env f

The use-env routine, if used, is called before initscr or newterm are called. When called with #f as an argument, the values of lines and columns specified in the terminfo database will be used, even if environment variables LINES and COLUMNS (used by default) are set, or if curses is running in a window (in which case default behavior would be to use the window size if LINES and COLUMNS are not set). Note that setting LINES or COLUMNS overrides the corresponding size which may be obtained from the operating system.

[Procedure] putwin win output-port The putwin routine writes all data associated with window win into an output port, such as a file or string port. This information can be later retrieved using the getwin function. The storage format is binary and is not described.

[Procedure] getwin input-port The getwin routine reads window related data from an input port. The input port should be backed by a file or string created by putwin. The routine then creates and initializes a new window using that data. It returns a new #<window> The storage format is binary and is not described.

[Procedure] delay-output ms

The delay-output routine inserts an ms millisecond pause in output. This routine should not be used extensively because padding characters are used rather than a CPU pause. The return value is unspecified.

[Procedure] flushinp

The flushinp routine throws away any typeahead that has been typed by the user and has not yet been read by the program. The return value is unspecified.

[Procedure] curses-version

Returns a string that indicates the version of ncurses being used, for example “ncurses 5.9”.

[Procedure] %filter

This procedure, when called before initscr, restricts the activity of curses to a single line, instead of the to the entire screen.

5.2.39 Window creation

[Procedure] newwin nlines ncols begin-y begin-x Calling newwin creates and returns a pointer to a new window with the given number of lines and columns. The upper left-hand corner of the window is at line begin-y, column begin-x. If either nlines or ncols is zero, they default to (- (lines) begin-y) and (- (cols) begin-x). A new full-screen window is created by calling (newwin 0

0 0 0).

[Procedure] delwin win

Calling delwin deletes the named window, freeing all memory associated with it (it does not actually erase the window’s screen image). Subwindows must be deleted before the main window can be deleted. This function is called implicitly if a window is garbage collected.

[Procedure] mvwin win y x

Calling mvwin moves the window so that the upper left-hand corner is at position x, y. If the move would cause the window to be off the screen, it is an error and the window is not moved. Moving subwindows is allowed, but should be avoided.

The return value is unspecified;

[Procedure] subwin orig nlines ncols begin-y begin-x Calling subwin creates and returns a pointer to a new window with the given number of lines, nlines, and columns, ncols. The window is at position (begin-y, begin-x) on the screen. (This position is relative to the screen, and not to the window orig.) The window is made in the middle of the window orig, so that changes made to one window will affect both windows. The subwindow shares memory with the window orig. When using this routine, it is necessary to call touchwin or touchline on orig before calling refresh on the subwindow.

[Procedure] is-subwin? win

Returns #t if win is a subwin. Otherwise, #f. If the underlying ncurses implementation is not capable of reporting whether a window is a subwindow, this function will always return #t. This can happen in older versions of ncurses that were compiled with the NCURSES_OPAQUE option enabled. To see if this is-subwin? procedure actually works, you can check the constant %issubwin-broken, which will be #f is is-subwin? actually works.

[Procedure] derwin orig nlines ncols begin-y begin-x Calling derwin is the same as calling subwin, except that begin-y and begin-x are relative to the origin of the window orig rather than the screen. There is no difference between the subwindows and the derived windows. It returns a window that shares memory with orig, or #f if the window could not be created.

[Procedure] mvderwin win par-y par-x Calling mvderwin moves a derived window (or subwindow) inside its parent window. The screen-relative parameters of the window are not changed. This routine is used to display different parts of the parent window at the same physical position on the screen. The return value is unspecified.

[Procedure] dupwin win

Calling dupwin returns a new window that is an exact duplicate of the window win.

[Procedure] syncup win

Calling syncup touches all locations in ancestors of win that are changed in win. If syncok! is called with second argument #t then syncup is called automatically whenever there is a change in the window.

[Procedure] syncok! win bf

If syncok! is called with the second argument #t then syncup is called automatically whenever there is a change in the window. The return value is unspecified.

[Procedure] syncdown win

The syncdown routine touches each location in win that has been touched in any of its ancestor windows. This routine is called by refresh, so it should almost never be necessary to call it manually.

[Procedure] cursyncup win

The routine cursyncup updates the current cursor position of all the ancestors of the window to reflect the current cursor position of the window. The return value is unspecified.

5.2.40 Terminal resizing

[Procedure] resizeterm lines columns The function resizeterm resizes the standard and current windows to the specified dimensions lines and columns. If the windows are increased in size, the extended areas are blank-filled. It returns #t on success or #f on failure. It could fail for memory allocation failures or for dimensions that are too small.

5.3 The Function Key Label Library: (ncurses slk)

The slk functions in the (ncurses slk) library manipulate the set of soft function-key labels that exist on many terminals. For those terminals that do not have soft labels, curses takes over the bottom line of stdscr, reducing the size of stdscr (and thus the return value of the procedure lines). curses standardizes on eight labels of up to eight characters each. In addition to this, the ncurses implementation supports a mode where it simulates 12 labels of up to five characters each. This is most common for today’s PC like end-user devices. Please note that ncurses simulates this mode by taking over up to two lines at the bottom of the screen, it doesn’t try to use any hardware support for this mode.

This API is definitely in flux.

[Procedure] slk-init fmt

The slk-init routine must be called before initscr or newterm is called. If initscr eventually uses a line from stdscr to emulate the soft labels, then fmt determines how the labels are arranged on the screen. Setting fmt to 0 indicates a 3-2-3 arrangement of the labels, 1 indicates a 4-4 arrangement and 2 indicates the PC-like 4-4-4 mode. If fmt is set to 3, it is again the PC like 4-4-4 mode, but in addition an index line is generated, helping the user to identify the key numbers easily. It returns #t if fmt is a valid value, or #f otherwise.

[Procedure] slk-set labnum label fmt The slk-set routine requires labnum to be a label number, from 1 to 8 (resp. 12); label must be the string to be put on the label, up to eight (resp. five) characters in length. fmt is either 0, 1, or 2, indicating whether the label is to be left-justified, centered, or right-justified, respectively, within the label. The procedure returns #f if the terminal or the softkeys were not initialized, or the labnum parameter is outside the range of label counts, or if the format parameter is outside the range 0..2, or if memory for the labels cannot be allocated. Otherwise it return #t

[Procedure] slk-refresh
[Procedure] slk-noutrefresh

These procedure do a refresh or noutrefresh operation on the slk window.

[Procedure] slk-label labnum

The slk-label routine returns the current label for label number labnum, with leading and trailing blanks stripped.

[Procedure] slk-clear

This routine clears the soft labels from the screen.

[Procedure] slk-restore

This procedure restores the soft labels after an slk-clear has been performed.

[Procedure] slk-touch

This routine forces all the soft labels to be output next time slk-noutrefresh is performed.

[Procedure] slk-attr-on! attrs [Procedure] slk-attr-off! attrs

[Procedure] slk-attr-set! attrs color-pair

[Procedure] slk-color! color-pair [Procedure] slk-attr

The slk-attr-on!, slk-attr-set!, and slk-attr-off! routines correspond to

attr-on!, attr-set!, attr-off!. They have an effect only if soft labels are simulated on the bottom line of the screen. The default highlight for soft keys is

A_STANDOUT.

The slk-color! routine corresponds to color-set!. It has an effect only if soft labels are simulated on the bottom line of the screen.

The procedure slk-attr returns the attributes and color pair currently set for the soft keys. They are returned as a list of two elements. The first element is a bitmap of attributes; the second element is a color pair number.

5.4 The curses panel library: (ncurses panel)

These are the functions in the (ncurses panel) module.

Panels are curses windows with the added feature of depth. Panel functions allow the use of stacked windows and ensure the proper portions of each window and the curses stdscr window are hidden or displayed when panels are added, moved, modified or removed. The set of currently visible panels is the stack of panels. The stdscr window is beneath all panels, and is not considered part of the stack.

A window is associated with every panel. The panel routines enable you to create, move, hide, and show panels, as well as position a panel at any desired location in the stack.

Panel routines are a functional layer added to curses, make only high-level curses calls, and work anywhere terminfo curses does.

[Procedure] new-panel win

This procedure allocates a PANEL structure, associates it with win, places the panel on the top of the stack (causes it to be displayed above any other panel) and returns the new panel.

[Procedure] update-panels

This procedure refreshes the virtual screen to reflect the relations between the panels in the stack, but does not call doupdate to refresh the physical screen. Use this function and not refresh or noutrefresh. update-panels may be called more than once before a call to doupdate, but doupdate is the procedure responsible for updating the physical screen.

[Procedure] del-panel pan

This procedure removes the given panel from the stack and invalidates pan

[Procedure] hide-panel pan

This procedure removes the given panel from the panel stack and thus hides it from view. The panel pan is not lost, merely removed from the stack.

[Procedure] panel-hidden? pan

This procedure returns #t if the panel is in the panel stack, #f if it is not.

[Procedure] show-panel pan

This procedure makes a hidden panel visible by placing it on top of the panels in the panel stack.

[Procedure] top-panel pan

This procedure puts the given visible panel on top of all panels in the stack.

[Procedure] bottom-panel pan

This procedure puts the panel at the bottom of all panels.

[Procedure] move-panel pan starty startx This procedure moves the given panel window so that its upper-left corner is at starty, startx. It does not change the position of the panel in the stack. Be sure to use this function, not mvwin, to move a panel window.

[Procedure] replace-panel! pan window

This procedure replaces the current window of panel with window (useful, for example if you want to resize a panel; It does not change the position of the panel in the stack.

[Procedure] panel-window panel

Returns the window of a given panel

5.5 The curses menu library: (ncurses menu)

These are the functions in the (ncurses menu) module.

5.5.1 Menus Overview

Menu items are individual menu entries. Each menu item has a name – which is the short, usually single word, description of the item – and a description, which is a longer description of the menu item. Depending on the format of the menu, the description may or many not be visible. Each menu item also holds a set of options about the visibility and selectability of the item.

A menu is a collection of menu items along with options on how it operates and is presented.

5.5.2 Menu Item Procedures

[Procedure] new-item name description Returns a new allocated menu item of type with a given name and description. It can throw an error if the strings are too large or if a memory allocation error occurs.

[Procedure] set-current-item! menu item

Given a menu and an item that has been attached to that menu, this procedure sets item to be the current item. It returns an integer that will be one of E_OK, E_BAD_ARGUMENT, E_NOT_CONNECTED if not menu items are attached to this menu, or

E_SYSTEM_ERROR.

[Procedure] current-item menu

Returns the current item for menu.

[Procedure] set-top-row! menu row

This procedure sets menu’s top row to be row. row is an integer where zero indicates that the first row is the top row. This procedure will only have an effect when there are more rows in the menu than are displayed on the screen. If all the menu items can be displayed, this procedure will have no effect and return E_BAD_ARGUMENT. The number of rows and columns that a menu displays is set by set-menu-format!. It will return one of E_OK, E_BAD_ARGUMENT, E_NOT_CONNECTED if no menu items are attached to this menu, or E_SYSTEM_ERROR.

[Procedure] top-row menu

Returns the number of the top menu row being displayed, or #f on failure.

[Procedure] item-index item

This procedure returns the zero-origin index of this item in its containing menu’s item list or #f on failure.

[Procedure] item-name item

Returns, as a string, the name of the menu item item.

[Procedure] item-description item

Returns, as a string, the description of the menu item item.

[Procedure] set-item-opts! item opts [Procedure] item-opts-on! item opts [Procedure] item-opts-off! item opts

These functions set or query the bit mask for item. There is only one option bit mask, O_SELECTABLE. When this is on, the item may be selected during menu processing. It defaults to on. The procedure set-item-opts sets the bit mask of item to opts. The procedure menu-opts-on turns the bits of opts on in item. The procedure menuopts-off turns off the bits on opts in item. These routines will return E_OK on success or E_SYSTEM_ERROR on failure.

[Procedure] item-opts item

This procedure returns the options bit mask of item.

[Procedure] set-item-value! item value [Procedure] item-value item

If the menu option O_ONEVALUE is turned off, the menu can have more than one item selected simultaneously. In this case, the procedure item-value can be used to query item to see if it is selected. It returns #t or #f.

An item can be selected with the procedure set-item-value!, which takes a menu item and a boolean value. It returns E_OK, E_SYSTEM_ERROR, or E_REQUEST_DENIED.

[Procedure] item-visible? item

Returns #t if item is mapped onto the screen. If a menu has more menu items than can be displayed at once, only some of the menu items will be mapped onto the screen.

5.5.3 Menu Procedures

5.5.3.1 Menu Colors and Attributes

[Procedure] set-menu-fore! menu attr [Procedure] set-menu-back! menu attr [Procedure] set-menu-grey! menu attr

The procedure set-menu-fore! sets the colors and attributes of selected menu items. The default is A_STANDOUT.

The procedure set-menu-back! sets the colors and attributes of selectable but not currently selected menu items. Its default is A_NORMAL.

The procedure set-menu-grey! sets the colors and attributes of unselectable menu items for menus that allow multiple selections. Its default is A_UNDERLINE.

[Procedure] menu-fore menu [Procedure] menu-back menu [Procedure] menu-grey menu

These procedures return the colors and attributes of selected, selectable, and unselectable menu items respectively.

[Procedure] set-menu-pad menu

This procedure sets the character used to fill the space between the name and the description parts of a menu.

[Procedure] menu-pad menu

This procedure returns the character used to fill the space between the name and description parts of a menu.

[Procedure] item-count menu

Returns the number of menu items attached to menu.

5.5.3.2 Positioning a menu’s cursor

[Procedure] pos-menu-cursor menu

This procedure repositions the cursor the position associated with the menu’s current selected item. This is useful if other (non-menu) routines have been called and have moved the cursor away from the menu.

5.5.3.3 The menu driver

[Procedure] menu-driver menu c

Once a menu has been posted (displayed), you should funnel input events to it through menu-driver. This routine has three major input cases:

• The input is a form navigation request. Navigation request codes are constants defined in the table below, which are distinct from the key and character codes returned by wgetch.

• The input is a printable character. Printable characters (which must be positive, less than 256) are checked according to the program’s locale settings.

• The input is the KEY MOUSE special key associated with an mouse event.

The menu driver request are as follows.

REQ_LEFT_ITEM

Move left to an item.

REQ_RIGHT_ITEM

Move right to an item.

REQ_UP_ITEM

Move up to an item.

REQ_DOWN_ITEM

Move down to an item.

REQ_SCR_ULINE

Scroll up a line.

REQ_SCR_DLINE

Scroll down a line.

REQ_SCR_DPAGE

Scroll down a page.

REQ_SCR_UPAGE

Scroll up a page.

REQ_FIRST_ITEM

Move to the first item.

REQ_LAST_ITEM

Move to the last item.

REQ_NEXT_ITEM

Move to the next item.

REQ_PREV_ITEM

Move to the previous item.

REQ_TOGGLE_ITEM

Select/deselect an item.

REQ_CLEAR_PATTERN

Clear the menu pattern buffer.

REQ_BACK_PATTERN

Delete the previous character from the pattern buffer.

REQ_NEXT_MATCH

Move to the next item matching the pattern match.

REQ_PREV_MATCH

Move to the previous item matching the pattern match.

If the second argument c is a printable character, the code appends it to the pattern buffer and attempts to move to the next item matching the new pattern. If there is no such match, menu-driver returns E_NO_MATCH and deletes the appended character from the buffer. If the second argument c is one of the above pre-defined requests, the corresponding action is performed. If the second argument is the KEY_MOUSE special key, the associated mouse event is translated into one of the above pre-defined requests. Currently only clicks in the user window (e.g. inside the menu display area or the decoration window) are handled. If you click above the display region of the menu:

• a REQ SCR ULINE is generated for a single click,

• a REQ SCR UPAGE is generated for a double-click and

• a REQ FIRST ITEM is generated for a triple-click.

If you click below the display region of the menu:

• a REQ SCR DLINE is generated for a single click,

• a REQ SCR DPAGE is generated for a double-click and

• a REQ LAST ITEM is generated for a triple-click.

If you click at an item inside the display area of the menu:

• the menu cursor is positioned to that item.

• If you double-click an item a REQ_TOGGLE_ITEM is generated and E_UNKNOWN_ COMMAND is returned. This return value makes sense, because a double click usually means that an item-specific action should be returned. It is exactly the purpose of this return value to signal that an application specific command should be executed.

• If a translation into a request was done, menu driver returns the result of this request.

If you clicked outside the user window or the mouse event could not be translated into a menu request an E REQUEST DENIED is returned. The procedure menu-driver returns one of the following error codes:

E_OK The routine succeeded.

E_SYSTEM_ERROR

System error occurred (see errno).

E_BAD_ARGUMENT

Routine detected an incorrect or out-of-range argument.

E_BAD_STATE

Routine was called from an initialization or termination function.

E_NOT_POSTED

The menu has not been posted.

E_UNKNOWN_COMMAND

The menu driver code saw an unknown request code.

E_NO_MATCH

Character failed to match.

E_REQUEST_DENIED

The menu driver could not process the request.

5.5.3.4 Menu sizes

[Procedure] set-menu-format! menu rows cols

The procedure set-menu-format! sets the maximum display size of the given menu. If this size is too small to display all menu items, the menu will be made scrollable. If this size is larger than the menus subwindow and the subwindow is too small to display all menu items, post-menu will fail. The default format is 16 rows, 1 column. A zero row or column argument is interpreted as a request not to change the current value. This procedure returns E_OK, E_SYSTEM_ERROR, E_BAD_ARGUMENT if the routine detected an incorrect or out-of-range argument, or E_POSTED if the menu is already posted.

[Procedure] menu-format

This procedure returns a two-element list containing the display size of the menu in rows and columns

5.5.3.5 Menu mark strings

[Procedure] set-menu-mark! menu str

In order to make menu selections visible on older terminals without highlighting or color capability, the menu library marks selected items in a menu with a prefix string. The function set-menu-mark! sets the mark string for the given menu. Calling setmenu-mark! with a null menu item will abolish the mark string. Note that changing the length of the mark string for a menu while the menu is posted is likely to produce unhelpful behavior. The default string is “-” (a dash).

[Procedure] menu-mark menu

Returns, as a string, the prefix string used by the menu library to designate the selected item.

5.5.4 Menu creation

[Procedure] new-menu items

Given a list of items, where each item was created by the procedure new-item, the procedure new-menu returns a menu. If the menu cannot be created, an error is thrown.

5.5.5 Menu options

[Procedure] set-menu-opts! menu options [Procedure] menu-opts-on! menu options [Procedure] menu-opts-off! menu options

The function set-menu-opts! sets all the given menu’s option bits (menu option bits may be logically-OR’ed together). The function menu-opts-on turns on the given option bits, and leaves others alone. The function menu-opts-off turns off the given option bits, and leaves others alone. The following options are defined (all are on by default):

O_ONEVALUE

Only one item can be selected for this menu.

O_SHOWDESC

Display the item descriptions when the menu is posted.

O_ROWMAJOR

Display the menu in row-major order.

O_IGNORECASE

Ignore the case when pattern-matching.

O_SHOWMATCH

Move the cursor to within the item name while pattern-matching.

O_NONCYCLIC

Don’t wrap around next-item and previous-item, requests to the other end of the menu.

The procedure will return E_OK, E_SYSTEM_ERROR, or E_POSTED if the menu is already posted.

[Procedure] menu-opts

Returns the bitmask of the menu’s options.

5.5.5.1 Menu pattern buffer

[Procedure] set-menu-pattern! menu pattern

Every menu has an associated pattern match buffer. As input events that are printable characters come in, they are appended to this match buffer and tested for a match, as described in menu-driver.

The function set-menu-pattern! sets the pattern buffer for the given menu and tries to find the first matching item. If it succeeds, that item becomes current; if not, the current item does not change. It may return the following error codes: E_OK, E_BAD_ARGUMENT, E_BAD_STATE if the outline was called from an initialization or termination function, E_NO_MATCH if the character failed to match, or E_SYSTEM_ERROR.

[Procedure] menu-pattern menu

The procedure menu-pattern returns the pattern buffer of the given menu.

5.5.5.2 Writing or erasing menus from window

[Procedure] post-menu menu [Procedure] unpost-menu menu

The procedure post-menu displays a menu to its associated subwindow. To trigger physical display of the subwindow, use refresh or some equivalent curses routine (the implicit doupdate triggered by an curses input request will do). post-menu resets the selection status of all items. The procedure unpost-menu erases menu from its associated subwindow. These routines return one of the following: E_OK, E_SYSTEM_ERROR, E_BAD_ARGUMENT, E_POSTED if the menu has already been posted, E_BAD_STATE if the routine was called from an initialization or termination function, E_NO_ROOM if the menu is too large for its window, or E_NOT_POSTED if the menu has not been posted.

5.5.5.3 Control spacing between menu items

[Procedure] set-menu-spacing! menu description rows cols

[Procedure] menu-spacing menu

The procedure set-menu-spacing! sets the spacing information for the menu. The parameter description controls the number of spaces between an item name and an item description. It must not be larger than TABSIZE. The menu system puts in the middle of this spacing area the pad character. The remaining parts are filled with spaces. The parameter rows controls the number of rows that are used for an item. It must not be larger than 3. The menu system inserts the blank lines between item rows, these lines will contain the pad character in the appropriate positions. The parameter cols controls the number of blanks between columns of items. It must not be larger than TABSIZE. A value of 0 for all the spacing values resets them to the default, which is 1 for all of them.

set-menu-spacing! will return E_OK, E_POSTED if the menu is posted, or E_BAD_

ARGUMENT if one of the spacing values is out of range. The procedure menu-spacing returns the spacing info for the menu as a list of three elements: description, row, and column spacing

5.5.5.4 Associate menus with window

[Procedure] set-menu-win! menu win [Procedure] set-menu-sub! menu subwin [Procedure] menu-win menu [Procedure] menu-sub menu [Procedure] scale-menu menu

Every menu has an associated pair of curses windows. The menu window displays any title and border associated with the window; the menu subwindow displays the items of the menu that are currently available for selection. The first four functions get and set those windows. It is not necessary to set either window; by default, the driver code uses stdscr for both. The procedure scale-menureturns the minimum size required for the subwindow of menu.

set-menu-win! and set-menu-sub! will return one of the following, E_OK, E_SYSTEM_

ERROR, E_BAD_ARGUMENT, or E_POSTED if the menu has already been posted.

5.6 The curses form library: (ncurses form)

These are the functions in the (ncurses form) module.

5.6.1 Forms Overview

The form library provides terminal-independent facilities for composing form screens on character-cell terminals. The library includes: field routines, which create and modify form fields; and form routines, which group fields into forms, display forms on the screen, and handle interaction with the user.

5.6.2 Positioning a form window cursor

[Procedure] pos-form-cursor form

The function pos-form-cursor restores the cursor to the position required for the forms driver to continue processing requests. This is useful after curses routines have been called to do screen-painting in response to a form operation. It returns E_OK if the routine succeeded, E_BAD_ARGUMENT if the routine detected an incorrect or out-of-range argument, E_NOT_POSTED if the form has not been posted, or E_SYSTEM_ERROR if a system error occurred.

5.6.3 Form data

[Procedure] data-ahead? form [Procedure] data-behind? form

These procedures test if there is ahead or behind the current screen of the given form form. They return #t or #f. This could be because a text entry has been declared to have variable size (not O STATIC) and not all text from the entry is being displayed.

5.6.4 Command-processing loop of the form system

[Constant] %is-form-driver-wide

This constant is #t if the form-driver function is capable of processing characters with representations that are greater than 8 bits.

[Procedure] form-driver form c

Once a form has been posted (displayed), you should funnel input events to it through form-driver. This routine has three major input cases:

• The input is a form navigation request. Navigation request codes are constants defined below, which are distinct from the key and character codes returned by

    wgetch.

• The input is a printable character.

• The input is the KEY MOUSE special key associated with an mouse event.

The form driver requests are as follows:

Name  Description
REQ_NEXT_PAGE  Move to the next page.
REQ_PREV_PAGE  Move to the previous page.
REQ_FIRST_PAGE  Move to the first page.
REQ_LAST_PAGE  Move to the last field.
REQ_NEXT_FIELD  Move to the next field.
REQ_PREV_FIELD  Move to the previous field.
REQ_FIRST_FIELD  Move to the first field.
REQ_LAST_FIELD  Move to the last field.
REQ_SNEXT_FIELD  Move to the sorted next field.
REQ_SPREV_FIELD  Move to the sorted previous field.
REQ_SFIRST_FIELD  Move to the sorted first field.
REQ_SLAST_FIELD  Move to the sorted last field.
REQ_LEFT_FIELD  Move left to a field.
REQ_RIGHT_FIELD  Move right to a field.
REQ_UP_FIELD  Move up to a field.
REQ_DOWN_FIELD  Move down to a field.
REQ_NEXT_CHAR  Move to the next char.
REQ_PREV_CHAR  Move to the previous char.
REQ_NEXT_LINE  Move to the next line.
REQ_PREV_LINE  Move to the previous line.
REQ_NEXT_WORD  Move to the next word.
REQ_PREV_WORD  Move to the previous word.
REQ_BEG_FIELD  Move to the beginning of the field.
REQ_END_FIELD  Move to the end of the field.
REQ_BEG_LINE  Move to the beginning of the line.
REQ_END_LINE  Move to the end of the line.
REQ_LEFT_CHAR  Move left in the field.
REQ_RIGHT_CHAR  Move right in the field.
REQ_UP_CHAR  Move up in the field.
REQ_DOWN_CHAR  Move down in the field.
REQ_NEW_LINE  Insert or overlay a newline.
REQ_INS_CHAR  Insert a blank at the cursor.
REQ_INS_LINE  Insert a blank line at the cursor.
REQ_DEL_CHAR  Delete character at the cursor.

REQ_DEL_PREV  Delete character before the cursor.
REQ_DEL_LINE  Delete line at the cursor.
REQ_DEL_WORD  Delete blank-delimited word at the cursor.
REQ_CLR_EOL  Clear to end of line from cursor.
REQ_CLR_EOF  Clear to end of field from cursor.
REQ_CLR_FIELD  Clear the entire field.
REQ_OVL_MODE  Enter overlay mode.
REQ_INS_MODE  Enter insert mode.
REQ_SCR_FLINE  Scroll the field forward a line.
REQ_SCR_BLINE  Scroll the field backward a line.
REQ_SCR_FPAGE  Scroll the field forward a page.
REQ_SCR_BPAGE  Scroll the field backward a page.
REQ_SCR_FHPAGE  Scroll the field forward half a page.
REQ_SCR_BHPAGE  Scroll the field backward half a page.
REQ_SCR_FCHAR  Scroll the field forward a character.
REQ_SCR_BCHAR  Scroll the field backward a character.
REQ_SCR_HFLINE  Horizontal scroll the field forward a line.
REQ_SCR_HBLINE  Horizontal scroll the field backward a line.
REQ_SCR_HFHALF  Horizontal scroll the field forward half a line.
REQ_SCR_HBHALF  Horizontal scroll the field backward half a line.

 REQ_VALIDATION  Validate field.

REQ_NEXT_CHOICE  Display next field choice.
REQ_PREV_CHOICE  Display previous field choice.

If the second argument is a printable character, the driver places it in the current position in the current field. If it is one of the forms requests listed above, that request is executed. It is important to note that when characters are Unicode codepoints greater than U+00FF, the form driver will only use them for versions of the wide ncurses form library that are 2014 or newer. For older version of ncurses, only printable characters (which must be positive, less than 256) are allowed. The %is-form-driver-wide constant can be used to check if form-driver can handle characters whose numerical representations are greater than 256. If the second argument is the KEY_MOUSE special key, the associated mouse event is translated into one of the above pre-defined requests. Currently only clicks in the user window (e.g. inside the form display area or the decoration window) are handled. If you click above the display region of the form:

• a REQ PREV FIELD is generated for a single click,

• a REQ PREV PAGE is generated for a double-click and

• a REQ FIRST FIELD is generated for a triple-click.

If you click below the display region of the form:

• a REQ NEXT FIELD is generated for a single click,

• a REQ NEXT PAGE is generated for a double-click and

• a REQ LAST FIELD is generated for a triple-click.

If you click at an field inside the display area of the form:

• the form cursor is positioned to that field.

• If you double-click a field, the form cursor is positioned to that field and E_ UNKNOWN_COMMAND is returned. This return value makes sense, because a double click usually means that an field-specific action should be returned. It is exactly the purpose of this return value to signal that an application specific command should be executed.

• If a translation into a request was done, form driver returns the result of this request.

If you clicked outside the user window or the mouse event could not be translated into a form request an E_REQUEST_DENIED is returned. If the second argument is neither printable nor one of the above predefined form requests, the driver assumes it is an application-specific command and returns E_ UNKNOWN_COMMAND. Application-defined commands should be defined relative to MAX_ COMMAND, the maximum value of these pre-defined requests. form-driver returns one of the following error codes: E_OK if the routine succeeded, E_BAD_ARGUMENT if the routine detected an incorrect or out-of-range argument, E_ BAD_STATE if the routine was called from an initialization or termination function, E_NOT_POSTED if the form has not been posted, E_INVALID_FIELD if the contents of field is invalid, E_REQUEST_DENIED if the form driver could not process the request, E_ SYSTEM_ERROR if a system error occurred, and E_UNKNOWN_COMMAND if the form driver code saw an unknown request code.

5.6.5 Making or breaking connections between forms and fields

[Procedure] set-form-fields form fields

This procedure takes a fields, which is a scheme list of field types, and assigns them to form. It returns an integer that will have the value E_OK on success, E_BAD_ARGUMENT, E_ CONNECTED if the fields are already connected to the form, E_POSTED if the form is already posted, or E_SYSTEM_ERROR if a system error occurred.

[Procedure] field-count form

Returns the number of fields in form.

[Procedure] move-field field row col

This moves the screen position of field to row, col. The field must not be connected to a form. It returns an integer that will have the value E_OK on success, E_BAD_ARGUMENT, E_ CONNECTED if the fields are already connected to the form, E_POSTED if the form is already posted, or E_SYSTEM_ERROR if a system error occurred.

5.6.6 Color and attribute control for form fields

[Procedure] set-field-fore! field attr [Procedure] set-field-back! field attr

The procedure set-field-fore! sets the attributes of the contents of a field, and set-field-back! set the attributes of the unused portion of a field. They return the constants E_OK on success, E_BAD_ARGUMENT, or E_SYSTEM_ERROR.

[Procedure] field-fore field [Procedure] field-back field

The procedure field-fore returns the attributes of the contents of a field, and field-back returns the attributes of the unused portion of a field.

[Procedure] set-field-pad! field pad

The procedure set-field-pad! sets the character used as blank in the field to the codepoint pad. They return the constants E_OK on success, E_BAD_ARGUMENT, or E_SYSTEM_ERROR.

[Procedure] field-pad field

The procedure field-pad returns the codepoint used as the blank in field.

5.6.7 Field buffer control

There are functions to set and get the text in a field, as well as to check if that field has been modified.

[Procedure] set-field-buffer! field buf value

The function set-field-buffer! sets the numbered buffer of the given field to contain a given string. Buffer 0 is the displayed value of the field; other numbered buffers may be allocated by applications through the nbuf argument of new-field but are not manipulated by the forms library.

[Procedure] field-buffer field buffer

Returns a string that is the contents of the field. buffer zero is the visible buffer on the screen. A field may have other buffers if so allocated when the field was created

using new-field.

[Procedure] field-status? field

The returns the field status of field. The field status is set to a non-zero value whenever the field changes.

[Procedure] set-field-status! field status

Sets the field status of field to status, a boolean.

[Procedure] set-max-field field max

This procedure sets the maximum size for a dynamic field. An argument of zero turns off any maximum size threshold for that field.

5.6.8 Retrieving field characteristics

[Procedure] field-info field

The procedure returns a six-element list of information about the field. The elements of the list are: height, width, row of upper-left corner, column of upper-left corner, number of off-screen rows, and number of working buffers. This information was stored when the field was created, and rows and columns may not be accurate.

[Procedure] dynamic-field-info field

This procedure returns of the actual size of the field and its maximum size as a list. The elements of the list are: rows, column, and max size.

5.6.9 Retrieving field characteristics

[Procedure] set-field-just! field justification This procedure sets the justification attribute of a field to one of NO_JUSTIFICATION,

JUSTIFY_RIGHT, JUSTIFY_LEFT, or JUSTIFY_CENTER.

[Procedure] field-just field

This procedure returns the justification attribute of field. It is one of

NO_JUSTIFICATION, JUSTIFY_RIGHT, JUSTIFY_LEFT, or JUSTIFY_CENTER.

5.6.10 Creating and destroying form fields

[Procedure] new-field height width toprow leftcol offscreen nbuffers The function new-field allocates a new field and initializes it from the parameters given: height, width, row of upper-left corner, column of upper-left corner, number off-screen rows, and number of additional working buffers.

[Procedure] dup-field field toprow leftcol The function dup-field duplicates a field at a new location. Most attributes (including current contents, size, validation type, buffer count, growth threshold, justification, foreground, background, pad character, options, and user pointer) are copied. Field status and the field page bit are not copied.

[Procedure] link-field field toprow leftcol This is like dup-field except that the new field shares buffers with its parent. Attribute data is separate.

[Procedure] free-field field

This explicitly releases the buffer in field, instead of waiting for the garbage collector to take care of it.

5.6.11 Setting and getting field options

[Procedure] set-field-opts! field opts [Procedure] field-opts-on! field opts [Procedure] field-opts-off! field opts [Procedure] field-opts field

The function set-field-opts! sets all the given field’s option bits (field option bits may be logically-OR’ed together). The function field-opts-on! turns on the given option bits, and leaves others alone. The function field-opts-off! turns off the given option bits, and leaves others alone. The function field-opts returns the field’s current option bits.

The following options are defined. All are on by default

O_VISIBLE

The field is displayed. If this option is off, display of the field is suppressed.

O_ACTIVE The field is visited during processing. If this option is off, the field will not be reachable by navigation keys. Please notice that an invisible field appears to be inactive also.

O_PUBLIC The field contents are displayed as data is entered.

O_EDIT The field can be edited.

O_WRAP Words that do not fit on a line are wrapped to the next line. Words are blankseparated.

O_BLANK The field is cleared whenever a character is entered at the first position.

O_AUTOSKIP

Skip to the next field when this one fills.

O_NULLOK Allow a blank field.

O_STATIC Field buffers are fixed to field’s original size. Turn this option off to create a dynamic field.

O_PASSOK Validate field only if modified by user.

5.6.12 Data type validation for fields

These functions set a field to be able to contain only a certain type of input. The user will not be able to move off the field if it contains invalid input.

[Procedure] set-field-type! field type … [Procedure] field-type field

The procedure set-field-type! declares a data type for a given form field. The procedure field-type returns the declared data type. The input type is a symbol that is one of the following list. This is the type checked by validation functions. The predefined types are as follows:

TYPE_ALNUM

Alphanumeric data. Requires a third int argument, a minimum field width.

TYPE_ALPHA

Character data. Requires a third int argument, a minimum field width.

TYPE_ENUM

Accept one of a specified set of strings. Requires a third argument pointing to a list of strings; a fourth integer flag argument to enable casesensitivity; and a fifth int flag argument specifying whether a partial match must be a unique one (if this flag is off, a prefix matches the first of any set of more than one list elements with that prefix). Please notice that the string list is not copied, only a reference to it is stored in the field. So you should avoid using a list that lives in automatic variables on the stack.

TYPE_INTEGER

Integer data. Requires a third int argument controlling the precision, a fourth long argument constraining minimum value, and a fifth long constraining maximum value. If the maximum value is less than or equal to the minimum value, the range is simply ignored.

TYPE_NUMERIC

Numeric data (may have a decimal-point part). Requires a third int argument controlling the precision, a fourth double argument con- straining minimum value, and a fifth double constraining maximum value. If your system supports locales, the decimal point character to be used must be the one specified by your locale. If the maximum value is less than or equal to the minimum value, the range is simply ignored.

TYPE_REGEXP

Regular expression data. Requires a regular expression (char *) third argument; the data is valid if the regular expression matches it. Regular expressions are in the format of regcomp and regexec. Please notice that the regular expression must match the whole field. If you have for example an eight character wide field, a regular expression "^[0-9]*$" always means that you have to fill all eight positions with digits. If you want to allow fewer digits, you may use for example "^[0-9]* *$" which is good for trailing spaces (up to an empty field), or "^*[0-9]* *$" which is good for leading and trailing spaces around the digits.

TYPE_IPV4

An Internet Protocol Version 4 address. This requires no additional argument. It is checked whether or not the buffer has the form a.b.c.d, where a,b,c and d are numbers between 0 and 255. Trailing blanks in the buffer are ignored. The address itself is not validated. Please note that this is an ncurses extension. This field type may not be available in other curses implementations.

5.6.13 Creating forms

[Procedure] new-form fields

Creates a new form given a list that contains fields. The fields are created using

new-field.

5.6.14 Form pagination functions

[Procedure] set-new-page field new-page-flag This procedure sets or resets a flag marking the given field as beginning a new page on its form.

[Procedure] new-page? field

Returns a flag indicating if the given field is the start of a new page on its form.

5.6.15 Setting and getting form options

[Procedure] set-form-opts form opts [Procedure] form-opts-on! form opts [Procedure] form-opts-off! form opts

The function set-form-opts sets all the given form’s option bits (form option bits may be logically-OR’ed together).

The function form-opts-on! turns on the given option bits, and leaves others alone.

The function form-opts-off! turns off the given option bits, and leaves others alone.

[Procedure] form-opts form

This returns the current option bits for the form.

The following options are defined, and all are on by default.

O_NL_OVERLOAD

Overload the REQ_NEW_LINE forms driver request so that calling it at the end of a field goes to the next field.

O_BS_OVERLOAD

Overload the REQ_DEL_PREV forms driver request so that calling it at the beginning of a field goes to the previous field.

5.6.16 Setting the page number and field

[Procedure] set-current-field! form field

This procedure sets the current field on the form.

[Procedure] field-index form

This returns the index into the field array that is current for the form.

[Procedure] set-form-page form n

Sets the page of the form to n.

[Procedure] form-page form

Returns the current form’s displayed page number.

5.6.17 Writing or erasing forms from associated subwindows

[Procedure] post-form form

This procedure displays a form to its associated subwindow. To trigger physical display of the subwindow, use refresh or some equivalent routine.

[Procedure] unpost-form form

This erases the form from its associated subwindow.

5.6.18 Printable form request names

[Procedure] form-request-name request

[Procedure] form-request-by-name name

These procedures convert a string containing the printable name of a form request to its integer value and vice versa.

5.6.19 Form windows

[Procedure] set-form-win form win [Procedure] form-win form [Procedure] set-form-sub form sub [Procedure] form-sub form [Procedure] scale-form form

Every form has an associated pair of curses windows. The form window displays any title and border associated with the window; the form sub-window displays the items of the form that are currently available for selection. The first four functions get and set those windows. It is not necessary to set either window; by default, the driver code uses stdscr for both. The function scale-form returns the a list of two elements: minimum size required for the subwindow of form.

Index 114

Index

#

#<field> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 #<file> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 #<form> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 #<item> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 #<menu> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 #<panel> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 #<screen>. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 57, 58 #<window> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

$

$TERM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

%

%filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57, 92

%guile-ncurses-shell-read-port. . . . . . . . . . . . . 10
%guile-ncurses-shell-stdscr . . . . . . . . . . . . . . . . . 9
%guile-ncurses-shell-write-port . . . . . . . . . . . 10

%is-form-driver-wide . . . . . . . . . . . . . . . . . . 104, 106 %is-pad-broken . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 %is-subwin-broken. . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

%scheme-char-from-c-char . . . . . . . . . . . . . . . . . . . 61
%scheme-char-from-c-wchar . . . . . . . . . . . . . . . . . . 62

%scheme-char-to-c-char . . . . . . . . . . . . . . . . . . . . . . 61

%scheme-char-to-c-wchar. . . . . . . . . . . . . . . . . . . . . 62

%ucs4-chars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 %wide-ncurses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

7

7-bit characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

8

8-bit characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

A

addch. . . . . . . . . . . . . . . . . . . . 56, 62, 65, 67, 81, 85, 88 addchstr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 addstr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 65 assume-default-colors . . . . . . . . . . . . . . . . . . . 70, 71 attr-get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 attr-off! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 attr-on! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 attr-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 16, 17, 65 Attributes, adding to a character . . . . . . . . . . . . . . 12 Attributes, querying . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Attributes, setting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

B

Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 Background rendition. . . . . . . . . . . . . . . . . . . . . . . . . . 68 baudrate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Baudrate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 beep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67, 72 Beeping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 bkgd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 bkgdset! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67, 69 blink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 blink-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 blink-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 Blocking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79, 80 bold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 bold-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 bold-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 border . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 bottom-panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 Buffering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

C

can-change-color?. . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 Carriage return . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 cbreak! . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55, 72, 78, 79 Character buffering . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Characters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Characters, output . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 chgat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 clear . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Clearing the screen. . . . . . . . . . . . . . . . . . . . . . . . . 69, 85 clearok! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69, 85 clrtobot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69, 85 clrtoeol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 69 color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 Color capabilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 Color capability . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 Color pairs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 25, 66 Color, default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 color-content . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 color-pair . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25, 66, 69 color-pairs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 color-set!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 COLOR_BLACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 COLOR_BLUE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 COLOR_CYAN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 COLOR_GREEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 COLOR_MAGENTA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 COLOR_RED . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 COLOR_WHITE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 COLOR_YELLOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 colors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70, 71 Colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Colors, default. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 Colors, defining. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 Colors, initialization . . . . . . . . . . . . . . . . . . . . . . . 24, 70 Colors, manipulation . . . . . . . . . . . . . . . . . . . . . . . . . . 69 Colors, modifying. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 cols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Complex characters, modifying . . . . . . . . . . . . . . . . 61 Complex characters, querying . . . . . . . . . . . . . . . . . . 61 Control character processing . . . . . . . . . . . . . . . 78, 80 Coordinates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Copying windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 copywin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 current-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 curs-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Curses mode, exiting . . . . . . . . . . . . . . . . . . . . . . . . . . 58 Curses, dumping to disk . . . . . . . . . . . . . . . . . . . . . . . 89 Curses, ending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57, 82 Curses, initialization . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Curses, Initialization. . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Curses, termination. . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 curses-error? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 curses-version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

curses-wrong-type-arg-error:arg . . . . . . . . . . . 59
curses-wrong-type-arg-error:expected-type

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

curses-wrong-type-arg-error? . . . . . . . . . . . . . . . 59

Cursor, hiding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Cursor, positioning . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 Cursor, visiblity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 cursyncup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

D

data-ahead? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 data-behind? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 def-prog-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 def-shell-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 Default colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 define-key. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 del-panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 delay-output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 Delaying output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 delch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 72 deleteln . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 delscreen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 delwin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 derwin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 dim. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 dim-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 dim-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 doupdate . . . . . . . . . . . . . . . . . . . . . . . . . . . 58, 82, 88, 89 Dumping to disk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 dup-field. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 dupwin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 dynamic-field-info . . . . . . . . . . . . . . . . . . . . . . . . . 108

E

echo! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 echochar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Echoing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 15, 78 endwin. . . . . . . . . . . . . . . . . . . . . . . . . . 55, 58, 59, 82, 89 erase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 erasechar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Erasing the screen . . . . . . . . . . . . . . . . . . . . . . . . . 69, 85 Exiting curses mode . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

F

field-back . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 field-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 field-count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 field-fore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 field-index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 field-info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 field-just . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 field-opts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 field-opts-off!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 field-opts-on! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 field-pad. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 field-status? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 field-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 flash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 Flashing the screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 Flushing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 Flushing input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 flushinp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 fopen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 form-driver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 form-opts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 form-opts-off! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 form-opts-on! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 form-page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 form-request-by-name . . . . . . . . . . . . . . . . . . . . . . . 112 form-request-name . . . . . . . . . . . . . . . . . . . . . . . . . . 112 form-sub . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 form-win . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 free-field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 Function keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 26

G

getbegx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 getbegy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 getbegyx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 getbkgd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 getch . . . . . . . . . . . . . . . . . . . . . . . 72, 73, 77, 78, 79, 80 getcurx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 getcury . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 getmaxx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 getmaxy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 getmaxyx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 getmouse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 getnstr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 getparx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 getpary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 getparyx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 getscrreg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 getwin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89, 92 getyx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 guile-ncurses-shell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

H

Half-delay mode . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 79 halfdelay!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 has-colors? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 has-ic? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 has-il? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 has-key? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 has-mouse?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Hello, world . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 hide-panel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 hline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 horizontal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 horizontal-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 horizontal-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

I

idcok! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 idlok!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85, 86 immedok! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 inch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 78 inchnstr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 init-color! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 init-pair!. . . . . . . . . . . . . . . . . . . . . . . . . 66, 69, 70, 71 Initialization . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 10, 57 Initializing the terminal . . . . . . . . . . . . . . . . . . . . . . . . . 8 initscr …. 8, 10, 55, 57, 58, 69, 70, 81, 89, 91, 94 Input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Input, blocking . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79, 80 Input, flushing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80, 92 Inputting characters . . . . . . . . . . . . . . . . . . . . . . . 15, 72 insch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 insdelln . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Insert character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Insert lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 insertln . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 insstr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 instr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 intrflush! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56, 79 inverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 inverse-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 inverse-on. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 invis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 invis-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 invis-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 is-cleared? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 is-idcok? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 is-idlok? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 is-immedok? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81, 86 is-keypad?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 is-leaveok? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 is-linetouched? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 is-meta? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 is-nodelay? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 is-notimeout? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 is-pad? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 is-scrollok? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 is-subwin?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 is-wintouched? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 isendwin? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 item-count. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 item-description. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 item-index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 item-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 item-opts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 item-opts-off! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 item-opts-on! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 item-value. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 item-visible? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

K

Key names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 key-defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 key-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Keycodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Keycodes, customizing . . . . . . . . . . . . . . . . . . . . . . . . . 76 Keycodes, lookup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 keyname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 keypad! . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56, 72, 73, 79 Keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 Keys, custom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 Keys, unmapped. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 killchar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90

L

leaveok! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85, 88 left . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 left-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 left-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 Line buffering . . . . . . . . . . . . . . . . . . . . . . 10, 78, 79, 80 Line feed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 Line graphics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 link-field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 longname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 low. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 low-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 low-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

M

mcprint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 menu-back . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 menu-driver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 menu-fore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 menu-format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 menu-grey . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 menu-mark. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 menu-opts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 menu-opts-off! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 menu-opts-on! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 menu-pad . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 menu-pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 menu-spacing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 menu-sub . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 menu-win . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 meta! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 mouse-trafo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 mouseinterval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 mousemask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56, 84 move-field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 move-panel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 mvderwin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 mvwin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

N

napms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Ncurses, version information . . . . . . . . . . . . . . . . . . . 92 new-field. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 new-form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 new-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 97 new-menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 102 new-page?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 new-panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 95 newpad . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 newterm . . . . . . . . . . . . . . . . . . . . . . 7, 55, 57, 58, 81, 94 newwin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 92 nl!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 nocbreak! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72, 78, 79 nodelay! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79, 80 noecho! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55, 72, 78 nonl!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56, 86 noqiflush!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 noraw! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 normal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 normal-off. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 normal-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 notimeout!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 noutrefresh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87, 88

O

Opaque types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Output procedures . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Output, delaying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 Output, flushing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 Outputting characters . . . . . . . . . . . . . . . . . . 12, 62, 88 Outputting strings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Outputting wide characters . . . . . . . . . . . . . . . . . . . . 13 overlay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 Overlaying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 Overlaying windows . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 overwrite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86

P

Pads . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 Pads, creating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 Pads, refreshing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 pair-content . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 pair-number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 panel-hidden? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 panel-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 pechochar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 pnoutrefresh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 pos-form-cursor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 pos-menu-cursor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 post-form. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 post-menu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 prefresh. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86, 87, 88 Print procedures. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 protect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 protect-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 protect-on. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 putwin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89, 92

Q

qiflush! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 Quitting Curses. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

R

raw! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78, 80 Redrawing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 Redrawing windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 redrawln . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 redrawwin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 refresh . . 56, 57, 58, 62, 69, 72, 84, 85, 87, 88, 91,

93 Refreshing windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Rendition. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 replace-panel! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 reset-prog-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 reset-shell-mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 resetty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 resizeterm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76, 94 Resizing windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 right-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 right-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 ripoffline. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

S

savetty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 scale-form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 scale-menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 scr-dump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 scr-init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 scr-restore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 scr-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 Screens, clearing . . . . . . . . . . . . . . . . . . . . . . . . . . . 69, 85 Screens, creation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 Screens, erasing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69, 85 Screens, freeing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Screens, switching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 scrl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 scroll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Scrolling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85, 86, 90 scrollok! . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 85, 86, 90 set-current-field! . . . . . . . . . . . . . . . . . . . . . . . . . 112 set-current-item!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 set-field-back!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 set-field-buffer! . . . . . . . . . . . . . . . . . . . . . . . . . . 108 set-field-fore!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 set-field-just!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 set-field-opts!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 set-field-pad! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 set-field-status! . . . . . . . . . . . . . . . . . . . . . . . . . . 108 set-field-type!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 set-form-fields. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 set-form-opts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 set-form-page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 set-form-sub . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 set-form-win . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 set-item-opts! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 set-item-value! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 set-max-field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 set-menu-back! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 set-menu-fore! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 set-menu-format!. . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 set-menu-grey! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 set-menu-mark! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 set-menu-opts! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 set-menu-pad . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 set-menu-pattern! . . . . . . . . . . . . . . . . . . . . . . . . . . 102 set-menu-spacing! . . . . . . . . . . . . . . . . . . . . . . . . . . 103 set-menu-sub! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 set-menu-win! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 set-new-page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 set-tabsize! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 set-term . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 58 set-top-row! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 set-xchar-attr! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 set-xchar-chars!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 set-xchar-color!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 setscrreg!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 show-panel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 Signals, disabling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 slk-attr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 slk-attr-off! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 slk-attr-on! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 slk-attr-set! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 slk-clear . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 slk-color!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 slk-init. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57, 84, 94 slk-label . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 slk-noutrefresh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 slk-refresh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 slk-restore . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 slk-set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 slk-touch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 SMOBs used in Curses. . . . . . . . . . . . . . . . . . . . . . . . . . 7 standend! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 standout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 standout! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 standout-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 standout-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 start-color! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 stdscr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56, 57 subpad . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 subwin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87, 93 Subwindows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 Switching terminals. . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 syncdown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 syncok! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 syncup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

T

tabsize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 term-attrs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 termattrs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 termcap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Terminal capabilities, color . . . . . . . . . . . . . . . . . 25, 71 Terminal capabilities, insert. . . . . . . . . . . . . . . . . . . . 90 Terminal capabilities, keys . . . . . . . . . . . . . . . . . . . . . 76 Terminal initialization . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Terminal modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Terminal modes, saving . . . . . . . . . . . . . . . . . . . . . . . . 81 Terminal, baudrate . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Terminals, resizing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 Terminals, switching . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 terminfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 termname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 timeout! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 top. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 top-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 top-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 top-panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 top-row . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 touchline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87, 91, 93 touchwin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87, 91, 93 tset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Typeahead . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 typeahead!. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

U

unctrl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 underline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 underline-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 underline-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 ungetch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 ungetmouse. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 Unicode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 unpost-form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 unpost-menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 untouchwin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 update-panels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 Updating automatically. . . . . . . . . . . . . . . . . . . . . . . . 85 use-default-colors . . . . . . . . . . . . . . . . . . . . . . . 70, 71 use-env . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57, 91

V

Version information. . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 vertical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 vertical-off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 vertical-on . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 vline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

W

wenclose? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 wgetch. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73, 79 Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Windows, background. . . . . . . . . . . . . . . . . . . . . . 67, 68 Windows, copying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 Windows, creation . . . . . . . . . . . . . . . . . . . . . . . . . 20, 92 Windows, derived. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 Windows, duplication. . . . . . . . . . . . . . . . . . . . . . . . . . 93 Windows, moving . . . . . . . . . . . . . . . . . . . . . . . . . . 32, 92 Windows, overlaying . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 Windows, redrawing . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 Windows, resizing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 Writing to disk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 wscrl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 wtouchln . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

X

xchar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 xchar-attr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 xchar-chars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 xchar-color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 xchar? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61